### v1

# Non-local planelike minimizers and $\Gamma$-convergence of periodic energies to a local anisotropic perimeter

Abstract.
We investigate a homogenization problem related to a non-local interface energy with a periodic forcing term. We show the existence of planelike minimizers for such energy.
Moreover, we prove that, under suitable assumptions on the non-local kernel and the external field, the sequence of rescaled energies $\Gamma$-converges to a suitable local anisotropic perimeter, where the anisotropy is defined as the limit of the normalized energy of a planelike minimizer in larger and larger cubes (i.e., what is called in jargon “stable norm”).
To obtain this, we also establish several auxiliary results, including: the minimality of the level sets of the minimizers, explicit bounds on the oscillations of the minimizers, density estimates for almost minimizers, and non-local perimeter estimates in the large.
1. Introduction
Planelike minimizers are objects (such as surfaces or interfaces) that minimize a given energy functional and lie at a bounded distance from a hyperplane. They tipically occur in periodic media, where the energy density repeats periodically in space, allowing for suitable cancellations of the inhomogeneous minutiae of the environment that produce stable structures which, albeit not being perfectly flat, appear as flat at a large scale. In this spirit, planelike minimizers show how a possibly complicated structure at a fine scale behaves like a simple, averaged object in the large, and this phenomenon in turn provides an essential building block to understand homogenization.
Moreover, planelike minimizers link the microscopic features of a medium to its macroscopic behavior, often reducing the analysis to that of “effective energies”, called in jargon “stable norms”, which play the role of homogenized “surface tensions” capturing the effective anisotropy of the medium at a large scale (see below for further details).
Also, when planelike minimizers possess some geometric organization (such as foliations or laminations of space), they can help describe the global geometry of minimizers.
In certain variational settings arising in the theory of dynamical systems (like twist maps or Frenkel-Kontorova models), planelike minimizers correspond to Aubry-Mather sets, namely to configurations that minimize action and exhibit periodic and quasi-periodic order (see [MR670747, MR719634, MR2356117]). In analogy to this, planelike minimizers have been constructed for elliptic functionals and minimal surfaces (see [MR0847308, MR1852978, MR2197072, MR2542727]), phase transition models and partial differential equations (see [MR2099113, MR2342272, MR2809349]), fluid jets (see [MR2126143]), statistical mechanics and spin systems (see [MR1620543, MR3652519]).
In this paper, we investigate a homogenization problem related to a non-local interface energy with a periodic forcing term, extending the results presented in [chambolle_thouroude] to a non-local setting. Considering some assumptions on the non-local kernel $K$ and the external field $g$, we show the existence of planelike minimizers for such energy, i.e. minimizers that are at bounded distance from a plane (see Definition 1.11 below). Moreover, we prove that the sequence of rescaled energies $\{\mathscr{F}_{\varepsilon}\}_{\varepsilon}$ $\Gamma$-converges, as $\varepsilon$ approaches 0, to a suitable local anisotropic perimeter $\mathscr{F}_{\phi}$. The anisotropy $\phi$ is also known as “stable norm” (see [MR3223561]), and is defined as the limit of the (normalized) energy of a planelike minimizer in larger and larger cubes.
Adapting some ideas presented in [MR1852978, chambolle_thouroude] to our setting, we construct planelike minimizers for our energy $\mathscr{F}$ via a cell-problem. Then, we show that the stable norm $\phi$ is well-defined and we prove the main $\Gamma$-convergence result.
Setting and main assumptions
The mathematical framework adopted in this paper goes as follows. Let $Q:=(0,1)^{n}$ be the $n$-dimensional cube of side $1$, and let $g\in L^{\infty}(\mathbb{R}^{n})$ be a $\mathbb{Z}^{n}$-periodic function such that
(1.1) $\int_{Q}g(x)\,dx=0.$
In all the paper, the following assumptions on the kernel $K$ will be in force.
Suppose [^1] Hypotheses (1.2) and (1.3) are very natural assumptions to work with. Compare for instance with [MR1634336, Paragraph 1.2]. that, for all $x$, $y$, $w\in\mathbb{R}^{n}$, and for any rotation $R\in SO(n)$,
(1.2) $K(y,x)=K(x,y)=K(x+w,y+w)=K(Rx,Ry)\geqslant 0,$
and
(1.3) $\int_{\mathbb{R}^{n}}|h|K(h,0)\,dh<+\infty.$
Namely, we require that the kernel under consideration is non-negative, symmetric, translation and rotation invariant, and has some integrability properties.
Moreover, we assume that there exist parameters $s_{1}$ and $s_{2}$, and positive constants $\delta$, $\kappa_{1}$, $\kappa_{2}$ and $\kappa_{3}$, such that
$0<s_{1}<\frac{1}{2}<s_{2}<1,\qquad\kappa_{1}\leqslant\kappa_{2},$
and, for every $x$, $y\in\mathbb{R}^{n}$,
(1.4) $\kappa_{1}\frac{\chi_{(0,\delta)}(|x-y|)}{|x-y|^{n+2s_{1}}}\leqslant K(x,y)\leqslant\kappa_{2}\min\left\{\frac{1}{|x-y|^{n+2s_{1}}},\frac{1}{|x-y|^{n+2s_{2}}}\right\}$
(1.5) and $\inf_{(x,y)\in Q\times Q}K(x,y)\geqslant\kappa_{3}.$
Remark 1.1.
We point out that, in light of (1.2), the lower bound in (1.5) is equivalent to
(1.6) $K(x,y)\geqslant\kappa_{3}\chi_{(0,\sqrt{n})}(|x-y|),\quad{\mbox{for all }}(x,y)\in\mathbb{R}^{n}\times\mathbb{R}^{n}.$
Indeed, for every $x$, $y\in Q$, we have that $|x-y|\leqslant\mbox{diam}(Q)=\sqrt{n}$. Thus, (1.6) yields (1.5).
Moreover, for every $x$, $y\in\mathbb{R}^{n}$ such that $|x-y|<\sqrt{n}$, we have that $x-y\in Q$. Hence, from (1.2) and (1.5), we infer that
$K(x,y)=K(x-y,0)\geqslant\inf_{(w,z)\in Q\times Q}K(w,z)\geqslant\kappa_{3},$
which entails (1.6).
We also point out that if the lower bound in (1.4) holds true with $\delta\geqslant\sqrt{n}$, then (1.5) and (1.6) are always verified with $\kappa_{3}:=\kappa_{1}\delta^{-n-2s_{1}}$.
As model cases of kernels that satisfy (1.2), (1.3), (1.4), and (1.5), one can think about
$K_{1}(x,y):=\frac{\chi_{(0,\sqrt{n})}(|x-y|)}{|x-y|^{n+2s}},$
$K_{2}(x,y):=\frac{\chi_{(0,\delta)}(|x-y|)}{|x-y|^{n+2s}}+\chi_{[\delta,+\infty)}(|x-y|)e^{-|x-y|},$
and $K_{3}(x,y):=\frac{\chi_{(0,\delta)}(|x-y|)}{|x-y|^{n+2s}}+\frac{\chi_{[\delta,+\infty)}(|x-y|)}{|x-y|^{n+2S}},$
with $0<s<\frac{1}{2}<S<1$.
Remark 1.2.
We point out that (1.4) entails
$K(x,y)\leqslant\frac{\kappa_{2}}{|x-y|^{n+2s_{2}}},\quad{\mbox{for all }}(x,y)\in\mathbb{R}^{n}\times\mathbb{R}^{n}.$
Hence, $K$ is integrable at infinity (when $x$ and $y$ are very far apart). As we will see, this condition yields a local energy in the $\Gamma$-limit.
From now on, we will call a domain any open and bounded set, not necessarily connected.
For our purposes, we now recall the notion of $K$-nonlocal interaction between disjoint sets $A$, $B\subseteq\mathbb{R}^{n}$, that is
$\mathcal{L}_{K}(A,B):=\iint_{A\times B}K(x,y)\,dx\,dy,$
and the notion of $K$-nonlocal perimeter of a set $E\subseteq\mathbb{R}^{n}$ with respect to a Lipschitz domain $\Omega\subseteq\mathbb{R}^{n}$ (see Definition C.1 for the precise notion set with Lipschitz boundary), which is defined as
$\begin{split}\operatorname{P}_{K}(E,\Omega)&:=\iint_{\Omega_{\sharp}}\frac{1}{2}|\chi_{E}(x)-\chi_{E}(y)|\,K(x,y)\,dx\,dy\\ &=\mathcal{L}_{K}(\Omega\cap E,\Omega\cap E^{c})+\mathcal{L}_{K}(\Omega\cap E,\Omega^{c}\cap E^{c})+\mathcal{L}_{K}(\Omega^{c}\cap E,\Omega\cap E^{c}),\end{split}$
where
$\Omega_{\sharp}:=(\Omega\times\Omega)\cup(\Omega^{c}\times\Omega)\cup(\Omega\times\Omega^{c}),$
with $\Omega^{c}:=\mathbb{R}^{n}\setminus\Omega$.
We will also use the notation
$\operatorname{P}_{K}(E):=\mathcal{L}_{K}(E,E^{c}).$
Moreover, we define the energy functional $\mathcal{J}$ in $\Omega$ as the perturbation of $\operatorname{P}_{K}$ with the periodic external forcing term $g$, given by
$\mathcal{J}(E,\Omega):=\operatorname{P}_{K}(E,\Omega)+\int_{E\cap\Omega}g(x)\,dx.$
However, taking into account the periodicity of $g$, in order to account for the energy contributions of $\partial\Omega$, we will also consider the energy functional $\mathscr{F}$ defined as
(1.7) $\mathscr{F}(E,\Omega):=\operatorname{P}_{K}(E,\Omega)+\int_{E\cap\mathcal{Q}(\Omega)_{1}}g(x)\,dx,$
where
$\mathcal{Q}(\Omega)_{1}:=\{k+Q\mbox{ s.t. }k\in\mathbb{Z}^{n},\ k+Q\subseteq\Omega\}.$
Construction of planelike minimizers
To construct minimizers for $\mathcal{J}$ and $\mathscr{F}$, given $p\in\mathbb{R}^{n}$, we define
$\text{{E}}_{p}(u):=\iint_{Q\times\mathbb{R}^{n}}\frac{1}{2}|u(x)-u(y)+p\cdot(x-y)|\,K(x,y)\,dx\,dy+\int_{Q}g(x)\,u(x)\,dx.$
Then, we will address the cell problem:
(1.8) $\begin{split}&\mbox{find}\quad u_{p}\in\mathcal{W}:=\left\{u\in L^{1}_{\rm{loc}}(\mathbb{R}^{n})\mbox{ s.t. $u$ is $\mathbb{Z}^{n}$-periodic and }\int_{Q}u(x)\,dx=0\right\}\\ &\mbox{such that}\quad\text{{E}}_{p}(u_{p})=\min_{\mathcal{W}}\text{{E}}_{p}.\end{split}$
Remark 1.3.
We stress that the choice of the domain of integration in the definition of $\text{{E}}_{p}$ is tailored so that the energy functional under consideration is additive with respect to unions of disjoint sets (compare $Q\times\mathbb{R}^{n}$ with $Q_{\sharp}$ and $Q\times Q$).
Moreover, this choice of $\text{{E}}_{p}$ is the right one for deriving the desired Euler–Lagrange equation (see Corollary 3.2) and for constructing (class-$A$) minimizers of $\mathcal{J}$.
The problem in (1.8) is well-defined, according to the following result:
Theorem 1.4 (Existence of minimizers for $\text{{E}}_{p}$).
Let $K:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ satisfy (1.2), (1.3), (1.4), and (1.5).
Then, there exists $\gamma>0$, depending only on $n$, $\kappa_{1}$, $\delta$, and $s_{1}$, such that, if $\|g\|_{L^{\frac{n}{2s_{1}}}(Q)}\leqslant\gamma$, there exists $u\in\mathcal{W}$ such that
(1.9) $\text{{E}}_{p}(u)=\min_{\mathcal{W}}\text{{E}}_{p}$
(1.10) and $u\in L^{\frac{n}{n-2s_{1}}}(Q).$
A key step to find minimizers for $\mathcal{J}$ is the following integral condition for minimizers $u_{p}$ constructed in Theorem 1.4.
Proposition 1.5.
Let $K:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ satisfy (1.2), (1.3), (1.4), and (1.5) and let $u_{p}$ be the minimizer for $\text{{E}}_{p}$ in $\mathcal{W}$ given by Theorem 1.4.
Then, there exists $z:\mathbb{R}^{n}\times\mathbb{R}^{n}\to[-1,1]$ such that, for all $\eta\in C^{\infty}(\mathbb{R}^{n})$ that are $\mathbb{Z}^{n}$-periodic, we have that
(1.11) $\iint_{Q\times\mathbb{R}^{n}}\frac{1}{2}z(x,y)(\eta(x)-\eta(y))\,K(x,y)\,dx\,dy+\int_{Q}g(x)\,\eta(x)\,dx=0.$
Moreover, for a.e. $(x,y)\in\mathbb{R}^{n}\times\mathbb{R}^{n}$ and for all $k\in\mathbb{Z}^{n}$,
(1.12) $u_{p}(x)-u_{p}(y)+p\cdot(x-y)=z(x,y)\big|u_{p}(x)-u_{p}(y)+p\cdot(x-y)\big|,$
(1.13) $z(x,y)=-z(y,x),$
(1.14) and $z(x+k,y+k)=z(x,y).$
A proof of Theorem 1.4 will be presented in Section 2, while Section 3 is devoted to showing Proposition 1.5.
We utilize the function $z$ constructed in Proposition 1.5 as a “calibration” to show that the level sets of the minimizers provided by Theorem 1.4 are in turn (class-$A$) minimizers of our geometric problem.
As a counterpart of the classical theory of minimal surfaces, and in the wake of [MR4117514, MR4142859], one can define calibrations in our setting as follows:
Definition 1.6.
Let $E\subseteq\mathbb{R}^{n}$ be a set of finite $K$-perimeter. We say that a function $z:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ is a calibration for $E$ in $\Omega$ if
$\operatorname{P}_{K}(E,\Omega)=\iint_{\Omega_{\sharp}}\frac{1}{2}z(x,y)\left(\chi_{E}(x)-\chi_{E}(y)\right)K(x,y)\,dx\,dy.$
By analogy with the notation introduced in [MR1852978], we also give the following definition of minimality for $\mathcal{J}$.
Definition 1.7.
We say that a set $E\subseteq\mathbb{R}^{n}$ is a class-$A$ minimizer for $\mathcal{J}$ if, for any domain $\Omega\subseteq\mathbb{R}^{n}$, we have that
$\mathcal{J}(E,\Omega)\leqslant\mathcal{J}(F,\Omega),\quad\mbox{ for any\penalty 10000\ $F$ such that }F\setminus\Omega=E\setminus\Omega.$
Remark 1.8.
Observe that if $E$ is a class-$A$ minimizer for $\mathcal{J}$, then, for any Lipschitz domain $\Omega$, it holds that
$\mathscr{F}(E,\Omega)\leqslant\mathscr{F}(F,\Omega),$
for every $F\subseteq\mathbb{R}^{n}$ such that $F\setminus\mathcal{Q}(\Omega)_{1}=E\setminus\mathcal{Q}(\Omega)_{1}$.
Indeed, by the properties of $\operatorname{P}_{K}$, for any $A$ and $B$ subsets of $\mathbb{R}^{n}$ such that $A\Delta B\subset\mathcal{Q}(\Omega)_{1}$, we have that
$\operatorname{P}_{K}(A,\Omega)-P_{K}(B,\Omega)=\operatorname{P}_{K}(A,\mathcal{Q}(\Omega)_{1})-\operatorname{P}_{K}(B,\mathcal{Q}(\Omega)_{1}).$
Alternatively, observe that if $F\setminus\mathcal{Q}(\Omega)_{1}=E\setminus\mathcal{Q}(\Omega)_{1}$, then in particular $F\setminus\Omega=E\setminus\Omega$ and
$\int_{E\cap\Omega}g(x)\,dx-\int_{F\cap\Omega}g(x)\,dx=\int_{E\cap\mathcal{Q}(\Omega)_{1}}g(x)\,dx-\int_{F\cap\mathcal{Q}(\Omega)_{1}}g(x)\,dx.$
For the rest of our discussion, we adopt the short notation
(1.15) $v_{p}(x):=u_{p}(x)+p\cdot x$
and define, for any $t\in\mathbb{R}$,
(1.16) $E_{p,t}:=\big\{{\mbox{$x\in\mathbb{R}^{n}$ s.t. $v_{p}(x)>t$}}\big\}.$
Then, we have the following result:
Theorem 1.9 (Minimality of level sets).
Let $K:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ satisfy (1.2), (1.3), (1.4), and (1.5) and let $u_{p}$ be a minimizer for $\text{{E}}_{p}$ in $\mathcal{W}$. Let also $v_{p}$ and $E_{p,t}$ be as in (1.15) and (1.16), respectively.
Then, for every $t\in\mathbb{R}$, the set $E_{p,t}$ is a class-$A$ minimizer for $\mathcal{J}$ in the sense of Definition 1.7.
A proof of Theorem 1.9 can be found in Section 4. The idea is to first prove a weaker version of Theorem 1.9, in which the conclusion holds for a.e. $t\in\mathbb{R}$ only. Then, exploiting the closedness of class-$A$ minimizers with respect to the $L^{1}_{\rm{loc}}$-convergence, we will obtain the desired result.
Thanks to suitable density estimates (see Proposition 5.1 below), we also infer that a minimizer $u_{p}$ of $\text{{E}}_{p}$ has controlled oscillations in $Q$. To this end, we recall that, for any $\varphi\in\mbox{BV}(Q)$,
$\operatorname*{osc}_{Q}(\varphi):=\operatorname*{ess\,sup}_{Q}\varphi-\operatorname*{ess\,inf}_{Q}\varphi.$
Theorem 1.10 (Minimizers of $\text{{E}}_{p}$ have controlled oscillations).
Let $K:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ satisfy (1.2), (1.3), (1.4), and (1.5) and let $u_{p}$ be a minimizer for $\text{{E}}_{p}$ in $\mathcal{W}$.
Let also $v_{p}$ be as (1.15).
Then, there exists a constant $c>0$ independent of $p$ such that
$\operatorname*{osc}_{Q}(v_{p})\leqslant c|p|\qquad{\mbox{and}}\qquad\operatorname*{osc}_{Q}(u_{p})\leqslant(c+\sqrt{n})|p|.$
As a consequence of Theorem 1.10, we obtain that the collection of level sets $\{E_{p,t}\}_{t\in\mathcal{T}_{p}}$, where
$\mathcal{T}_{p}:=\{t\in\mathbb{R}\mbox{ s.t. }\partial E_{p,t}\cap Q\neq\varnothing\},$
can be trapped within a strip of height $M$, for some $M>0$ independent of $t$ and $p$. For this reason, we call such level sets “planelike”, according to the setting below.
Definition 1.11.
Let $p\in\mathbb{R}\setminus\{0\}$ and let $E\subseteq\mathbb{R}^{n}$ be a class-$A$ minimizer for the energy functional $\mathcal{J}$.
We say that $E$ is a planelike minimizer for $\mathcal{J}$ in direction $p/|p|$ if there exists $M>0$ (independent of $p$) such that
either $\{x\in\mathbb{R}^{n}\mbox{ s.t. }x\cdot p\leqslant-M|p|\}\subseteq E\subseteq\{x\in\mathbb{R}^{n}\mbox{ s.t. }x\cdot p\leqslant M|p|\}$
or $\{x\in\mathbb{R}^{n}\mbox{ s.t. }x\cdot p\geqslant M|p|\}\subseteq E\subseteq\{x\in\mathbb{R}^{n}\mbox{ s.t. }x\cdot p\geqslant-M|p|\}.$
In both cases, it also holds
$\partial E\subseteq\{x\in\mathbb{R}^{n}\mbox{ s.t. }|x\cdot p|\leqslant M|p|\}.$
Corollary 1.12 (Level sets are planelike minimizers).
Let $p\in\mathbb{R}^{n}$ and let $K:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ satisfy (1.2), (1.3), (1.4), and (1.5). Let $u_{p}$ be a minimizer for $\text{{E}}_{p}$ in $\mathcal{W}$ and let also $v_{p}$ be as in (1.15), $t\in\mathbb{R}$ and $E_{p,t}$ be defined as in (1.16).
If $t\in\mathcal{T}_{p}$, then $E_{p,t}$ is a planelike minimizer in direction $p/|p|$ in the sense of Definition 1.11. Namely, there exists $M>0$ independent of $p$ and $t$ such that
(1.17) $E_{p,t}\subseteq\{x\in\mathbb{R}^{n}\mbox{ s.t. }x\cdot p\geqslant-M|p|\}$
(1.18) and $\{x\in\mathbb{R}^{n}\mbox{ s.t. }x\cdot p\geqslant M|p|\}\subseteq E_{p,t}.$
In particular, it holds
$\partial E_{p,t}\subseteq\{x\in\mathbb{R}^{n}\mbox{ s.t. }|x\cdot p|\leqslant M|p|\}.$
Section 5 is committed to showing the density estimates needed for Theorem 1.10, while a proof of the latter and Corollary 1.12 is contained in Section 6.
$\Gamma$-convergence results and the stable norm
With the planelike minimizers of Corollary 1.12, we are able to prove the main result of this paper, which is the $\Gamma$-convergence of a sequence of rescaled energies to the so-called stable norm.
Let $\Omega\subseteq\mathbb{R}^{n}$ be a Lipschitz domain and let, for any $\varepsilon>0$,
$\mathcal{Q}(\Omega)_{\varepsilon}:=\{\varepsilon(k+Q)\mbox{ s.t. }k\in\mathbb{Z}^{n},\ \varepsilon(k+Q)\subseteq\Omega\}.$
Then, for any set $E$, we define the rescaled energy $\mathscr{F}_{\varepsilon}$ in $\Omega$ as
(1.19) $\begin{split}\mathscr{F}_{\varepsilon}(E,\Omega)&:=\iint_{\Omega_{\sharp}}\frac{1}{2}|\chi_{E}(x)-\chi_{E}(y)|K_{\varepsilon}(x,y)\,dx\,dy+\int_{\mathcal{Q}(\Omega)_{\varepsilon}\cap E}g_{\varepsilon}(x)\,dx\\ &=\operatorname{P}_{K_{\varepsilon}}(E,\Omega)+\int_{\mathcal{Q}(\Omega)_{\varepsilon}\cap E}g_{\varepsilon}(x)\,dx,\end{split}$
where
$K_{\varepsilon}(x,y):=\frac{1}{\varepsilon^{n+1}}K\left(\frac{x}{\varepsilon},\frac{y}{\varepsilon}\right)\qquad{\mbox{and}}\qquad g_{\varepsilon}(x):=\frac{1}{\varepsilon}g\left(\frac{x}{\varepsilon}\right).$
For later convenience, we also set
(1.20) $\begin{split}\mathscr{E}_{\varepsilon}(E,\Omega)&:=\mathscr{F}_{\varepsilon}(E,\Omega)-\mathcal{L}_{K_{\varepsilon}}(E\cap\Omega,E^{c}\cap\Omega^{c})\\ &=\iint_{\Omega\times\mathbb{R}^{n}}\chi_{E}(x)\chi_{E^{c}}(y)K_{\varepsilon}(x,y)\,dx\,dy+\int_{\mathcal{Q}(\Omega)_{\varepsilon}\cap E}g_{\varepsilon}(x)\,dx.\end{split}$
We will show that the sequence $\{\mathscr{F}_{\varepsilon}\}_{\varepsilon>0}$ $\Gamma$-converges to a local anisotropic perimeter.
To this end, let $p\in\mathbb{S}^{n-1}$ and let $Q^{p}$ be a cube of side $1$ with a face perpendicular to $p$. Moreover, let $E_{p}$ be a planelike minimizer for $\mathcal{J}$. Then, in the wake of [MR3223561], we define the stable norm $\phi:\mathbb{S}^{n-1}\to\mathbb{R}$ as
(1.21) $\begin{split}\phi(p):=\lim_{\varepsilon\to 0^{+}}\mathscr{F}_{\varepsilon}(\varepsilon E_{p},Q^{p})=\lim_{R\to+\infty}R^{1-n}\mathscr{F}_{1}(E_{p},Q^{p}_{R}),\end{split}$
whenever such limit exists. Here, we are adopting the notation $Q^{p}_{R}$ for the cube of side $R$ with a face perpendicular to $p$.
Besides, for any set $E\subseteq\Omega$ of finite perimeter, we define
(1.22) $\mathscr{F}_{\phi}(E,\Omega):=\int_{\partial E\cap\Omega}\phi(\nu_{E}(x))d\mathcal{H}^{n-1}(x).$
Then, the following holds true:
Theorem 1.13 ($\Gamma$-convergence of the energies $\mathscr{F}_{\varepsilon}$).
We have that $\mathscr{F}_{\varepsilon}\xrightarrow{\Gamma}\mathscr{F}_{\phi}$, as $\varepsilon\to 0$, with respect to the $L^{1}_{\rm{loc}}(\mathbb{R}^{n})$-convergence of sets. Namely, for every set $E\subseteq\mathbb{R}^{n}$, we have that
  1. (i)
    for any sequence of sets $\{E_{\varepsilon}\}_{\varepsilon}$ such that $E_{\varepsilon}\to E$ in $L^{1}_{\rm{loc}}(\mathbb{R}^{n})$ (i.e. $|E_{\varepsilon}\Delta E|\to 0$ as $\varepsilon\to 0$),
    (1.23) $\liminf_{\varepsilon\to 0}\mathscr{F}_{\varepsilon}(E_{\varepsilon},\Omega)\geqslant\mathscr{F}_{\phi}(E,\Omega);$
  2. (ii)
    there exists a sequence of sets $\{E_{\varepsilon}\}_{\varepsilon}$ such that $E_{\varepsilon}\to E$ in $L^{1}_{\rm{loc}}(\mathbb{R}^{n})$ and
    (1.24) $\limsup_{\varepsilon\to 0}\mathscr{F}_{\varepsilon}(E_{\varepsilon},\Omega)\leqslant\mathscr{F}_{\phi}(E,\Omega).$
Some preliminary results needed for Theorem 1.13 can be found in Section 7, while we will present some properties of $\phi$ such as well-definedness in Section 8 and continuity in Section 10. Then, Sections 9 and 11 are devoted to proving Theorem 1.13-(i) and Theorem 1.13-(ii) respectively.
We stress that Proposition 7.3 in Section 7 entails that the energy $\mathscr{F}_{1}$ is non-negative in every large domain. However, there is no Poincaré Inequality available for generic non-local kernels as the ones considered by our [^2] Check [MR3732175, Proposition 4.1] for a Poincaré Inequality for a non-local kernel $K$ under some additional assumptions. setting. Thus, given a sequence of sets $\{E_{\varepsilon}\}_{\varepsilon}$ with $\sup_{\varepsilon}\mathscr{F}_{\varepsilon}(E_{\varepsilon},\Omega)<+\infty$, to the best of our knowledge, Proposition 7.3 does not provide any useful uniform bound either on the $BV$-norm $\|\chi_{E_{\varepsilon}}\|_{BV(\Omega)}$, or on any Gagliardo seminorm $\chi_{E_{\varepsilon}}]_{W^{s,p}(\Omega)}$ (compare with [chambolle_thouroude, Remark 2.2]). For this reason, it remains an open question whether a compactness result holds for sequences of sets with uniformly bounded energy.
Complementary results
Additionally, we state and prove in the appendices of this work a few interesting results concerning some uniform density estimates for almost minimizers (defined in a suitable sense, to be compared with [sequoia, Definition 1.1]) of $\operatorname{P}_{K}$ and the behavior of our non-local perimeter in large balls (i.e. in $B_{R}$, with $R\to+\infty$).
Comparison with the existing literature
In relation with the existing literature for the non-local setting, we recall that problems entangled with phase-transition, both at the macro-scale and at the meso-scale, have been addressed in [MR1612250, MR1634336] and [pagliardini_non_local_planelike], respectively. However, our setting is quite different, since in [MR1612250, MR1634336] the authors only consider a model with no external fields, while in [pagliardini_non_local_planelike] no $\Gamma$-convergence results have been investigated. Moreover, we drop the presence of the double-well potential, that is peculiar of the Allen-Cahn model.
In our setting, we have to face two main difficulties. First, we observe that the forcing term $g$ requires a careful treatment. Indeed, even though the energy contribution of $g$ does not need to be positive, we still have to guarantee that the energies $\mathscr{F}_{\varepsilon}$ eventually are. Moreover, we want to avoid the contribution due to $g$ near the boundary of the domain.
The second main concern is that, since the kernel $K$ accounts for long-range interactions, we will need specific assumptions to ensure that $K$ “localizes in the limit”. In fact, on the one hand, to have a well-defined stable norm $\phi$, we need that the “volume term” measured by $K$ and the contribution due to $g$ scale compatibly. On the other hand, we want to avoid an infinite energy contribution coming from outside the domain.
Further developments
A possible further research direction could be investigating the existence of class-$A$ minimizers for $\mathcal{J}$ under a mass constraint. Suppose that such minimizers exist and let $\{E_{m}\}_{m>0}$ be a sequence of class-$A$ minimizers for $\mathcal{J}$ such that $|E_{m}|=m$. Then, in the fashion of the classical case, it is reasonable to expect that a suitable rescaling of $E_{m}$ converges, as $m\to+\infty$, to the Wulff shape (i.e. the isoperimetric set) for $\operatorname{P}_{\phi}$.
As another further development of this work, we point out that it would be interesting to investigate the differentiability of the stable norm. We refer to [MR2197072] for the analogous in the classical case on manifolds. See also [MR3223561] for a different approach using the local version of the cell problem (1.8).
Finally, it would be interesting to determine whether it is possible to prove the existence of planelike minimizers when the kernel $K$ is integrable (dropping the lower bound in (1.4)). For instance, in [MR4360596], the authors have proved that a suitable rescaling of the fractional perimeter $\mbox{Per}_{s}$ $\Gamma$-converges, as $s\to 0$, to an integral operator with an $L^{1}$ kernel, now commonly referred to as 0-Perimeter. However, addressing this problem goes beyond the purpose of this paper.
2. Existence of minimizers for $\text{{E}}_{p}$ - Proof of Theorem 1.4
The first step to construct planelike minimizers for the energy functional $\mathcal{J}$ is to prove that the cell problem (1.8) is well-defined, as stated in Theorem 1.4.
To do this, we observe that
$\begin{split}&\text{{E}}_{p}(0)=\iint_{Q\times\mathbb{R}^{n}}\frac{1}{2}|p\cdot(x-y)|\,K(x,y)\,dx\,dy\leqslant\frac{|p|}{2}\iint_{Q\times\mathbb{R}^{n}}|x-y|\,K(x,y)\,dx\,dy\\ &\qquad\qquad=\frac{|p|\,|Q|}{2}\int_{\mathbb{R}^{n}}|h|K(h,0)\,dh<+\infty,\end{split}$
thanks to (1.2) and (1.3). Hence, the minimization process takes place on a nonempty set of competitors.
Accordingly, we take a minimizing sequence $u_{j}\in\mathcal{W}$ with $\text{{E}}_{p}(u_{j})\leqslant\text{{E}}_{p}(0)$. Thus, using (1.4), the fact that $g\in L^{\infty}(\mathbb{R}^{n})\subseteq L^{\frac{n}{2s_{1}}}(Q)$, and the Hölder inequality, we obtain
(2.1) $\begin{split}\text{{E}}_{p}(0)&\geqslant\iint_{Q\times\mathbb{R}^{n}}\frac{1}{2}|u_{j}(x)-u_{j}(y)|\,K(x,y)\,dx\,dy-\iint_{Q\times\mathbb{R}^{n}}\frac{1}{2}|p\cdot(x-y)|\,K(x,y)\,dx\,dy\\ &\quad-\int_{Q}|g(x)|\,|u_{j}(x)|\,dx\\ &\geqslant\frac{\kappa_{1}}{2}\int_{(Q\times\mathbb{R}^{n})\cap\{|x-y|<\delta\}}\frac{|u_{j}(x)-u_{j}(y)|}{|x-y|^{n+2s_{1}}}\,dx\,dy-\text{{E}}_{p}(0)-\|g\|_{L^{\frac{n}{2s_{1}}}(Q)}\|u_{j}\|_{L^{\frac{n}{n-2s_{1}}}(Q)}.\end{split}$
Also, by the fractional Poincaré inequality (see e.g. [MR4567945, Theorem 6.33]),
(2.2) $\|u_{j}\|_{L^{1}(Q)}\leqslant C\iint_{Q\times Q}\frac{|u_{j}(x)-u_{j}(y)|}{|x-y|^{n+2s_{1}}}\,dx\,dy$
and, thanks to the fractional Sobolev inequality (see e.g. [MR2944369, Theorem 6.7]),
(2.3) $\|u_{j}\|_{L^{\frac{n}{n-2s_{1}}}(Q)}\leqslant C\left(\|u_{j}\|_{L^{1}(Q)}+\iint_{Q\times Q}\frac{|u_{j}(x)-u_{j}(y)|}{|x-y|^{n+2s_{1}}}\,dx\,dy\right),$
for some constant $C>0$ depending only on $n$ and $s_{1}$ and possibly changing from line to line.
Combining (2.2) and (2.3), we infer that
$\|u_{j}\|_{L^{\frac{n}{n-2s_{1}}}(Q)}\leqslant C\iint_{Q\times Q}\frac{|u_{j}(x)-u_{j}(y)|}{|x-y|^{n+2s_{1}}}\,dx\,dy.$
This, together with Lemma A.1, gives that
$\begin{split}\|u_{j}\|_{L^{\frac{n}{n-2s_{1}}}(Q)}&\leqslant C\iint_{(Q\times Q)\cap\{|x-y|<\delta\}}\frac{|u_{j}(x)-u_{j}(y)|}{|x-y|^{n+2s_{1}}}\,dx\,dy\\ &\leqslant C\int_{(Q\times\mathbb{R}^{n})\cap\{|x-y|<\delta\}}\frac{|u_{j}(x)-u_{j}(y)|}{|x-y|^{n+2s_{1}}}\,dx\,dy,\end{split}$
up to renaming $C>0$, in dependence only of $\delta$, $n$, and $s_{1}$.
Therefore, plugging the latter inequality into (2.1), we deduce that
(2.4) $\begin{split}&\frac{\kappa_{1}}{4}\int_{(Q\times\mathbb{R}^{n})\cap\{|x-y|<\delta\}}\frac{|u_{j}(x)-u_{j}(y)|}{|x-y|^{n+2s_{1}}}\,dx\,dy\\ &\qquad\leqslant\left(\frac{\kappa_{1}}{2}-C\|g\|_{L^{\frac{n}{2s_{1}}}(Q)}\right)\int_{(Q\times\mathbb{R}^{n})\cap\{|x-y|<\delta\}}\frac{|u_{j}(x)-u_{j}(y)|}{|x-y|^{n+2s_{1}}}\,dx\,dy\leqslant 2\text{{E}}_{p}(0),\end{split}$
as long as $\|g\|_{L^{\frac{n}{2s_{1}}}(Q)}\leqslant\frac{\kappa_{1}}{4C}$.
Lemma A.1, (2.2), and (2.4) provide a uniform bound for $u_{j}$ in $W^{s_{1},1}(Q)$. Accordingly, by (2.3), the sequence $\{u_{j}\}_{j}$ is also bounded in $L^{\frac{n}{n-2s_{1}}}(Q)$, and hence precompact in $L^{1}(Q)$ (see e.g. [MR2944369, Theorem 7.1]). It thereby follows that, up to a subsequence, the sequence $\{u_{j}\}_{j}$ converges to some $u$ in $L^{1}(Q)$ and, by periodicity, a.e. in $\mathbb{R}^{n}$. In particular, we have that $u\in\mathcal{W}$.
Now, by construction, and thanks to the Fatou Lemma, we have that
(2.5) $\begin{split}&\inf_{\mathcal{W}}\text{{E}}_{p}=\lim_{j\to+\infty}\text{{E}}_{p}(u_{j})\\ &\qquad\geqslant\iint_{Q\times\mathbb{R}^{n}}\frac{1}{2}|u(x)-u(y)+p\cdot(x-y)|\,K(x,y)\,dx\,dy+\lim_{j\to+\infty}\int_{Q}g(x)\,u_{j}(x)\,dx.\end{split}$
Notice also that, again by the Fatou Lemma,
(2.6) $S:=\sup_{j\in\mathbb{N}}\|u_{j}\|_{L^{\frac{n}{n-2s_{1}}}(Q)}\geqslant\lim_{j\to+\infty}\|u_{j}\|_{L^{\frac{n}{n-2s_{1}}}(Q)}\geqslant\|u\|_{L^{\frac{n}{n-2s_{1}}}(Q)}.$
Moreover, given $\varepsilon>0$, using the absolute continuity of the Lebesgue integral (see e.g. [MR1681462, Corollary 3.6]), we find $\eta>0$ such that, whenever a subset $V$ of $Q$ has measure less than $\eta$, we have that
(2.7) $\|g\|_{L^{\frac{n}{2s_{1}}}(V)}\leqslant\varepsilon.$
At this stage, we invoke the Severini-Egorov Theorem (see e.g. [MR1681462, Theorem 2.33]) to find a set $U\subseteq Q$ of measure less than $\eta$ such that $u_{j}$ converges to $u$ uniformly in $Q\setminus U$. As a result, we come to
$\begin{split}&\lim_{j\to+\infty}\left|\int_{Q}g(x)\,\big(u_{j}(x)-u(x)\big)\,dx\right|=\lim_{j\to+\infty}\left|\int_{U}g(x)\,\big(u_{j}(x)-u(x)\big)\,dx\right|\\ &\qquad\qquad\leqslant\lim_{j\to+\infty}\left(\|u_{j}\|_{L^{\frac{n}{n-2s_{1}}}(U)}+\|u\|_{L^{\frac{n}{n-2s_{1}}}(U)}\right)\|g\|_{L^{\frac{n}{2s_{1}}}(U)}\leqslant 2S\|g\|_{L^{\frac{n}{2s_{1}}}(U)}.\end{split}$
From this and (2.7), we infer that
$\lim_{j\to+\infty}\left|\int_{Q}g(x)\,\big(u_{j}(x)-u(x)\big)\,dx\right|\leqslant\varepsilon.$
By the arbitrariness of $\varepsilon$, we find that
$\lim_{j\to+\infty}\int_{Q}g(x)\,\big(u_{j}(x)-u(x)\big)\,dx=0.$
Plugging this information into (2.5), we obtain that
$\inf_{\mathcal{W}}\text{{E}}_{p}\geqslant\iint_{Q\times\mathbb{R}^{n}}|u(x)-u(y)+p\cdot(x-y)|\,K(x,y)\,dx\,dy+\int_{Q}g(x)\,u(x)\,dx=\text{{E}}_{p}(u),$
showing that $u$ is a minimizer (i.e. (1.9)). The statement in (1.10) is a consequence of (2.6).
Remark 2.1.
We point out that, by inspection of the proofs of Theorem 1.4 and Lemma A.1, we infer that
$\|g\|_{L^{\frac{n}{2s_{1}}}(Q)}\leqslant\frac{\kappa_{1}}{4C}\leqslant\widetilde{c}\,\delta^{n+2s_{1}}\varphi(\delta),$
where $\kappa_{1}$, $s_{1}$, and $\delta$ are as in (1.4), $C$ is a multiple of the constant appearing in Lemma A.1, $\widetilde{c}=\widetilde{c}(n,s_{1},\kappa_{1})$ is a positive constant, and $\varphi$ is a function such that
$\lim_{\delta\to 0^{+}}\varphi(\delta)=0.$
3. Proof of Proposition 1.5 and Euler-Lagrange equation for the cell problem (1.8)
Here, we present the proof of Proposition 1.5.
Proof of Proposition 1.5.
We first show (1.11) for every $\eta\in C^{\infty}(\mathbb{R}^{n})$ that is $\mathbb{Z}^{n}$-periodic, and such that $\int_{Q}\eta(x)\,dx=0$. Let us consider
$\begin{split}\mathcal{X}:=\Bigg\{u\in L^{\frac{n}{n-2s_{1}}}(Q)\quad&\mbox{s.t. $u$ is $\mathbb{Z}^{n}$-periodic, }\int_{Q}u(x)\,dx=0,\\ &\mbox{and }\iint_{Q\times\mathbb{R}^{n}}\frac{1}{2}|u(x)-u(y)|\,K(x,y)\,dx\,dy<+\infty\Bigg\}.\end{split}$
For any $u\in{\mathcal{X}}$, in the spirit of [MR3491533, Theorem 2.3] (see also [MR3930619, MR3996039, MR4537323, MR4645235]), we define the subdifferential of $\text{{E}}_{p}$ at $u$ as the collection of all $\varphi\in{\mathcal{X}}$ satisfying
(3.1) $\text{{E}}_{p}(w)-\text{{E}}_{p}(u)\geqslant\int_{Q}\big(w(x)-u(x)\big)\,\varphi(x)\,dx,\quad\mbox{ for all\penalty 10000\ $w\in{\mathcal{X}}$,}$
and we denote it by $\partial\text{{E}}_{p}(u)$.
Now, given a minimizer $u_{p}$ for $\text{{E}}_{p}$ (as constructed in Theorem 1.4), let us define a function $z:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ as
(3.2) $z(x,y):=\begin{cases}\displaystyle\frac{u_{p}(x)-u_{p}(y)+p\cdot(x-y)}{|u_{p}(x)-u_{p}(y)+p\cdot(x-y)|},\quad&{\mbox{if }}u_{p}(x)-u_{p}(y)+p\cdot(x-y)\neq 0,\\ \\ 0,\quad&{\mbox{ otherwise.}}\end{cases}$
We have that
(3.3) $\left|u_{p}(x)-u_{p}(y)+p\cdot(x-y)\right|z(x,y)=u_{p}(x)-u_{p}(y)+p\cdot(x-y),$
namely (1.12) holds true.
Also,
$z(x,y)=-z(y,x)\in[-1,1]\quad\mbox{for a.e. }(x,y)\in\mathbb{R}^{n}\times\mathbb{R}^{n},$
and $z(x+k,y+k)=z(x,y)\quad\mbox{for all }k\in\mathbb{Z}^{n},$
that are (1.13) and (1.14), respectively.
Now, we claim that, for all $\varphi\in\partial\text{{E}}_{p}(u_{p})$ and $\eta\in{\mathcal{X}}\cap C^{\infty}(\mathbb{R}^{n})$, the function $z$ satisfies
(3.4) $\iint_{Q\times\mathbb{R}^{n}}\frac{1}{2}z(x,y)(\eta(x)-\eta(y))\,K(x,y)\,dx\,dy+\int_{Q}g(x)\,\eta(x)\,dx=\int_{Q}\varphi(x)\,\eta(x)\,dx.$
To check this, we recall the notation for $v_{p}$ in (1.15) and we stress that, by (3.3),
$|v_{p}(x)-v_{p}(y)|\,z(x,y)=v_{p}(x)-v_{p}(y).$
Now, let $\varepsilon\in(-1,1)\setminus\{0\}$ and $w:=u_{p}+\varepsilon\eta$. Notice that $w\in{\mathcal{X}}$, since both $u_{p}$ and $\eta$ are (recall (1.10) in Theorem 1.4), therefore we can employ (3.1) with $u:=u_{p}$ and $w$ as above and we obtain that
(3.5) $\begin{split}&\varepsilon\int_{Q}\varphi(x)\,\eta(x)\,dx-\varepsilon\int_{Q}g(x)\,\eta(x)\,dx\\ \leqslant\;&\iint_{Q\times\mathbb{R}^{n}}\frac{1}{2}\Big(|v_{p}(x)-v_{p}(y)+\varepsilon(\eta(x)-\eta(y))|-|v_{p}(x)-v_{p}(y)|\Big)\,K(x,y)\,dx\,dy\\ =\;&\iint_{Q\times\mathbb{R}^{n}}\frac{1}{2}\Big(\big||v_{p}(x)-v_{p}(y)|z(x,y)+\varepsilon(\eta(x)-\eta(y))\big|-|v_{p}(x)-v_{p}(y)|\Big)\,K(x,y)\,dx\,dy\\ =\;&\iint_{Q\times\mathbb{R}^{n}}\frac{1}{2}\left(\int_{0}^{1}\frac{d}{dt}\Big||v_{p}(x)-v_{p}(y)|z(x,y)+t\varepsilon(\eta(x)-\eta(y))\Big|\,dt\right)\,K(x,y)\,dx\,dy\\ =\;&\iint_{Q\times\mathbb{R}^{n}}\frac{1}{2}\left(\int_{0}^{1}\varepsilon\frac{\Big(|v_{p}(x)-v_{p}(y)|z(x,y)+t\varepsilon(\eta(x)-\eta(y))\Big)(\eta(x)-\eta(y))}{\Big||v_{p}(x)-v_{p}(y)|z(x,y)+t\varepsilon(\eta(x)-\eta(y))\Big|}\,dt\right)\,K(x,y)\,dx\,dy.\end{split}$
In addition, for all $t\in[0,1]$,
(3.6) $\begin{split}&\left|\frac{\Big(|v_{p}(x)-v_{p}(y)|z(x,y)+t\varepsilon(\eta(x)-\eta(y))\Big)(\eta(x)-\eta(y))}{\Big||v_{p}(x)-v_{p}(y)|z(x,y)+t\varepsilon(\eta(x)-\eta(y))\Big|}\right|\\ &\qquad\qquad\qquad\leqslant|\eta(x)-\eta(y)|\leqslant 2\|\eta\|_{C^{1}(\mathbb{R}^{n})}\min\{|x-y|,1\}.\end{split}$
On this account, dividing by $\varepsilon$ in (3.5) and discussing the sign of $\varepsilon$, thanks to the Dominated Convergence Theorem (whose application is justified by (1.3) and (3.6)), we deduce that
$\begin{split}&\int_{Q}\varphi(x)\,\eta(x)\,dx-\int_{Q}g(x)\,\eta(x)\,dx\\ &\qquad=\lim_{\varepsilon\to 0}\iint_{Q\times\mathbb{R}^{n}}\frac{1}{2}\left(\int_{0}^{1}\frac{\Big(|v_{p}(x)-v_{p}(y)|z(x,y)+t\varepsilon(\eta(x)-\eta(y))\Big)(\eta(x)-\eta(y))}{\Big||v_{p}(x)-v_{p}(y)|z(x,y)+t\varepsilon(\eta(x)-\eta(y))\Big|}\,dt\right)\,K(x,y)\,dx\,dy\\ &\qquad=\iint_{Q\times\mathbb{R}^{n}}\frac{1}{2}\left(\int_{0}^{1}\frac{|v_{p}(x)-v_{p}(y)|z(x,y)(\eta(x)-\eta(y))}{|v_{p}(x)-v_{p}(y)|}\,dt\right)\,K(x,y)\,dx\,dy\\ &\qquad=\iint_{Q\times\mathbb{R}^{n}}\frac{1}{2}z(x,y)(\eta(x)-\eta(y))\,K(x,y)\,dx\,dy,\end{split}$
showing (3.4).
Now, by the minimality of $u_{p}$, for all $w\in\mathcal{X}\subseteq\mathcal{W}$, we have that $\text{{E}}_{p}(w)-\text{{E}}_{p}(u)\geqslant 0$, which entails that $0\in\partial\text{{E}}_{p}(u_{p})$. Hence, employing (3.4) with $\varphi\equiv 0$ yields (1.11) for every $\eta\in C^{\infty}(\mathbb{R}^{n})$ that is $\mathbb{Z}^{n}$-periodic, and such that $\int_{Q}\eta(x)\,dx=0$.
Let now $\eta\in C^{\infty}(\mathbb{R}^{n})$ be $\mathbb{Z}^{n}$-periodic, and define
$\widetilde{\eta}(x):=\eta(x)-\int_{Q}\eta(y)\,dy.$
In particular, $\widetilde{\eta}\in\mathcal{C}^{\infty}(\mathbb{R}^{n})\cap\mathcal{X}$ and (1.11) holds true for $\widetilde{\eta}$.
Therefore, taking advantage of (1.1) and the definition of $\widetilde{\eta}$, we infer that
$\begin{split}&\iint_{Q\times\mathbb{R}^{n}}\frac{1}{2}z(x,y)(\eta(x)-\eta(y))\,K(x,y)\,dx\,dy+\int_{Q}g(x)\,\eta(x)\,dx\\ &\qquad=\iint_{Q\times\mathbb{R}^{n}}\frac{1}{2}z(x,y)(\widetilde{\eta}(x)-\widetilde{\eta}(y))\,K(x,y)\,dx\,dy+\int_{Q}g(x)\,\widetilde{\eta}(x)\,dx=0.\end{split}$
concluding the proof. ∎
By a direct computation, from Proposition 1.5, we also obtain the following:
Corollary 3.1.
Let $K:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ satisfy (1.2), (1.3), (1.4), and (1.5) and let $u_{p}$ be the minimizer for $\text{{E}}_{p}$ in $\mathcal{W}$ given by Theorem 1.4. Let also $z:\mathbb{R}^{n}\times\mathbb{R}^{n}\to[-1,1]$ be as in Proposition 1.5.
Then, for all $\eta\in C^{\infty}(\mathbb{R}^{n})$ that are $\mathbb{Z}^{n}$-periodic, we have that
(3.7) $\begin{split}&\iint_{Q\times Q}\frac{1}{2}z(x,y)(\eta(x)-\eta(y))\,K(x,y)\,dx\,dy\\ &\qquad+\iint_{Q\times Q^{c}}z(x,y)\eta(x)\,K(x,y)\,dx\,dy+\int_{Q}g(x)\,\eta(x)\,dx=0.\end{split}$
Proof.
Let $\eta\in C^{\infty}(\mathbb{R}^{n})$ be $\mathbb{Z}^{n}$-periodic. Observe that (3.7) will follow from (1.11) if we show that
$\begin{split}&\iint_{Q\times\mathbb{R}^{n}}\frac{1}{2}z(x,y)(\eta(x)-\eta(y))\,K(x,y)\,dx\,dy\\ &\qquad\qquad=\iint_{Q\times Q}\frac{1}{2}z(x,y)(\eta(x)-\eta(y))\,K(x,y)\,dx\,dy+\iint_{Q\times Q^{c}}z(x,y)\eta(x)\,K(x,y)\,dx\,dy,\end{split}$
that is, if we show that
(3.8) $\iint_{Q\times Q^{c}}z(x,y)\eta(y)\,K(x,y)\,dx\,dy=-\iint_{Q\times Q^{c}}z(x,y)\eta(x)\,K(x,y)\,dx\,dy.$
To achieve this, notice that $Q^{c}=\bigcup_{j\in\mathbb{Z}^{n}\setminus\{0\}}j+Q$. Moreover, according to [MR3732175, Remark 2.1], we have that
$\iint_{Q\times Q^{c}}\left|z(x,y)\eta(y)\,K(x,y)\right|\,dx\,dy\leqslant\|\eta\|_{L^{\infty}(Q)}\operatorname{P}_{K}(Q)<+\infty$
Thus, using Fubini-Tonelli’s Theorem together with the periodicity of $\eta$, $z$, and $K$ (see (1.14) and (1.2)), we obtain
$\begin{split}&\iint_{Q\times Q^{c}}z(x,y)\eta(y)\,K(x,y)\,dx\,dy=\sum_{j\in\mathbb{Z}^{n}\setminus\{0\}}\iint_{Q\times(j+Q)}z(x,y)\eta(y)\,K(x,y)\,dx\,dy\\ &\qquad\qquad=\sum_{j\in\mathbb{Z}^{n}\setminus\{0\}}\iint_{(Q-j)\times Q}z(x,y)\eta(y)\,K(x,y)\,dx\,dy=\iint_{Q^{c}\times Q}z(x,y)\eta(y)\,K(x,y)\,dx\,dy\\ &\qquad\qquad=\iint_{Q\times Q^{c}}z(y,x)\eta(x)\,K(y,x)\,dx\,dy=-\iint_{Q\times Q^{c}}z(x,y)\eta(x)\,K(x,y)\,dx\,dy,\end{split}$
where we have also used (1.13) and the symmetry of $K$ to obtain the last equality. This shows (3.8), concluding the proof. ∎
As a byproduct of Corollary 3.1, we deduce the Euler-Lagrange equation for the cell problem (1.8).
Corollary 3.2 (Euler-Lagrange equation for the cell problem (1.8)).
Let $K:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ satisfy (1.2), (1.3), (1.4), and (1.5). Let also $z:\mathbb{R}^{n}\times\mathbb{R}^{n}\to[-1,1]$ be as in Proposition 1.5.
Then, for every $\eta\in C_{c}^{\infty}(\mathbb{R}^{n})$, we have that
(3.9) $\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\frac{1}{2}z(x,y)(\eta(x)-\eta(y))\,K(x,y)\,dx\,dy+\int_{\mathbb{R}^{n}}g(x)\,\eta(x)\,dx=0.$
Proof.
We use the notation $Q^{(j)}:=k_{j}+Q$, for some $k_{j}\in\mathbb{Z}^{n}$ and we first claim that
(3.10) formula (3.9) holds true for $\eta\in C^{\infty}_{c}(Q^{(j)})$.
To do this, we extend $\eta$ periodically in $\mathbb{R}^{n}$ and we call this new function $\widetilde{\eta}$.
Taking advantage of the periodicity of $\widetilde{\eta}$, $z$, and $K$, we obtain
(3.11) $\begin{split}&\iint_{Q^{(j)}\times Q^{(j)}}\frac{1}{2}z(x,y)(\eta(x)-\eta(y))\,K(x,y)\,dx\,dy+\iint_{Q^{(j)}\times(Q^{(j)})^{c}}z(x,y)\eta(x)\,K(x,y)\,dx\,dy\\ =&\iint_{Q^{(j)}\times Q^{(j)}}\frac{1}{2}z(x,y)(\widetilde{\eta}(x)-\widetilde{\eta}(y))\,K(x,y)\,dx\,dy+\iint_{Q^{(j)}\times(Q^{(j)})^{c}}z(x,y)\widetilde{\eta}(x)\,K(x,y)\,dx\,dy\\ =&\iint_{Q\times Q}\frac{1}{2}z(x,y)(\widetilde{\eta}(x)-\widetilde{\eta}(y))\,K(x,y)\,dx\,dy+\iint_{Q\times Q^{c}}z(x,y)\widetilde{\eta}(x)\,K(x,y)\,dx\,dy.\end{split}$
Besides, from the periodicity of $\widetilde{\eta}$ and $g$, we deduce that
(3.12) $\int_{Q^{(j)}}g(x)\,\eta(x)\,dx=\int_{Q^{(j)}}g(x)\,\widetilde{\eta}(x)\,dx=\int_{Q}g(x)\,\widetilde{\eta}(x)\,dx.$
Therefore, using (3.7) with $\eta:=\widetilde{\eta}$, (3.11), and (3.12), we infer that
(3.13) $\begin{split}&\iint_{Q^{(j)}\times Q^{(j)}}\frac{1}{2}z(x,y)(\eta(x)-\eta(y))\,K(x,y)\,dx\,dy\\ &\qquad+\iint_{Q^{(j)}\times\left(Q^{(j)}\right)^{c}}z(x,y)\eta(x)\,K(x,y)\,dx\,dy+\int_{Q^{(j)}}g(x)\,\eta(x)\,dx=0.\end{split}$
Now, recall that
$(Q^{(j)})_{\sharp}=(\mathbb{R}^{n}\times\mathbb{R}^{n})\setminus\big((Q^{(j)})^{c}\times(Q^{(j)})^{c}\big).$
Thus, since $\mbox{supp}(\eta)\Subset Q^{(j)}$, (3.13) entails that
$\begin{split}&\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\frac{1}{2}z(x,y)\left(\eta(x)-\eta(y)\right)\,K(x,y)\,dx\,dy+\int_{\mathbb{R}^{n}}g(x)\,\eta(x)\,dx\\ &\qquad=\iint_{(Q^{(j)})_{\sharp}}\frac{1}{2}z(x,y)\left(\eta(x)-\eta(y)\right)\,K(x,y)\,dx\,dy+\int_{Q^{(j)}}g(x)\,\eta(x)\,dx\\ &\qquad=\iint_{Q^{(j)}\times Q^{(j)}}\frac{1}{2}z(x,y)\left(\eta(x)-\eta(y)\right)\,K(x,y)\,dx\,dy\\ &\qquad\qquad+\iint_{Q^{(j)}\times(Q^{(j)})^{c}}z(x,y)\eta(x)\,K(x,y)\,dx\,dy+\int_{Q^{(j)}}g(x)\,\eta(x)\,dx\\ &\qquad=0,\end{split}$
where we have also used (1.13) to obtain the second equality. This establishes the claim in (3.10).
Let now $\eta\in C^{\infty}_{c}(\mathbb{R}^{n})$ with $\mbox{supp}(\eta)\Subset\Omega$, for some domain $\Omega\subseteq\mathbb{R}^{n}$, and consider a finite covering of $\Omega$ made of unit cubes $Q^{(1)},\dots,Q^{(N)}$, with $Q^{(j)}=k_{j}+Q$, for some $k_{j}\in\mathbb{Z}^{n}$.
Let also $\xi_{m}$ be a sequence of functions in $C_{c}^{\infty}(Q,[0,1])$ such that $\xi_{m}\to\chi_{Q}$ (pointwise), as $m\to+\infty$.
We set
$\eta_{m,j}(x):=\eta(x)\xi_{m}(x-k_{j})$
and observe that
(3.14) $\eta_{m,j}\in C^{\infty}_{c}(Q^{(j)}).$
Moreover, for all $x\in\mathbb{R}^{n}$,
(3.15) $\begin{split}&\eta(x)=\eta(x)\sum_{j=1}^{N}\chi_{Q^{(j)}}(x)=\eta(x)\sum_{j=1}^{N}\chi_{Q}(x-k_{j})\\ &\qquad=\lim_{m\to+\infty}\eta(x)\sum_{j=1}^{N}\xi_{m}(x-k_{j})=\lim_{m\to+\infty}\sum_{j=1}^{N}\eta_{m,j}(x).\end{split}$
In light of (3.14) and (3.10), we can employ (3.9) with $\eta:=\eta_{m,j}$, obtaining that
(3.16) $\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\frac{1}{2}z(x,y)\left(\eta_{m,j}(x)-\eta_{m,j}(y)\right)\,K(x,y)\,dx\,dy+\int_{\mathbb{R}^{n}}g(x)\,\eta_{m,j}(x)\,dx=0.$
Therefore, recalling (3.15), we sum (3.16) over $j$ and pass to the limit as $m\to+\infty$ (using the Dominated Convergence Theorem) to obtain the desired result. ∎
4. Minimality of level sets - Proof of Theorem 1.9
This section is committed to the proof of Theorem 1.9. Our strategy goes as follows: we first prove a weaker version of Theorem 1.9, in which the conclusion holds only for a.e. $t\in\mathbb{R}$, then we conclude thanks to the closedness of class-$A$ minimizers with respect to the $L^{1}_{\rm{loc}}$-convergence. Therefore, we now focus on showing the following:
Proposition 4.1.
Let $K:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ satisfy (1.2), (1.3), (1.4), and (1.5) and let $u_{p}$ be a minimizer for $\text{{E}}_{p}$ in $\mathcal{W}$. Let also $v_{p}$ and $E_{p,t}$ be as in (1.15) and (1.16), respectively.
Then, for a.e. $t\in\mathbb{R}$, the set $E_{p,t}$ is a class-$A$ minimizer for $\mathcal{J}$ in the sense of Definition 1.7.
Our argument relies on the following technical results:
Lemma 4.2.
Let $K:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ satisfy (1.3) and (1.4).
Then, for every set $E$, we have that
$\operatorname{P}_{K}(E,\Omega)<+\infty.$
Proof.
Thanks to the assumptions on $K$ in (1.3) and (1.4), we have that
$\begin{split}&\operatorname{P}_{K}(E,\Omega)=\iint_{{\Omega}_{\sharp}}\frac{1}{2}|\chi_{E}(x)-\chi_{E}(y)|\,K(x,y)\,dx\,dy=\iint_{{\Omega}_{\sharp}}\chi_{E}(x)\chi_{E^{c}}(y)K(x,y)\,dx\,dy\\ &\qquad=\iint_{{\Omega}_{\sharp}\cap\{|x-y|<\delta\}}\chi_{E}(x)\chi_{E^{c}}(y)K(x,y)\,dx\,dy+\iint_{{\Omega}_{\sharp}\cap\{|x-y|\geqslant\delta\}}\chi_{E}(x)\chi_{E^{c}}(y)K(x,y)\,dx\,dy\\ &\qquad\leqslant\iint_{{\Omega}_{\sharp}}\frac{\kappa_{2}\chi_{E}(x)\chi_{E^{c}}(y)}{|x-y|^{n+2s_{1}}}\,dx\,dy+\frac{2}{\delta}\iint_{\Omega\times\mathbb{R}^{n}}|x-y|K(x,y)\,dx\,dy\\ &\qquad\leqslant\iint_{E\times E^{c}}\frac{\kappa_{2}}{|x-y|^{n+2s_{1}}}\,dx\,dy+\frac{2|\Omega|}{\delta}\int_{\mathbb{R}^{n}}|h|K(h,0)\,dh\\ &\qquad<+\infty,\end{split}$
as desired. ∎
Lemma 4.3.
Let $K:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ satisfy (1.3) and (1.4) and let $u_{p}$ be a minimizer for $\text{{E}}_{p}$ in $\mathcal{W}$. Let also $v_{p}$ and $E_{p,t}$ be as in (1.15) and (1.16), respectively.
Then, for a.e. $t\in\mathbb{R}$,
(4.1) $\operatorname{P}_{K}(E_{p,t},\Omega)=\iint_{\Omega_{\sharp}}\frac{1}{2}z(x,y)\Big(\chi_{E_{p,t}}(x)-\chi_{E_{p,t}}(y)\Big)\,K(x,y)\,dx\,dy<+\infty.$
Proof.
First, observe that, thanks to Lemma 4.2,
$\operatorname{P}_{K}(E_{p,t},\Omega)<+\infty.$
Now, by the layer cake representation (or by direct inspection of the integrals), for any function $f$, we have that
$f(x)-f(y)=\int_{-\infty}^{+\infty}\Big(\chi_{\{f>t\}}(x)-\chi_{\{f>t\}}(y)\Big)\,dt,$
as well as
$|f(x)-f(y)|=\int_{-\infty}^{+\infty}\Big|\chi_{\{f>t\}}(x)-\chi_{\{f>t\}}(y)\Big|\,dt.$
Thus, by virtue of (1.12), we come to
(4.2) $\begin{split}z(x,y)\int_{-\infty}^{+\infty}\Big(\chi_{E_{p,t}}(x)-\chi_{E_{p,t}}(y)\Big)\,dt&=z(x,y)\int_{-\infty}^{+\infty}\Big(\chi_{\{v_{p}>t\}}(x)-\chi_{\{v_{p}>t\}}(y)\Big)\,dt\\ &=z(x,y)\big(v_{p}(x)-v_{p}(y)\big)\\ &=\big|v_{p}(x)-v_{p}(y)\big|\\ &=\int_{-\infty}^{+\infty}\Big|\chi_{\{v_{p}>t\}}(x)-\chi_{\{v_{p}>t\}}(y)\Big|\,dt\\ &=\int_{-\infty}^{+\infty}\Big|\chi_{E_{p,t}}(x)-\chi_{E_{p,t}}(y)\Big|\,dt.\end{split}$
Let us introduce the short notation
$\Upsilon(t):=\Big|\chi_{E_{p,t}}(x)-\chi_{E_{p,t}}(y)\Big|-z(x,y)\Big(\chi_{E_{p,t}}(x)-\chi_{E_{p,t}}(y)\Big).$
Then, using that $|z(x,y)|\leqslant 1$ and (4.2), we infer that $\Upsilon\geqslant 0$, and
$\int_{-\infty}^{+\infty}\Upsilon(t)\,dt=0.$
As a consequence, $\Upsilon$ vanishes for a.e. $t\in\mathbb{R}$, and hence
$\Big|\chi_{E_{p,t}}(x)-\chi_{E_{p,t}}(y)\Big|=z(x,y)\Big(\chi_{E_{p,t}}(x)-\chi_{E_{p,t}}(y)\Big).$
This leads to
$\operatorname{P}_{K}(E_{p,t},\Omega)=\iint_{{\Omega}_{\sharp}}\frac{1}{2}z(x,y)\Big(\chi_{E_{p,t}}(x)-\chi_{E_{p,t}}(y)\Big)\,K(x,y)\,dx\,dy,$
as desired. ∎
Proof of Proposition 4.1.
Let $t\in\mathbb{R}$. We consider a set $F$ such that $F\setminus\Omega=E_{p,t}\setminus\Omega$. By Lemma 4.2, we have that $\operatorname{P}_{K}(F,\Omega)<+\infty$.
We define the function $\widetilde{\chi}:=\chi_{E_{p,t}}-\chi_{F}$. We notice that $\widetilde{\chi}=0$ in $\Omega^{c}$ and
$\|\widetilde{\chi}\|_{L^{1}(\Omega)}+\iint_{{\Omega_{\sharp}}}\frac{1}{2}|\widetilde{\chi}(x)-\widetilde{\chi}(y)|K(x,y)\,dx\,dy\leqslant 2|\Omega|+\operatorname{P}_{K}(E_{p,t},\Omega)+\operatorname{P}_{K}(F,\Omega)<+\infty.$
Thus, by [MR3310082, Theorem 6], there exists [^3] On the one hand, we point out that Theorem 6 in [MR3310082] is stated and proved for general open sets $\Omega\subseteq\mathbb{R}^{n}$ with continuous boundary and the space $X^{s,p}_{0}:=\left\{u\in L^{p}(\Omega)\mbox{ s.t. }\|u\|_{L^{p}(\Omega)}+\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}|u(x)-u(y)|K(x-y)\,dx\,dy<+\infty\mbox{ and }u=0\mbox{ in }\Omega^{c}\right\},$ for some $p\geqslant 1$, and $K:\mathbb{R}^{n}\setminus\{0\}\to\mathbb{R}$. In particular, the shape of the kernel $K$ plays a role in proving [MR3310082, Theorem 6] (see in particular [MR3310082, Lemma 12]). On the other hand, by inspection of the proofs contained there, we see that analogous results with a more general kernel $K:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ satisfying (1.2), (1.3), (1.4), and (1.5) can be established using the same arguments and considering the space $X^{s,p}_{0}\cap L^{\infty}(\Omega)$, as in the setting studied here. a sequence of functions $\eta_{\ell}\in C_{c}^{\infty}(\Omega,[-1,1])$ such that
(4.3) $\lim_{\ell\to+\infty}\iint_{{\Omega}_{\sharp}}\frac{1}{2}\Big|\big(\chi_{E_{p,t}}-\chi_{F}-\eta_{\ell}\big)(x)-\big(\chi_{E_{p,t}}-\chi_{F}-\eta_{\ell}\big)(y)\Big|\,K(x,y)\,dx\,dy=0$
(4.4) and $\lim_{\ell\to+\infty}\|\chi_{E_{p,t}}-\chi_{F}-\eta_{\ell}\|_{L^{1}(\Omega)}=0.$
Owing to Corollary 3.2, we have that, for every $\ell\in\mathbb{N}$,
$\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\frac{1}{2}z(x,y)\big(\eta_{\ell}(x)-\eta_{\ell}(y)\big)\,K(x,y)\,dx\,dy+\int_{\mathbb{R}^{n}}g(x)\,\eta_{\ell}(x)\,dx=0.$
In this way, using also the Dominated Convergence Theorem, we obtain
(4.5) $\begin{split}&\int_{F\cap\Omega}g(x)\,dx-\int_{E_{p,t}\cap\Omega}g(x)\,dx=-\int_{\mathbb{R}^{n}}g(x)\,\big(\chi_{E_{p,t}}(x)-\chi_{F}(x)\big)\,dx\\ &\qquad=-\lim_{\ell\to+\infty}\int_{\mathbb{R}^{n}}g(x)\,\eta_{\ell}(x)\,dx\\ &\qquad=\lim_{\ell\to+\infty}\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\frac{1}{2}z(x,y)\big(\eta_{\ell}(x)-\eta_{\ell}(y)\big)\,K(x,y)\,dx\,dy.\end{split}$
Now, from Lemma 4.3, we know that
$\operatorname{P}_{K}(E_{p,t},\Omega)=\iint_{{\Omega}_{\sharp}}\frac{1}{2}z(x,y)\Big(\chi_{E_{p,t}}(x)-\chi_{E_{p,t}}(y)\Big)\,K(x,y)\,dx\,dy.$
Since $|z|\leqslant 1$, we also have that
$\operatorname{P}_{K}(F,\Omega)\geqslant\iint_{{\Omega}_{\sharp}}\frac{1}{2}z(x,y)\Big(\chi_{F}(x)-\chi_{F}(y)\Big)\,K(x,y)\,dx\,dy$
and hence
$\operatorname{P}_{K}(E_{p,t},\Omega)-\operatorname{P}_{K}(F,\Omega)\leqslant\iint_{{\Omega}_{\sharp}}\frac{1}{2}z(x,y)\Big(\big(\chi_{E_{p,t}}-\chi_{F}\big)(x)-\big(\chi_{E_{p,t}}-\chi_{F}\big)(y)\Big)\,K(x,y)\,dx\,dy.$
Using this and (4.3), we find that
$\begin{split}\operatorname{P}_{K}(E_{p,t},\Omega)-\operatorname{P}_{K}(F,\Omega)&\leqslant\lim_{\ell\to+\infty}\iint_{{\Omega}_{\sharp}}\frac{1}{2}z(x,y)\big(\eta_{\ell}(x)-\eta_{\ell}(y)\big)\,K(x,y)\,dx\,dy\\ &=\lim_{\ell\to+\infty}\iint_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\frac{1}{2}z(x,y)\big(\eta_{\ell}(x)-\eta_{\ell}(y)\big)\,K(x,y)\,dx\,dy,\end{split}$
where the last equality is due to the fact that $\eta_{\ell}$ vanishes outside $\Omega$.
This, in tandem with (4.5), yields
(4.6) $\operatorname{P}_{K}(E_{p,t},\Omega)-\operatorname{P}_{K}(F,\Omega)\leqslant\int_{F\cap\Omega}g(x)\,dx-\int_{E_{p,t}\cap\Omega}g(x)\,dx,$
showing the class-$A$ minimality of $E_{p,t}$. ∎
Now, since we think it is interesting per se, we show that the functional $\mathcal{J}$ is continuous with respect to the $L^{1}_{\rm{loc}}$-convergence of class-$A$ minimizers. As a byproduct, we obtain that class-$A$ minimizers are closed with respect to the $L^{1}_{\rm{loc}}$-convergence, i.e. the limit of a convergent sequence of class-$A$ minimizers in $L^{1}_{\rm{loc}}$ is a class-$A$ minimizer itself. Our proof follows the ideas of [MR2675483, Theorem 3.3] (see also [sequoia, Proposition 2.1]).
Lemma 4.4.
Let $K:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ satisfy (1.3) and (1.4) and let $\{E_{j}\}_{j}$ be a sequence of class-$A$ minimizers for $\mathcal{J}$ such that $E_{j}\to E$ in $L^{1}_{\rm{loc}}(\mathbb{R}^{n})$, for some set $E\subseteq\mathbb{R}^{n}$.
Then, $E$ is a class-$A$ minimizer for $\mathcal{J}$.
Moreover, for any domain $\Omega\subseteq\mathbb{R}^{n}$, we have
(4.7) $\lim_{j\to+\infty}\mathcal{J}(E_{j},\Omega)=\mathcal{J}(E,\Omega).$
Proof.
Let $\Omega$ be a smooth domain of $\mathbb{R}^{n}$. Let also $F\subseteq\mathbb{R}^{n}$ be a competitor for $E$ in $\Omega$, namely $F\setminus\Omega=E\setminus\Omega$. Define the sequence of sets $\{F_{j}\}_{j}$ as
$F_{j}:=(F\cap\Omega)\cup(E_{j}\cap\Omega^{c}).$
In particular, we have $F_{j}\setminus\Omega=E_{j}\setminus\Omega$ and therefore, by the class-$A$ minimality of $E_{j}$,
$\mathcal{J}(E_{j},\Omega)\leqslant\mathcal{J}(F_{j},\Omega).$
Consequently, thanks to the lower semicontinuity of $\operatorname{P}_{K}$ and the $L^{1}_{\rm{loc}}$-convergence of $E_{j}$ to $E$ (together with the fact that $g$ is bounded), we infer that
(4.8) $\mathcal{J}(E,\Omega)\leqslant\liminf_{j\to+\infty}\mathcal{J}(E_{j},\Omega)\leqslant\liminf_{j\to+\infty}\mathcal{J}(F_{j},\Omega).$
Now we observe that
$\begin{split}\mathcal{J}(F_{j},\Omega)&=\operatorname{P}_{K}(F_{j},\Omega)+\int_{F_{j}\cap\Omega}g(x)\,dx\\ &=\operatorname{P}_{K}(F,\Omega)+2\iint_{\Omega\times\Omega^{c}}\chi_{F^{c}}(x)\big(\chi_{E_{j}}(y)-\chi_{F}(y)\big)K(x,y)\,dx\,dy+\int_{F\cap\Omega}g(x)\,dx\\ &=\mathcal{J}(F,\Omega)+2\iint_{\Omega\times\Omega^{c}}\chi_{F^{c}}(x)\big(\chi_{E_{j}}(y)-\chi_{E}(y)\big)K(x,y)\,dx\,dy.\end{split}$
This, recalling also (1.4), yields that
(4.9) $\begin{split}|\mathcal{J}(F_{j},\Omega)-\mathcal{J}(F,\Omega)|\leqslant\;&2\iint_{\Omega\times\Omega^{c}}\chi_{F^{c}}(x)\chi_{E_{j}\Delta E}(y)K(x,y)\,dx\,dy\\ \leqslant\;&2\kappa_{2}\iint_{\Omega\times\Omega^{c}}\frac{\chi_{E_{j}\Delta E}(y)}{|x-y|^{n+2s_{1}}}\,dx\,dy\\ =:\;&b_{j}.\end{split}$
Now, we have that
$\lim_{j\to+\infty}b_{j}=0,$
see [sequoia, Appendix A] (see also the proof of [MR2675483, Theorem 3.3]).
It follows from this and (4.9) that
(4.10) $\limsup_{j\to+\infty}\mathcal{J}(F_{j},\Omega)\leqslant\limsup_{j\to+\infty}\left(\mathcal{J}(F,\Omega)+b_{j}\right)=\mathcal{J}(F,\Omega).$
This and (4.8) give that
(4.11) $\mathcal{J}(E,\Omega)\leqslant\mathcal{J}(F,\Omega),$
hence $E$ is a minimizer for $\mathcal{J}$ in $\Omega$.
In addition, choosing
$F_{j}:=(E\cap\Omega)\cup(E_{j}\cap\Omega^{c}),$
we deduce from (4.8) and (4.10) that
$\mathcal{J}(E,\Omega)\leqslant\liminf_{j\to+\infty}\mathcal{J}(E_{j},\Omega)\leqslant\limsup_{j\to+\infty}\mathcal{J}(E_{j},\Omega)\leqslant\mathcal{J}(E,\Omega),$
which entails (4.7) for every smooth domain $\Omega$.
Moreover, if $\partial\Omega$ is not $\mathcal{C}^{\infty}$, by interior smooth approximation, there exists a sequence $\{\Omega^{h}\}_{h}$ of smooth domains such that $\Omega^{h}\subseteq\Omega$, for every $h$, and $\Omega^{h}\to\Omega$ locally in $\mathbb{R}^{n}$. Then, since $\operatorname{P}_{K}$ is continuous with respect to the $L^{1}_{\rm{loc}}$-convergence of the domain, we obtain (4.7) and (4.11) for every domain $\Omega\subseteq\mathbb{R}^{n}$.
By the arbitrariness of $\Omega$, we conclude that $E$ is class-$A$ minimizer for $\mathcal{J}$. ∎
Proof of Theorem 1.9.
Let $t\in\mathbb{R}$ and consider a sequence $\{t_{j}\}_{j}\subseteq\mathbb{R}$ such that $t_{j}\to t$, as $j\to+\infty$, and Proposition 4.1 holds true for every $t_{j}$. Moreover, up to take a subsequence, suppose that
$t\leqslant t_{j}\leqslant t_{j-1}\leqslant\dots\leqslant t_{0}.$
Thus, recalling that $E_{p,t}=\{v_{p}>t\}$ (see (1.16)), we have that, for a.e. $x\in\mathbb{R}^{n}$,
$\lim_{j\to+\infty}|\chi_{E_{p,t}}(x)-\chi_{E_{p,t_{j}}}(x)|=\lim_{j\to+\infty}|\chi_{\{v_{p}>t\}}(x)-\chi_{\{v_{p}>t_{j}\}}(x)|=\lim_{j\to+\infty}\chi_{\{t<v_{p}\leqslant t_{j}\}}(x)=0.$
Therefore, for any domain $\Omega\subseteq\mathbb{R}^{n}$, thanks to the Dominated Convergence Theorem (with dominant $\chi_{E_{p,t}}\in L^{1}(\Omega)$), we obtain
$\lim_{j\to+\infty}\int_{\Omega}|\chi_{E_{p,t}}(x)-\chi_{E_{p,t_{j}}}(x)|\,dx=0.$
Hence,
$E_{p,t_{j}}\to E_{p,t}\quad\mbox{in }L^{1}_{\rm{loc}}(\mathbb{R}^{n}).$
Thus, thanks to Lemma 4.4, we infer that also $E_{p,t}$ is a class-$A$ minimizer for $\mathcal{J}$ and this concludes the proof. ∎
5. Density estimates for minimizers of the functional $\mathscr{F}$
In this section we show that the level sets $E_{p,t}$ defined in (1.16) satisfy some uniform density estimates. Heuristically, if $x_{0}\in\partial E_{p,t}$, then, for any small $r>0$, the sets $B_{r}(x_{0})\cap E_{p,t}$ and $B_{r}(x_{0})\cap E_{p,t}^{c}$ have comparable measure. The precise statement goes as follows:
Proposition 5.1.
Let $K:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ satisfy (1.2), (1.3), (1.4), and (1.5) and let $u_{p}$ be a minimizer for $\text{{E}}_{p}$ in $\mathcal{W}$.
Let also $v_{p}(x)$ and $E_{p,t}$ be defined as in (1.15) and (1.16), respectively.
Then, there exists a constant $c_{0}\in(0,1)$, depending only on $n$, $s_{1}$, and $\kappa_{2}$, such that, for any $x_{0}\in(\partial E)\cap Q$ and $r\in\left(0,\min\{\delta/4,\mbox{dist}(x_{0},\partial Q)\}\right)$,
(5.1) $c_{0}r^{n}\leqslant|E_{p,t}\cap B_{r}(x_{0})|\leqslant(1-c_{0})r^{n},$
whenever $\|g\|_{L^{n/2s_{1}}(Q)}<(c_{0}/2)^{-n/2s_{1}}$.
This result is a direct consequence of the fact that every set $E_{p,t}$ is almost minimal in the following sense:
Definition 5.2 (Almost-minimality).
Let $\Omega\subseteq\mathbb{R}^{n}$ be a Lipschitz domain. Let $\Lambda>0$ and $\xi\in\mathbb{R}$. We say that a set $E\subseteq\mathbb{R}^{n}$ with finite $K$-perimeter is $(\Lambda,\xi)$-minimal in $\Omega$ if
$P_{K}(E,\Omega)\leqslant\operatorname{P}_{K}(F,\Omega)+\Lambda|E\Delta F|^{1-\xi},$
for any set $F\subseteq\mathbb{R}^{n}$ such that $F\setminus\Omega=E\setminus\Omega$.
In [sequoia], the authors have proved the Hölder-regularity for $(\Lambda,0)$-minimizers of the $2s$-perimeter (compare Definition 5.2 here with Definition 1.1 in [sequoia]). In particular, they have shown in [sequoia, Theorem 2.2] that $(\Lambda,0)$-minimal sets satisfy suitable uniform density estimates. A slight modification of the argument used there produces uniform density estimates for $\left(\Lambda,\frac{2s}{n}\right)$-minimal sets, for any $s\in(0,1/2)$. We refer to Appendix B for a detailed proof. Here, we focus on proving the following:
Lemma 5.3.
Let $K:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ satisfy (1.2), (1.3), (1.4), and (1.5) and let $u_{p}$ be a minimizer for $\text{{E}}_{p}$ in $\mathcal{W}$.
Let also $v_{p}$ and $E_{p,t}$ be defined as in (1.15) and (1.16), respectively.
Then, $E_{p,t}$ is a $\left(\|g\|_{L^{n/2s_{1}}(Q)},\frac{2s_{1}}{n}\right)$-minimal set in $Q$ (in the sense of Definition 5.2).
Proof.
Let $F\subseteq\mathbb{R}^{n}$ be a set such that $F\setminus Q=E_{p,t}\setminus Q$. Then, by Theorem 1.9 and using the Hölder inequality, we infer that
$\begin{split}&\operatorname{P}_{K}(E_{p,t},Q)-\operatorname{P}_{K}(F,Q)=\mathcal{J}(E_{p,t},Q)-\mathcal{J}(F,Q)-\int_{Q}g(x)\left(\chi_{E_{p,t}}(x)-\chi_{F}(x)\right)dx\\ &\qquad\qquad\leqslant\int_{Q}|g(x)|\left|\chi_{E_{p,t}}(x)-\chi_{F}(x)\right|dx\leqslant\|g\|_{L^{n/2s_{1}}(Q)}|E_{p,t}\Delta F|^{1-\frac{2s_{1}}{n}}.\qed\end{split}$
With this, we can now complete the proof of Proposition 5.1.
Proof of Proposition 5.1.
According to Lemma 5.3, $E_{p,t}$ is a $\left(\|g\|_{L^{n/2s_{1}}(Q)},\frac{2s_{1}}{n}\right)$-minimal set in $Q$ in the sense of Definition 5.2. Thus, Proposition B.1 yields the desired result provided that $\|g\|_{L^{n/2s_{1}}(Q)}<(c_{0}/2)^{-n/2s_{1}}$ (e.g up to considering $\gamma$ in Theorem 1.4 small enough). ∎
6. Minimizers of $\text{{E}}_{p}$ have controlled oscillations - Proof of Theorem 1.10
In this section, we focus on the proof of Theorem 1.10. Our argument relies on the following result which provides a lower bound for $\operatorname{P}_{K}$.
In the following, we will use the notation $\widetilde{Q}_{{q},\zeta}:={q}+(-\zeta,\zeta)^{n}$, with $q\in\mathbb{R}^{n}$ and $\zeta>0$.
Lemma 6.1.
Let $E\subseteq\mathbb{R}^{n}$ satisfy the uniform density estimates in Proposition 5.1 in a Lipschitz domain $\Omega\subseteq\mathbb{R}^{n}$.
Moreover, suppose that there exists a cube $\widetilde{Q}_{\bar{q},\bar{\zeta}}\subseteq\Omega$, for some $\bar{q}\in\Omega$ and some small $\bar{\zeta}>0$, such that $\widetilde{Q}_{\bar{q},3\bar{\zeta}}\subseteq\Omega$ and $\widetilde{Q}_{\bar{q},\bar{\zeta}}\cap\partial E\neq\varnothing$.
Then, there exists a constant $C>0$, depending only on $n$, $s_{1}$, $\kappa_{1}$, $\kappa_{2}$, $\delta$, and $\Omega$, such that
(6.1) $P_{K}(E,\Omega)\geqslant C.$
Proof.
Let us decompose $\mathbb{R}^{n}$ into disjoint cubes of side $\zeta$, for some given $\zeta\in\left(0,\min\left\{\frac{\delta}{6\sqrt{n}},\bar{\zeta}\right\}\right)$, with $\delta$ as in (1.4). Consider the set
$\text{{Q}}_{\ \zeta}:=\big\{\widetilde{Q}_{q,\zeta}\text{ such that }\widetilde{Q}_{q,\zeta}\subseteq\widetilde{Q}_{q,3\zeta}\subseteq\Omega,\ \widetilde{Q}_{q,\zeta}\cap\partial E\neq\varnothing\big\}.$
Notice that $\text{{Q}}_{\ \zeta}\neq\varnothing$ since $\widetilde{Q}_{\bar{q},\bar{\zeta}}\in\text{{Q}}_{\ \zeta}$. Define $N_{\zeta}:=\#(\text{{Q}}_{\ \zeta})$, namely the number of elements of $\text{{Q}}_{\ \zeta}$.
Let $x_{0}\in\widetilde{Q}_{q,\zeta}\cap\partial E$, for some $\widetilde{Q}_{q,\zeta}\in\text{{Q}}_{\ \zeta}$. Then, we have that $B_{\zeta}(x_{0})\subseteq\widetilde{Q}_{q,3\zeta}\subseteq\Omega$. Thus, since $E$ satisfies (5.1) by assumption, we obtain that
$|E\cap\widetilde{Q}_{q,3\zeta}|\geqslant c_{0}\zeta^{n}\qquad{\mbox{and}}\qquad|E^{c}\cap\widetilde{Q}_{q,3\zeta}|\geqslant c_{0}\zeta^{n}.$
From this, using also that $|x-y|\leqslant 3\sqrt{n}\zeta<\delta$, for every $x$, $y\in\widetilde{Q}_{q,3\zeta}$, and (1.4), we deduce that
(6.2) $\begin{split}\mathcal{L}_{K}(E\cap\widetilde{Q}_{q,3\zeta},E^{c}\cap\widetilde{Q}_{q,3\zeta})&\geqslant\kappa_{1}\int_{E\cap\widetilde{Q}_{q,3\zeta}}\int_{E^{c}\cap\widetilde{Q}_{q,3\zeta}}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}\\ &\geqslant\kappa_{1}\int_{E\cap\widetilde{Q}_{q,3\zeta}}\int_{E^{c}\cap\widetilde{Q}_{q,3\zeta}}\big(3\sqrt{n}\zeta\big)^{-n-2s_{1}}\,dx\,dy\\ &=\kappa_{1}\big(3\sqrt{n}\zeta\big)^{-n-2s_{1}}|E\cap\widetilde{Q}_{q,3\zeta}|\,|E^{c}\cap\widetilde{Q}_{q,3\zeta}|\\ &\geqslant C\zeta^{n-2s_{1}},\end{split}$
for some $C>0$, depending only on $n$, $s_{1}$, $\kappa_{1}$, and $\kappa_{2}$.
Set now $\text{{R}}_{\zeta}$ to be the family of all cubes $\widetilde{Q}_{q,\zeta}$ such that $\widetilde{Q}_{q,3\zeta}\subseteq\Omega$, and notice that $\text{{Q}}_{\ \zeta}\subseteq\text{{R}}_{\zeta}$. Thus, using (6.2), we see that
$\begin{split}&\mathcal{L}_{K}(E\cap\Omega,E^{c}\cap\Omega)\geqslant\sum_{\begin{subarray}{c}\widetilde{Q}_{q,\zeta}\in\text{{R}}_{\zeta},\\ \widetilde{Q}_{q^{\prime},\zeta}\in\text{{R}}_{\zeta}\end{subarray}}\mathcal{L}_{K}(E\cap\widetilde{Q}_{q,\zeta},E^{c}\cap\widetilde{Q}_{q^{\prime},\zeta})\geqslant 3^{-2n}\sum_{\begin{subarray}{c}\widetilde{Q}_{q,\zeta}\in\text{{R}}_{\zeta},\\ \widetilde{Q}_{q^{\prime},\zeta}\in\text{{R}}_{\zeta}\end{subarray}}\mathcal{L}_{K}(E\cap\widetilde{Q}_{q,3\zeta},E^{c}\cap\widetilde{Q}_{q^{\prime},3\zeta})\\ &\qquad\geqslant 3^{-2n}\sum_{\widetilde{Q}_{q,\zeta}\in\text{{R}}_{\zeta}}\mathcal{L}_{K}(E\cap\widetilde{Q}_{q,3\zeta},E^{c}\cap\widetilde{Q}_{q,3\zeta})\geqslant 3^{-2n}\sum_{\widetilde{Q}_{q,\zeta}\in\text{{Q}}_{\ \zeta}}\mathcal{L}_{K}(E\cap\widetilde{Q}_{q,3\zeta},E^{c}\cap\widetilde{Q}_{q,3\zeta})\\ &\qquad\geqslant 3^{-2n}CN_{\zeta}\zeta^{n-2s_{1}}.\end{split}$
From this, we infer that
$\operatorname{P}_{K}(E,\Omega)\geqslant 3^{-2n}CN_{\zeta}\zeta^{n-2s_{1}},$
which entails the desired estimate. ∎
Before proving Theorem 1.10, we observe that, since $u_{p}$ is $\mathbb{Z}^{n}$-periodic by assumption, we have that
$\operatorname*{osc}_{Q}(u_{p})=\operatorname*{osc}(u_{p}):=\operatorname*{osc}_{\mathbb{R}^{n}}(u_{p})=\operatorname*{ess\,sup}_{\mathbb{R}^{n}}u_{p}-\operatorname*{ess\,inf}_{\mathbb{R}^{n}}u_{p}.$
Thus, in what follows, we will simply write $\operatorname*{osc}(u_{p})$.
Proof of Theorem 1.10.
Let us recall that $E_{p,t}=\{v_{p}>t\}$ (see (1.16)). In virtue of Proposition 5.1, we have that $E_{p,t}$ satisfies the assumptions of Lemma 6.1 with $\Omega:=Q$. Thus, integrating (6.1) over the set
$\mathcal{T}_{p}:=\{t\in\mathbb{R}\mbox{ s.t. }E_{p,t}\cap Q\neq\varnothing\},$
we obtain
(6.3) $\int_{\mathcal{T}_{p}}\operatorname{P}_{K}(E_{p,t},Q)\,dt\geqslant C|\mathcal{T}_{p}|=C\operatorname*{osc}_{Q}(v_{p}),$
where $C$ is as in Lemma 6.1.
On the other hand, by the layer cake representation we have that
(6.4) $\begin{split}&\int_{\mathcal{T}_{p}}\operatorname{P}_{K}(E_{p,t},Q)\,dt\leqslant\int_{-\infty}^{+\infty}\left(\iint_{Q_{\sharp}}\frac{1}{2}|\chi_{E_{p,t}}(x)-\chi_{E_{p,t}}(y)|K(x,y)\,dx\,dy\right)\,dt\\ &\qquad\qquad=\iint_{Q_{\sharp}}\left(\int_{-\infty}^{+\infty}\frac{1}{2}|\chi_{E_{p,t}}(x)-\chi_{E_{p,t}}(y)|K(x,y)\,dt\right)\,dx\,dy\\ &\qquad\qquad=\iint_{Q_{\sharp}}\frac{1}{2}|v_{p}(x)-v_{p}(y)|K(x,y)\,dx\,dy\\ &\qquad\qquad\leqslant 2\iint_{Q\times\mathbb{R}^{n}}\frac{1}{2}|v_{p}(x)-v_{p}(y)|K(x,y)\,dx\,dy\\ &\qquad\qquad=2\left(\text{{E}}_{p}(u_{p})-\int_{Q}g(x)u_{p}(x)dx\right).\end{split}$
Also, by the minimality of $u_{p}$ and (1.3), we deduce that
(6.5) $\text{{E}}_{p}(u_{p})\leqslant\text{{E}}_{p}(0)\leqslant c|p|,$
for some constant $c>0$ independent of $p$.
Moreover, since $\int_{Q}u_{p}(x)\,dx=0$, we have that
$\operatorname*{ess\,inf}_{\mathbb{R}^{n}}u_{p}=\operatorname*{ess\,inf}_{Q}u_{p}\leqslant 0\leqslant\operatorname*{ess\,sup}_{Q}u_{p}=\operatorname*{ess\,sup}_{\mathbb{R}^{n}}u_{p},$
and hence $|u_{p}|\leqslant\operatorname*{osc}(u_{p})$. It thereby follows that
$\begin{split}&-\int_{Q}g(x)u_{p}(x)dx\leqslant\int_{Q}|g(x)||u_{p}(x)|dx\leqslant|Q|^{(n-s_{1})/n}\|g\|_{L^{n/s_{1}}(Q)}\operatorname*{osc}(u_{p})\\ &\qquad\qquad\leqslant\|g\|_{L^{n/s_{1}}(Q)}\left(\operatorname*{osc}_{Q}(v_{p})+\sqrt{n}|p|\right).\end{split}$
Plugging this and (6.5) into (6.4), we obtain
$\int_{\mathcal{T}_{p}}\operatorname{P}_{K}(E_{p,t},Q)dt\leqslant c|p|+2\|g\|_{L^{n/s_{1}}(Q)}\operatorname*{osc}_{Q}(v_{p}),$
up to renaming $c$.
This and (6.3) entail that
$\operatorname*{osc}_{Q}(v_{p})\leqslant c|p|,$
provided that $\|g\|_{L^{n/s_{1}}(Q)}\leqslant C/4$ (where $C$ is as in Lemma 6.1).
Besides,
$\operatorname*{osc}(u_{p})\leqslant\operatorname*{osc}_{Q}(v_{p})+\sqrt{n}|p|\leqslant(c+\sqrt{n})|p|.\qed$
Proof of Corollary 1.12.
For any $t\in\mathcal{T}_{p}$ (i.e. $t\in\mathbb{R}$ such that $E_{p,t}\cap Q\neq\varnothing$), by Theorem 1.9 we have that $E_{p,t}$ is a class-$A$ minimizer for $\mathcal{J}$.
Moreover, notice that for any $x_{1}\in E_{p,t}$, we have by definition that
(6.6) $p\cdot x_{1}>t-u_{p}(x_{1}).$
Now, since $\partial E_{p,t}\cap Q\neq\varnothing$, there exist $y_{1}\in Q$ such that
$u_{p}(y_{1})+p\cdot y_{1}\leqslant t.$
In particular, we have that
$u_{p}(y_{1})-\sqrt{n}|p|\leqslant t.$
Hence, using also (6.6), we infer that
$p\cdot x_{1}>t-u_{p}(x_{1})\geqslant u_{p}(y_{1})-u_{p}(x_{1})-\sqrt{n}|p|\geqslant-\operatorname*{osc}(u_{p})-\sqrt{n}|p|.$
From this and Theorem 1.10, it follows that
$p\cdot x_{1}>-\big(\operatorname*{osc}(u_{p})+\sqrt{n}|p|\big)\geqslant-(c+2\sqrt{n})|p|.$
Thus, setting $M:=c+2\sqrt{n}$, we obtain (1.17).
Analogously, for any $x_{2}\in E_{p,t}^{c}$, we have that
$p\cdot x_{2}\leqslant t-u_{p}(x_{2}).$
Moreover, from $\partial E_{p,t}\cap Q\neq\varnothing$, we deduce that there exist $y_{2}\in Q$ such that
$u_{p}(y_{2})+\sqrt{n}|p|\geqslant u_{p}(y_{2})+p\cdot y_{2}\geqslant t.$
We infer from the last two displays that
$p\cdot x_{2}\leqslant t-u_{p}(x_{2})\leqslant u_{p}(y_{2})-u_{p}(x_{2})+\sqrt{n}|p|\leqslant\operatorname*{osc}(u_{p})+\sqrt{n}|p|\leqslant M.$
This entails that
$E_{p,t}^{c}\subseteq\{x\in\mathbb{R}^{n}\mbox{ s.t. }x\cdot p\leqslant M|p|\}$
which reads (1.18), concluding the proof. ∎
7. Preliminary results for the $\Gamma$-convergence
In this section, we state and prove some preliminary results which are useful for the proof of the $\Gamma$-convergence stated in Theorem 1.13.
In what follows, we will often use a stronger version of [MR1634336, Theorem 2.8] to show that the energy contribution due to the interactions with the outside of the domain are negligible in the limit. We recall such result for the convenience of the reader.
Lemma 7.1 ([MR1634336, Theorem 2.8]).
Let $A$ and $B$ be open subsets of $\mathbb{R}^{n}$ and assume that $A$ is bounded with Lipschitz boundary. Moreover, define $\Sigma:=\partial A\cap\partial B$.
Let also $K$ satisfy (1.2), (1.3), (1.4), and (1.5) and let $\{\varepsilon_{j}\}_{j}$ be an infinitesimal sequence.
Consider a sequence of functions $\{u_{j}:A\cup B\to[-1,1]\}_{j}$ and assume that there exists an infinitesimal sequence $\{\eta_{j}\}_{j}$ such that
$\lim_{j\to+\infty}u_{j}(x+\eta_{j}h)=v(x),\quad\mbox{for a.e.\penalty 10000\ $x\in\Sigma$ and a.e\penalty 10000\ $h\in A$,}$
and $\lim_{j\to+\infty}u_{j}(x+\eta_{j}h)=w(x),\quad\mbox{for a.e.\penalty 10000\ $x\in\Sigma$ and a.e.\penalty 10000\ $h\in B$,}$
for some functions $v$, $w:\Sigma\to[-1,1].$
Then, there exists a constant $C_{K}>0$ depending only on $K$ such that
$\limsup_{j\to+\infty}\iint_{A\times B}|u_{j}(x)-u_{j}(y)|K_{\varepsilon_{j}}(x,y)\,dx\,dy\leqslant C_{K}\int_{\Sigma}|v(x)-w(x)|\,d\mathcal{H}^{n-1}(x).$
Notice that, in our setting, a suitable rescaled sequence of planelike minimizers (for $\mathcal{J}$) satisfies the assumptions of Lemma 7.1.
For later convenience, for any $p\in\mathbb{S}^{n-1}$ let us define the set
(7.1) $\mathcal{I}_{p}:=\{x\cdot p<0\}.$
Then, we have the following:
Lemma 7.2.
Let $p\in\mathbb{S}^{n-1}$ and let $E_{p}$ be a set such that
(7.2) $\{x\cdot p\leqslant-M\}\subseteq E_{p}\subseteq\{x\cdot p\leqslant M\},$
for some $M>0$.
Let $\{\varepsilon_{j}\}_{j}$ be an infinitesimal sequence and define $u_{j}:=\chi_{\varepsilon_{j}E_{p}}$.
Then, for every Lipschitz domain $\Omega$, we have that, for a.e. $x\in\partial\Omega$ and a.e. $h\in\mathbb{R}^{n}$,
$\lim_{j\to+\infty}u_{j}(x+\varepsilon_{j}^{2}h)=\chi_{\mathcal{I}_{p}}(x).$
Proof.
We observe that, thanks to (7.2),
(7.3) $\{x\cdot p\leqslant-\varepsilon_{j}M\}\subseteq\varepsilon_{j}E_{p}\subseteq\{x\cdot p\leqslant\varepsilon_{j}M\}.$
Now, let $x\in\partial\Omega\setminus\partial\mathcal{I}_{p}$, then $x\cdot p\neq 0$ and therefore $|x\cdot p|>\varepsilon_{j}(M+1)$ as soon as $j$ is large enough. Moreover, if $h\in(\partial\Omega\cap\partial\mathcal{I}_{p})^{c}$, possibly taking $j$ larger, we can assume that $\varepsilon_{j}|h|\leqslant 1$.
Suppose first that $x\cdot p>\varepsilon_{j}(M+1)$. In this case, we see that
$(x+\varepsilon_{j}^{2}h)\cdot p=x\cdot p+\varepsilon_{j}^{2}h\cdot p\geqslant x\cdot p-\varepsilon_{j}^{2}|h|>\varepsilon_{j}(M+1)-\varepsilon_{j}=\varepsilon_{j}M.$
It thereby follows from this and (7.3) that, for any $j$ large enough,
$u_{j}(x+\varepsilon_{j}^{2}h)-\chi_{\mathcal{I}_{p}}(x)=\chi_{\varepsilon_{j}E_{p}}(x+\varepsilon_{j}^{2}h)-\chi_{\mathcal{I}_{p}}(x)=0.$
If instead $x\cdot p<-\varepsilon_{j}(M+1)$, then
$(x+\varepsilon_{j}^{2}h)\cdot p=x\cdot p+\varepsilon_{j}^{2}h\cdot p\leqslant x\cdot p+\varepsilon_{j}^{2}|h|<-\varepsilon_{j}(M+1)+\varepsilon_{j}=-\varepsilon_{j}M.$
Thus, recalling (7.3), we get that, for any $j$ large enough,
$u_{j}(x+\varepsilon_{j}^{2}h)-\chi_{\mathcal{I}_{p}}(x)=\chi_{\varepsilon_{j}E_{p}}(x+\varepsilon_{j}^{2}h)-\chi_{\mathcal{I}_{p}}(x)=1-1=0.$
These considerations establish the desired limit. ∎
Another crucial step to prove Theorem 1.13 is showing that, for every set $F\subseteq\mathbb{R}^{n}$,
$\mathscr{E}(F,B_{R}):=\iint_{B_{R}\times\mathbb{R}^{n}}\chi_{F}(x)\chi_{F^{c}}(y)K(x,y)\,dx\,dy+\int_{\mathcal{Q}(B_{R})_{1}\cap F}g(x)\,dx$
(namely (1.20) with $\varepsilon:=1$ and $\Omega:=B_{R}$) is non-negative in any large ball.
Proposition 7.3.
Let $K$ satisfy (1.2), (1.3), (1.4), and (1.5).
Then, if $\|g\|_{L^{\infty}(\mathbb{R}^{n})}\leqslant\frac{\kappa_{3}}{2}$, we have that
$\mathscr{E}(F,B_{R})\geqslant 0,$
for every set $F\subseteq\mathbb{R}^{n}$ and for every $R>1$.
This result is a consequence of the following lemma.
Lemma 7.4.
Let $K$ satisfy (1.2), (1.3), (1.4), and (1.5).
Then, for every set $F\subseteq\mathbb{R}^{n}$,
$\iint_{Q\times Q}\chi_{F}(x)\chi_{F^{c}}(y)K(x,y)\,dx\,dy+\int_{Q\cap F}g(x)\,dx\geqslant\left(\frac{\kappa_{3}}{2}-\|g\|_{L^{\infty}(\mathbb{R}^{n})}\right)\min\{|F\cap Q|,|F^{c}\cap Q|\}.$
Proof.
Notice that
(7.4) $\begin{split}|F\cap Q|\,|F^{c}\cap Q|&=\min\{|F\cap Q|,|F^{c}\cap Q|\}\Big(1-\min\{|F\cap Q|,|F^{c}\cap Q|\}\Big)\\ &\geqslant\frac{1}{2}\min\{|F\cap Q|,|F^{c}\cap Q|\}.\end{split}$
Moreover, recalling (1.1), we have that
$\int_{Q\cap F}g(x)\,dx=-\int_{Q\cap F^{c}}g(x)\,dx\geqslant-\|g\|_{L^{\infty}(\mathbb{R}^{n})}\min\{|F\cap Q|,|F^{c}\cap Q|\}.$
Thus, from this, (1.5), and (7.4), we obtain that
$\begin{split}&\iint_{Q\times Q}\chi_{F}(x)\chi_{F^{c}}(y)K(x,y)\,dx\,dy+\int_{Q\cap F}g(x)\,dx\\ &\qquad\qquad\geqslant\kappa_{3}|F\cap Q|\,|F^{c}\cap Q|-\|g\|_{L^{\infty}(\mathbb{R}^{n})}\min\{|F\cap Q|,|F^{c}\cap Q|\}\\ &\qquad\qquad\geqslant\left(\frac{\kappa_{3}}{2}-\|g\|_{L^{\infty}(\mathbb{R}^{n})}\right)\min\{|F\cap Q|,|F^{c}\cap Q|\},\end{split}$
showing the result. ∎
Proof of Proposition 7.3.
Notice that
$\begin{split}\mathscr{E}(F,B_{R})&=\iint_{B_{R}\times\mathbb{R}^{n}}\chi_{F}(x)\chi_{F^{c}}(y)K(x,y)\,dx\,dy+\int_{\mathcal{Q}(B_{R})_{1}\cap F}g(x)\,dx\\ &\geqslant\iint_{\mathcal{Q}(B_{R})_{1}\times\mathbb{R}^{n}}\chi_{F}(x)\chi_{F^{c}}(y)K(x,y)\,dx\,dy+\int_{\mathcal{Q}(B_{R})_{1}\cap F}g(x)\,dx\\ &\geqslant\sum_{\begin{subarray}{c}j\in\mathbb{Z}^{n}\\ j+Q\subseteq B_{R}\end{subarray}}\left(\iint_{(F\cap(j+Q))\times(F^{c}\cap(j+Q))}K(x,y)\,dx\,dy+\int_{F\cap(j+Q)}g(x)\,dx\right).\end{split}$
In light if this, the claim in Proposition 7.3 will follow if we show that
(7.5) $\iint_{(F\cap(j+Q))\times(F^{c}\cap(j+Q))}K(x,y)dxdy+\int_{F\cap(j+Q)}g(x)dx\geqslant 0,$
for every $j\in\mathbb{Z}^{n}$ such that $j+Q\subseteq B_{R}$.
Now, using Lemma 7.4 in combination with (1.2) and the periodicity $g$, we deduce that
$\begin{split}&\iint_{(F\cap(j+Q))\times(F^{c}\cap(j+Q))}K(x,y)\,dx\,dy+\int_{F\cap(j+Q)}g(x)\,dx\\ &\qquad\qquad=\iint_{((F-j)\cap Q)\times((F^{c}-j)\cap Q)}K(x,y)\,dx\,dy+\int_{(F-j)\cap Q}g(x)\,dx\\ &\qquad\qquad\geqslant\left(\frac{\kappa_{3}}{2}-\|g\|_{L^{\infty}(\mathbb{R}^{n})}\right)\min\{|(F-j)\cap Q|,|(F^{c}-j)\cap Q|\}\\ &\qquad\qquad\geqslant 0,\end{split}$
whenever $\|g\|_{L^{\infty}(\mathbb{R}^{n})}\leqslant\frac{\kappa_{3}}{2}$. This shows (7.5), as desired. ∎
8. Well-definedness of the stable norm $\phi$
In this section, we prove that the stable norm $\phi$ given in (1.21) is well-defined. The results presented below are the analogues in our setting of Lemma A.1 and Corollary A.2 in [chambolle_thouroude].
Proposition 8.1.
Let $K$ satisfy (1.2), (1.3), (1.4), and (1.5). Let $g\in L^{\infty}(Q)$ be a $\mathbb{Z}^{n}$-periodic function such that $\int_{Q}g=0$.
Assume that for any $p\in\mathbb{S}^{n-1}$ there exists a planelike minimizer $E_{p}$ for $\mathcal{J}$ in direction $p$ such that
(8.1) $\partial E_{p}\subseteq\{|x\cdot p|\leqslant M\},$
for some $M>0$ independent of $p$.
Then, there exists a function $\phi:\mathbb{S}^{n-1}\to\mathbb{R}$ such that, for any infinitesimal sequence $\{\varepsilon_{j}\}_{j}$ and any planelike minimizer $E^{\prime}_{p}$ for $\mathcal{J}$ in direction $p$ satisfying (8.1), it holds
(8.2) $\phi(p)=\lim_{j\to+\infty}\mathscr{F}_{\varepsilon_{j}}(\varepsilon_{j}E^{\prime}_{p},Q^{p})=\lim_{j\to+\infty}\varepsilon_{j}^{n-1}\mathscr{F}\left(E^{\prime}_{p},\frac{1}{\varepsilon_{j}}Q^{p}\right).$
In order to prove Proposition 8.1, we need the auxiliary results in Lemmata 8.2 and 8.3 here below.
Lemma 8.2.
Let $E\subseteq\mathbb{R}^{n}$ satisfy (8.1). Let also $\alpha\in(0,2s_{2}-1)$, where $s_{2}$ is as in (1.4).
Then,
(8.3) $\lim_{R\to+\infty}\frac{\mathscr{F}(E,A_{R,R^{1+\alpha},R})}{R^{n-1}}=0,$
where $A_{L_{1},L_{2},R}:=\{L_{1}<|x\cdot p|<L_{2},\ |x\cdot p^{\perp}|<R\}$.
Proof.
Let us call for simplicity
$A_{R,R^{1+\alpha},R}^{+}:=A_{R,R^{1+\alpha},R}\cap\{x\cdot p>0\}.$
Without loss of generality, we suppose that
(8.4) $\{x\cdot p>M\}\subseteq E.$
Let us also take $R$ so large that
$M\leqslant\frac{R}{2}.$
We claim that
(8.5) $E\cap A_{R,R^{1+\alpha},R}=A_{R,R^{1+\alpha},R}^{+}.$
To check this, we first show that
(8.6) $x\cdot p>R\qquad{\mbox{for every\penalty 10000\ $x\in E\cap A_{R,R^{1+\alpha},R}$.}}$
Indeed, if $x\in A_{R,R^{1+\alpha},R}$, we have that $|x\cdot p|>R$, namely either $x\cdot p<-R$ or $x\cdot p>R$.
Let now $x\in E\cap A_{R,R^{1+\alpha},R}$ and suppose that $x\cdot p<-R$. Then thanks to (8.1) and (8.4) we have that
$R>x\cdot p>-M\geqslant-\frac{R}{2},$
which gives the desired contradiction and establishes (8.6).
In light of (8.6), we see that
(8.7) $E\cap A_{R,R^{1+\alpha},R}\subseteq A_{R,R^{1+\alpha},R}^{+}.$
Furthermore, for every $x\in A_{R,R^{1+\alpha},R}^{+}$, we have that $x\cdot p>R>M$, which entails that $x\in E$, thanks to (8.4). This proves that
$A_{R,R^{1+\alpha},R}^{+}\subseteq E\cap A_{R,R^{1+\alpha},R}.$
The claim in (8.5) then follows from this and (8.7).
As a consequence of (8.5) we also have that
$E\cap\mathcal{Q}(A_{R,R^{1+\alpha},R})_{1}=\mathcal{Q}\big(A_{R,R^{1+\alpha},R}^{+}\big)_{1}.$
Hence, recalling also (1.1), we infer that
(8.8) $\int_{E\cap\mathcal{Q}(A_{R,R^{1+\alpha},R})_{1}}g(x)\,dx=\int_{\mathcal{Q}\big(A_{R,R^{1+\alpha},R}^{+}\big)_{1}}g(x)\,dx=0.$
Additionally, using (8.4) and (8.5), and recalling (1.4), we obtain
(8.9) $\begin{split}&\iint_{A_{R,R^{1+\alpha},R}\times\mathbb{R}^{n}}\chi_{E}(x)\chi_{E^{c}}(y)K(x,y)\,dx\,dy\leqslant\kappa_{2}\iint_{(E\cap A_{R,R^{1+\alpha},R})\times E^{c}}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}\\ &\quad\leqslant\kappa_{2}\iint_{A_{R,R^{1+\alpha},R}^{+}\times\{y\cdot p\leqslant M\}}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}\\ &\quad=\kappa_{2}\iint_{A_{R,R^{1+\alpha},R}^{+}\times(\{y\cdot p\leqslant M\}\cap B_{R})}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}+\kappa_{2}\iint_{A_{R,R^{1+\alpha},R}^{+}\times(\{y\cdot p\leqslant M\}\cap B_{R}^{c})}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}\\ &\quad=:\mathscr{I}_{1}+\mathscr{I}_{2},\end{split}$
and we aim to estimate $\mathscr{I}_{1}$ and $\mathscr{I}_{2}$.
First, notice that for every $x\in A_{R,R^{1+\alpha},R}^{+}$ and $y\in\{y\cdot p\leqslant M\}$, we have that $x\cdot p>R$ and
$|x-y|\geqslant(x-y)\cdot p=x\cdot p-y\cdot p>R-M\geqslant\frac{R}{2}.$
It thereby follows that
(8.10) $A_{R,R^{1+\alpha},R}^{+}\times\{y\cdot p\leqslant M\}\subseteq\left\{(x,y)\in\mathbb{R}^{n}\times\mathbb{R}^{n}\;:\;|x-y|>\frac{R}{2}\right\}.$
Using this, we see that
(8.11) $\begin{split}\mathscr{I}_{1}&\leqslant\kappa_{2}\iint_{\begin{subarray}{c}A_{R,R^{1+\alpha},R}^{+}\times(\{y\cdot p\leqslant M\}\cap B_{R})\\ \{|x-y|>R/2\}\end{subarray}}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}\\ &\leqslant\frac{2^{n+2s_{2}}\kappa_{2}|A_{R,R^{1+\alpha},R}|\,|\{y\cdot p\leqslant M\}\cap B_{R}|}{R^{n+2s_{2}}}\\ &\leqslant\frac{c(R^{1+\alpha}-R)R^{n-1}\,R^{n}}{R^{n+2s_{2}}}\\ &=cR^{n+\alpha-2s_{2}},\end{split}$
for some positive constant $c=c(n,s_{2},\kappa_{2},M)$, possibly changing at every step of the calculation.
Besides, exploiting again (8.10) and changing variable h:=yxh:=y-x,
$\begin{split}&\mathscr{I}_{2}\leqslant\kappa_{2}\iint_{\begin{subarray}{c}A_{R,R^{1+\alpha},R}^{+}\times(\{y\cdot p\leqslant M\}\cap B_{R}^{c})\\ \{|x-y|>R/2\}\end{subarray}}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}\leqslant\kappa_{2}|A_{R,R^{1+\alpha},R}|\int_{B_{R/2}^{c}}\frac{dh}{|h|^{n+2s_{2}}}\\ &\qquad\leqslant c(R^{1+\alpha}-R)R^{n-1}\int_{R/2}^{+\infty}\frac{d\rho}{\rho^{1+2s_{2}}}=cR^{n+\alpha-2s_{2}}.\end{split}$
Gathering this, (8.9), and (8.11) together, we deduce that
$\iint_{A_{R,R^{1+\alpha},R}\times\mathbb{R}^{n}}\chi_{E}(x)\chi_{E^{c}}(y)K(x,y)\,dx\,dy\leqslant cR^{n+\alpha-2s_{2}}.$
From this and (8.8), we conclude that
$\left|\frac{\mathscr{F}(E,A_{R,R^{1+\alpha},R})}{R^{n-1}}\right|=\frac{1}{R^{n-1}}\left|\operatorname{P}_{K}\big(E,A_{R,R^{1+\alpha},R}\big)+\int_{E\cap\mathcal{Q}(A_{R,R^{1+\alpha},R})_{1}}g(x)\,dx\right|\leqslant cR^{1+\alpha-2s_{2}},$
which entails (8.3). ∎
Lemma 8.3.
Let $E$ be a planelike minimizer for $\mathcal{J}$ satisfying (8.1).
Then, there exist $C_{1}$, $C_{2}\in\mathbb{R}$ independent of $\varepsilon$ such that
(8.12) $C_{1}\varepsilon^{1-n}\leqslant\mathscr{F}\left(E,\frac{1}{\varepsilon}Q^{p}\right)\leqslant C_{2}\varepsilon^{1-n}.$
Moreover, let $\eta\in(0,1)$ such that $\eta^{1+\alpha}<\varepsilon<\eta$, for some parameter $\alpha\in(0,2s_{2}-1)$.
Then, there exists a constant $c$ independent of $\varepsilon$ and $\eta$ such that
(8.13) $\mathscr{F}\left(E,\frac{1}{\varepsilon}Q^{p}\setminus\frac{1}{\eta}Q^{p}\right)\leqslant c\big(\varepsilon^{1-n}-\eta^{1-n}+\eta^{2s_{2}-1-\alpha}\big),$
whenever $\eta$ is small enough (and hence $\varepsilon$ is small enough).
Proof.
Since $E$ satisfies (8.1), we suppose, without loss of generality, that
(8.14) $\big\{x\cdot p>M\big\}\subseteq E.$
We will use the notation $\widetilde{Q}:=k+Q$, for some $k\in\mathbb{Z}^{n}$. Notice that $\mathcal{Q}(\widetilde{Q})_{1}=\widetilde{Q}$. Thus, by the class-$A$ minimality of $E$ (for $\mathcal{J}$), we find
$\mathscr{F}(E,\widetilde{Q})=\mathcal{J}(E,\widetilde{Q})\leqslant\mathcal{J}(E\setminus\widetilde{Q},\widetilde{Q})=P_{K}(E\setminus\widetilde{Q},\widetilde{Q})+\int_{(E\setminus\widetilde{Q})\cap\widetilde{Q}}g(x)\,dx\leqslant P_{K}(\widetilde{Q})=P_{K}(Q).$
Hence,
(8.15) $\operatorname{P}_{K}(E,\widetilde{Q})=\mathscr{F}(E,\widetilde{Q})-\int_{E\cap\widetilde{Q}}g(x)\,dx\leqslant\operatorname{P}_{K}(Q)+\|g\|_{L^{1}(Q)}.$
Now, let $S_{M}:=\{|x\cdot p|\leqslant M\}$ and take a (finite) covering of the strip $S_{M}\cap Q^{p}/\varepsilon$, namely
$\mathcal{B}:=\big\{k+Q\mbox{ s.t. }k\in\mathbb{Z}^{n},\ (k+Q)\cap(S_{M}\cap Q^{p}/\varepsilon)\neq\varnothing\big\}.$
Observe that $\sharp(\mathcal{B})\leqslant 2(M+\sqrt{n})\varepsilon^{1-n}$.
Then, using (8.15) together with the monotonicity and the subadditivity of $\operatorname{P}_{K}$ with respect to the domain, we infer that
(8.16) $\varepsilon^{n-1}\operatorname{P}_{K}\left(E,S_{M}\cap\frac{1}{\varepsilon}Q^{p}\right)\leqslant\varepsilon^{n-1}\sum_{\widetilde{Q}\in\mathcal{B}}\operatorname{P}_{K}(E,\widetilde{Q})\leqslant 2(M+\sqrt{n})\Big(\operatorname{P}_{K}(Q)+\|g\|_{L^{1}(Q)}\Big)$
and
(8.17) $\varepsilon^{n-1}\left|\int_{E\cap\mathcal{Q}(Q^{p}/\varepsilon)_{1}}g(x)\,dx\right|=\varepsilon^{n-1}\sum_{\widetilde{Q}\in\mathcal{B}}\left|\int_{E\cap\mathcal{Q}(Q^{p}/\varepsilon)_{1}\cap\widetilde{Q}}g(x)\,dx\right|\leqslant 2(M+\sqrt{n})\|g\|_{L^{1}(Q)}.$
In particular, (8.17) yields
(8.18) $\varepsilon^{n-1}\mathscr{F}\left(E,\frac{1}{\varepsilon}Q^{p}\right)\geqslant\varepsilon^{n-1}\int_{E\cap\mathcal{Q}(Q^{p}/\varepsilon)_{1}}g(x)\,dx\geqslant-2(M+\sqrt{n})\|g\|_{L^{1}(Q)}.$
Moreover, we claim that
(8.19) $\varepsilon^{n-1}\operatorname{P}_{K}\left(E,\frac{1}{\varepsilon}Q^{p}\setminus S_{M}\right)\leqslant c.$
To achieve this, let $m\in\mathbb{N}$ be such that $2^{-m}\leqslant\varepsilon\leqslant 2^{-m+1}$. We define, for every $j\in\{0,\dots,m-1\}$,
$C_{j}:=\frac{1}{\varepsilon}Q^{p}\cap\Big\{M2^{j}\leqslant|x\cdot p|\leqslant M2^{j+1}\Big\}\subseteq\mathbb{R}^{n}.$
We recall (8.14) and we also set
$S_{M}^{+}:=\big\{x\cdot p>M\big\}\subseteq E$
and $S_{M}^{-}:=\big\{x\cdot p<-M\big\}\subseteq E^{c}.$
Then, we have
$\frac{1}{\varepsilon}Q^{p}\setminus S_{M}\subseteq\bigcup_{j=0}^{m-1}C_{j}.$
Moreover, notice that every $C_{j}$ can be covered with at most $M2^{m(n-1)+j+1}$ cubes of side $1$ centered at points in $\mathbb{Z}^{n}$ (let us call $\mathcal{B}_{j}$ such covering).
Let $\widetilde{Q}:=k+Q\in\mathcal{B}_{j}$, for some $k\in\mathbb{Z}^{n}$ and some $j\in\{0,\dots,m-1\}$. Without loss of generality, we can suppose that $\widetilde{Q}\cap S_{M}^{+}\neq\varnothing$, and hence $\widetilde{Q}\cap S_{M}^{-}=\varnothing$. Notice also that $\widetilde{Q}^{c}\cap S_{M}^{-}\subseteq{B}_{(2^{j}+1)M}^{c}(k)$. Thus,
(8.20) $\begin{split}&\iint_{\widetilde{Q}_{\sharp}}\chi_{S_{M}^{+}}(x)\chi_{S_{M}^{-}}(y)K(x,y)\,dx\,dy=\iint_{\widetilde{Q}\times(\widetilde{Q}^{c}\cap S_{M}^{-})}K(x,y)\,dx\,dy\\ &\qquad\qquad\leqslant\kappa_{2}\iint_{\widetilde{Q}\times(\widetilde{Q}^{c}\cap S_{M}^{-})}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}\leqslant\kappa_{2}\iint_{\widetilde{Q}\times{B}_{(2^{j}+1)M}^{c}(k)}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}\\ &\qquad\qquad=\kappa_{2}\iint_{Q\times{B}_{(2^{j}+1)M}^{c}}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}=\kappa_{2}\iint_{Q\times{B}_{(2^{j}+1)M}^{c}(x)}\frac{dx\,dh}{|h|^{n+2s_{2}}}\\ &\qquad\qquad\leqslant\kappa_{2}|Q|\int_{{B}_{2^{j}M}^{c}}\frac{dh}{|h|^{n+2s_{2}}}\leqslant c_{1}\,2^{-2s_{2}j},\end{split}$
for some positive $c_{1}=c_{1}(n,s_{2},\kappa_{2},M)$ independent of $m$.
Now, observe that, by (8.14) and the definitions of $S_{M}^{+}$ and $S_{M}^{-}$, we have that
$E\cap S_{M}^{c}=S_{M}^{+}\qquad\mbox{and}\qquad E^{c}\cap S_{M}^{c}=S_{M}^{-}.$
As a consequence,
$\begin{split}&\operatorname{P}_{K}\left(E,\frac{1}{\varepsilon}Q^{p}\setminus S_{M}\right)=\iint_{(Q^{p}/\varepsilon\setminus S_{M})_{\sharp}}\chi_{E}(x)\chi_{E^{c}}(y)K(x,y)\,dx\,dy\\ =&\iint_{(Q^{p}/\varepsilon\setminus S_{M})\times(Q^{p}/\varepsilon\setminus S_{M})}\chi_{E}(x)\chi_{E^{c}}(y)K(x,y)\,dx\,dy\\ &+\iint_{((Q^{p}/\varepsilon)^{c}\setminus S_{M})\times(Q^{p}/\varepsilon\setminus S_{M})}\chi_{E}(x)\chi_{E^{c}}(y)K(x,y)\,dx\,dy\\ &+\iint_{(Q^{p}/\varepsilon\setminus S_{M})\times((Q^{p}/\varepsilon)^{c}\setminus S_{M})}\chi_{E}(x)\chi_{E^{c}}(y)K(x,y)\,dx\,dy\\ &+\iint_{S_{M}\times(Q^{p}/\varepsilon\setminus S_{M})}\chi_{E}(x)\chi_{E^{c}}(y)K(x,y)\,dx\,dy+\iint_{(Q^{p}/\varepsilon\setminus S_{M})\times S_{M}}\chi_{E}(x)\chi_{E^{c}}(y)K(x,y)\,dx\,dy\\ =&\iint_{(Q^{p}/\varepsilon\setminus S_{M})_{\sharp}}\chi_{S_{M}^{+}}(x)\chi_{S_{M}^{-}}(y)K(x,y)\,dx\,dy\\ &+\iint_{S_{M}\times(Q^{p}/\varepsilon\setminus S_{M})}\chi_{E}(x)\chi_{S_{M}^{-}}(y)K(x,y)\,dx\,dy+\iint_{(Q^{p}/\varepsilon\setminus S_{M})\times S_{M}}\chi_{S_{M}^{+}}(x)\chi_{E^{c}}(y)K(x,y)\,dx\,dy\\ \leqslant&\iint_{(Q^{p}/\varepsilon\setminus S_{M})_{\sharp}}\chi_{S_{M}^{+}}(x)\chi_{S_{M}^{-}}(y)K(x,y)\,dx\,dy+\iint_{S_{M}\times(Q^{p}/\varepsilon\setminus S_{M})}K(x,y)\,dx\,dy.\end{split}$
Moreover, by Proposition C.2, there exists a positive constant $c_{2}$, independent of $\varepsilon$, such that
(8.21) $\iint_{S_{M}\times(Q^{p}/\varepsilon\setminus S_{M})}K(x,y)\,dx\,dy\leqslant\operatorname{P}_{K}(Q^{p}/\varepsilon\cap S_{M}^{+})+\operatorname{P}_{K}(Q^{p}/\varepsilon\cap S_{M}^{-})\leqslant c_{2}\varepsilon^{1-n}.$
Thus, the estimate in (8.20), in combination with (8.21) and the subadditivity of $P_{K}$ with respect to the domain, entails that
$\begin{split}&\varepsilon^{n-1}\operatorname{P}_{K}\left(E,\frac{1}{\varepsilon}Q^{p}\setminus S_{M}\right)\leqslant\varepsilon^{n-1}\left(\iint_{(Q^{p}/\varepsilon\setminus S_{M})_{\sharp}}\chi_{S_{M}^{+}}(x)\chi_{S_{M}^{-}}(y)K(x,y)\,dx\,dy+c_{2}\varepsilon^{1-n}\right)\\ &\qquad\leqslant 2^{-(m-1)(n-1)}\iint_{(Q^{p}/\varepsilon\setminus S_{M})_{\sharp}}\chi_{S_{M}^{+}}(x)\chi_{S_{M}^{-}}(y)K(x,y)\,dx\,dy+c_{2}\\ &\qquad\leqslant 2^{-(m-1)(n-1)}\sum_{j=0}^{m}\sum_{\widetilde{Q}\in\mathcal{B}_{j}}\iint_{\widetilde{Q}_{\sharp}}\chi_{S_{M}^{+}}(x)\chi_{S_{M}^{-}}(y)K(x,y)\,dx\,dy+c_{2}\\ &\qquad\leqslant 2^{-(m-1)(n-1)}\sum_{j=0}^{m}c_{1}2^{m(n-1)+j}2^{-2s_{2}j}\ +c_{2}\leqslant c_{1}\sum_{j=0}^{\infty}2^{(1-2s_{2})j}\ +c_{2}=:c<+\infty,\end{split}$
showing (8.19).
Therefore, gathering (8.16), (8.18) and (8.19), we obtain
$2(M+\sqrt{n})\|g\|_{L^{1}(Q)}\leqslant\varepsilon^{n-1}\mathscr{F}\left(E,\frac{1}{\varepsilon}Q^{p}\right)\leqslant 2(M+\sqrt{n})\Big(\operatorname{P}_{K}(Q)+2\|g\|_{L^{1}(Q)}\Big)+c,$
showing (8.12).
Now, to prove (8.13), let us consider $\eta^{1+\alpha}<\varepsilon<\eta$, for some $\alpha\in(0,2s_{2}-1)$. We assume again that $2^{-m}\leqslant\varepsilon\leqslant 2^{-m+1}$, for some $m\in\mathbb{N}$, and we decompose $\frac{1}{\varepsilon}Q^{p}\setminus\frac{1}{\eta}Q^{p}$ as
$\frac{1}{\varepsilon}Q^{p}\setminus\frac{1}{\eta}Q^{p}=\left(\frac{1}{\varepsilon}Q^{p}\setminus\frac{1}{\eta}Q^{p}\right)\cap S_{M}\cup\bigcup_{j=0}^{m-1}\widetilde{C}_{j},$
where, for all $j\in\{0,\dots,m-1\}$,
$\widetilde{C}_{j}:=\left(\frac{1}{\varepsilon}Q^{p}\setminus\frac{1}{\eta}Q^{p}\right)\cap\Big\{M2^{j}\leqslant|x\cdot p|\leqslant M2^{j+1}\Big\}.$
Notice that $\left(\frac{1}{\varepsilon}Q^{p}\setminus\frac{1}{\eta}Q^{p}\right)\cap S_{M}$ can be covered with at most $2M(\varepsilon^{1-n}-\eta^{1-n})$ cubes of side $1$. Similarly, each cylinder $\widetilde{C}_{j}$ can be covered with at most $2^{j}(\varepsilon^{1-n}-\eta^{1-n})$ translated unit cubes.
Therefore, arguing as for (8.12) and recalling Lemma 8.2, we deduce that, for every $\eta$ small enough,
$\mathscr{F}\left(E,\frac{1}{\varepsilon}Q^{p}\setminus\frac{1}{\eta}Q^{p}\right)\leqslant c(\varepsilon^{1-n}-\eta^{1-n}+\eta^{2s_{2}-1-\alpha}),$
for some positive constant $c$ independent of $\eta$ and $\varepsilon$, concluding the proof. ∎
Proof of Proposition 8.1.
Let $p\in\mathbb{S}^{n-1}$. In order to prove that the limit function $\phi$ in (8.2) is well-defined, we show that if $E_{p}$ and $E^{\prime}_{p}$ are two planelike minimizers for $\mathcal{J}$ with respect to some $p$ such that
$\partial E_{p}\cap\partial E^{\prime}_{p}\subseteq\{|x\cdot p|\leqslant M\},$
and $\varepsilon$, $\varepsilon^{\prime}>0$ are infinitesimal sequences such that the limits
$\ell:=\lim_{\varepsilon\to 0}\varepsilon^{n-1}\mathscr{F}\left(E_{p},\frac{1}{\varepsilon}Q^{p}\right)\qquad\mbox{and}\qquad\ell^{\prime}:=\lim_{\varepsilon^{\prime}\to 0}(\varepsilon^{\prime})^{n-1}\mathscr{F}\left(E^{\prime}_{p},\frac{1}{\varepsilon^{\prime}}Q^{p}\right)$
exist, then it must be $\ell=\ell^{\prime}$. To this purpose, up to taking subsequences, let us assume that $\varepsilon^{\prime}\ll\varepsilon$ (e.g. $\varepsilon^{\prime}\sim\varepsilon^{1+\alpha}$, for some $\alpha\in(0,2s_{2}-1)$).
Following the footsteps of [chambolle_thouroude], we recall the notation for $\mathcal{I}_{p}$ in (7.1) and we consider a finite covering of $\mathcal{I}_{p}\cap\mathcal{Q}(Q^{p}/\varepsilon^{\prime})_{1}$ made of $N:=\left\lfloor\left(\frac{\varepsilon}{\varepsilon^{\prime}(1+2\varepsilon\sqrt{n})}\right)^{n-1}\right\rfloor$ cubes that are translations of $(2\sqrt{n}+1/\varepsilon)Q^{p}$ centered at $\mathcal{I}_{p}$. Notice that each of this cube contains at least one translated cube $Q^{(j)}:=j+Q^{p}/\varepsilon$, with $j\in\mathbb{Z}^{n}$. Let also
$E_{j}:=j+\left(E_{p}\cap\frac{1}{\varepsilon}Q^{p}\right)\subseteq Q^{(j)}.$
Also, we define
(8.22) $R:=\bigcup_{j=1}^{N}Q^{(j)}\subseteq\mathcal{Q}(Q^{p}/\varepsilon^{\prime})_{1}\qquad\mbox{and}\qquad F^{\prime}:=\left(\bigcup_{j=1}^{N}E_{j}\right)\cup(E^{\prime}_{p}\setminus R).$
In particular, we have $F^{\prime}\setminus\mathcal{Q}(Q^{p}/\varepsilon^{\prime})_{1}=E^{\prime}_{p}\setminus\mathcal{Q}(Q^{p}/\varepsilon^{\prime})_{1}$ (see Figure 1).
Figure 1. The construction of the set $F^{\prime}$ as in (8.22).
Thus, by Remark 1.8, the subadditivity of $\operatorname{P}_{K}$ with respect to the domain, and (1.2), we infer that
(8.23) $\begin{split}&\mathscr{F}\left(E^{\prime}_{p},\frac{1}{\varepsilon^{\prime}}Q^{p}\right)\leqslant\mathscr{F}\left(F^{\prime},\frac{1}{\varepsilon^{\prime}}Q^{p}\right)=\mathscr{F}\left(F^{\prime},R\cup\left(\frac{1}{\varepsilon^{\prime}}Q^{p}\setminus R\right)\right)\\ &\qquad\leqslant\sum_{j=1}^{N}\mathscr{F}\left(E_{j},Q^{(j)}\right)+\iint_{(Q^{p}/\varepsilon^{\prime}\setminus R)_{\sharp}}\chi_{F^{\prime}}(x)\chi_{(F^{\prime})^{c}}(y)K(x,y)\,dx\,dy\\ &\qquad\qquad+\int_{F^{\prime}\cap(\mathcal{Q}(Q^{p}/\varepsilon^{\prime})_{1}\setminus\bigcup_{j=1}^{N}\mathcal{Q}(Q^{(j)})_{1})}g(x)\,d\\ &\qquad\leqslant N\mathscr{F}\left(E_{p},\frac{1}{\varepsilon}Q^{p}\right)+\iint_{(Q^{p}/\varepsilon^{\prime}\setminus R)_{\sharp}}\chi_{F^{\prime}}(x)\chi_{(F^{\prime})^{c}}(y)K(x,y)\,dx\,dy\\ &\qquad\qquad+\int_{F^{\prime}\cap(\mathcal{Q}(Q^{p}/\varepsilon^{\prime})_{1}\setminus\bigcup_{j=1}^{N}\mathcal{Q}(Q^{(j)})_{1})}g(x)\,dx.\end{split}$
At this stage, observe that
$\partial F^{\prime}\cap(Q^{p}/\varepsilon^{\prime}\setminus R)\subseteq\left\{|x\cdot p|\leqslant M+2\sqrt{n}\right\}$
can be covered with $\left\lfloor(M+2\sqrt{n})\left((\varepsilon^{\prime})^{1-n}-N\varepsilon^{1-n}\right)\right\rfloor$ cubes $Q$ of side $1$ and centered at $\mathbb{Z}^{n}$. Therefore, it follows from (8.13) that
(8.24) $\begin{split}&\iint_{(Q^{p}/\varepsilon^{\prime}\setminus R)_{\sharp}}\chi_{F^{\prime}}(x)\chi_{(F^{\prime})^{c}}(y)K(x,y)\,dx\,dy\\ \leqslant\;&c(M+2\sqrt{n})\Big((\varepsilon^{\prime})^{1-n}-N\varepsilon^{1-n}+\varepsilon^{2s_{2}-1-\alpha}\Big)\\ \leqslant\;&\frac{c(M+2\sqrt{n})}{(\varepsilon^{\prime})^{n-1}}\left(1-\left(\frac{\varepsilon^{\prime}}{\varepsilon}\right)^{n-1}\left\lfloor\left(\frac{\varepsilon}{\varepsilon^{\prime}(1+2\varepsilon\sqrt{n})}\right)^{n-1}\right\rfloor\right)+c(M+2\sqrt{n})\varepsilon^{2s_{2}-1-\alpha}\\ =:\;&\frac{\xi(\varepsilon,\varepsilon^{\prime})}{(\varepsilon^{\prime})^{n-1}},\end{split}$
for some function $\xi$ such that $\xi(\varepsilon,\varepsilon^{\prime})\to 0$, as $\varepsilon$, $\varepsilon^{\prime}$, and $\varepsilon^{\prime}/\varepsilon\sim\varepsilon^{\alpha}\to 0$.
Furthermore, since
$\partial F^{\prime}\cap(\mathcal{Q}(Q^{P}/\varepsilon^{\prime})_{1}\setminus\bigcup_{j=1}^{N}\mathcal{Q}(Q^{(j)})_{1}))\subseteq\{|x\cdot p|\leqslant M+2\sqrt{n}\}$
can be covered with $\left\lfloor(M+2\sqrt{n})\left((\varepsilon^{\prime})^{1-n}-N\lfloor\varepsilon^{1-n}\rfloor\right)\right\rfloor$ cubes $Q^{p}$ of side $1$ and centered at $\mathbb{Z}^{n}$, we also have
(8.25) $\begin{split}&\int_{F^{\prime}\cap(\mathcal{Q}(Q^{P}/\varepsilon^{\prime})_{1}\setminus\bigcup_{j=1}^{N}\mathcal{Q}(Q^{(j)})_{1}))}g(x)\,dx\leqslant\|g\|_{L^{1}(Q)}(M+2\sqrt{n})\Big((\varepsilon^{\prime})^{1-n}-N\lfloor\varepsilon^{1-n}\rfloor\Big)\\ &\qquad=\frac{\|g\|_{L^{1}(Q)}(M+2\sqrt{n})}{(\varepsilon^{\prime})^{n-1}}\left(1-(\varepsilon^{\prime})^{n-1}\lfloor\varepsilon^{1-n}\rfloor\left\lfloor\left(\frac{\varepsilon}{\varepsilon^{\prime}(1+2\varepsilon\sqrt{n})}\right)^{n-1}\right\rfloor\right)=:\frac{\zeta(\varepsilon,\varepsilon^{\prime})}{(\varepsilon^{\prime})^{n-1}},\end{split}$
for some function $\zeta$ such that $\zeta(\varepsilon,\varepsilon^{\prime})\to 0$, as $\varepsilon\to 0$ (and also $\varepsilon^{\prime},\varepsilon^{\prime}/\varepsilon\to 0$). Indeed,
$\begin{split}&\lim_{\begin{subarray}{c}\varepsilon^{\prime}\to 0\\ \varepsilon\to 0\end{subarray}}\ (\varepsilon^{\prime})^{n-1}\lfloor\varepsilon^{1-n}\rfloor\left\lfloor\left(\frac{\varepsilon}{\varepsilon^{\prime}(1+2\varepsilon\sqrt{n})}\right)^{n-1}\right\rfloor\\ =\;&\lim_{\begin{subarray}{c}\varepsilon^{\prime}\to 0\\ \varepsilon\to 0\end{subarray}}\ (\varepsilon^{\prime})^{n-1}\big(\varepsilon^{1-n}-\{\varepsilon^{1-n}\}\big)\left(\left(\frac{\varepsilon}{\varepsilon^{\prime}(1+2\varepsilon\sqrt{n})}\right)^{n-1}-\left\{\left(\frac{\varepsilon}{\varepsilon^{\prime}(1+2\varepsilon\sqrt{n})}\right)^{n-1}\right\}\right)\\ =\;&1,\end{split}$
where $\{\cdot\}:=(\cdot)$-$\lfloor\cdot\rfloor$ denotes the fractional part of a real number.
Now, recalling the definitions of $\ell$ and $\ell^{\prime}$, and putting (8.23), (8.24) and (8.25) together, we obtain
$(\varepsilon^{\prime})^{n-1}\mathscr{F}\left(E^{\prime}_{p},\frac{1}{\varepsilon^{\prime}}Q^{p}\right)\leqslant(\varepsilon^{\prime})^{n-1}\left\lfloor\left(\frac{\varepsilon}{\varepsilon^{\prime}(1+2\varepsilon\sqrt{n})}\right)^{n-1}\right\rfloor\mathscr{F}\left(E_{p},\frac{1}{\varepsilon}Q^{p}\right)+\xi(\varepsilon,\varepsilon^{\prime})+\zeta(\varepsilon,\varepsilon^{\prime}).$
Hence, taking the limits for $\varepsilon$, $\varepsilon^{\prime}$, and $\varepsilon^{\prime}/\varepsilon\sim\varepsilon^{\alpha}\to 0$, we deduce that $\ell^{\prime}\leqslant\ell$.
Interchanging the roles of $\varepsilon$ and $\varepsilon^{\prime}$, we also get that $\ell\leqslant\ell^{\prime}$, concluding the proof. ∎
We stress that the construction of the cubes $Q^{(j)}$ in the proof of Proposition 8.1 remains unaltered if we replace $Q^{p}/\varepsilon^{\prime}$ with one of its translated along $\partial\mathcal{I}_{p}$. Indeed, notice that when we consider translations of the cube $(2\sqrt{n}+1/\varepsilon)Q^{p}$, we do not ask any requirement on those translations but the fact that they are centered at $\partial\mathcal{I}_{p}$. Therefore, we infer that $\phi(p)$ is independent of translations of the domain in directions orthogonal to $p$.
The precise statement goes as follows:
Corollary 8.4.
Let $K$ satisfy (1.2), (1.3), (1.4), and (1.5). Let $g\in L^{\infty}(Q)$ be a $\mathbb{Z}^{n}$-periodic function such that $\int_{Q}g=0$.
Let $p\in\mathbb{S}^{n-1}$ and let $E_{p}$ be a planelike minimizer for $\mathcal{J}$ in direction $p$ that satisfies (8.1) . Let also $v\in\partial\mathcal{I}_{p}$, i.e. $v\cdot p=0$.
Then,
$\phi(p)=\lim_{\varepsilon\to 0}\varepsilon^{n-1}\mathscr{F}\left(E_{p},v+\frac{1}{\varepsilon}Q^{p}\right).$
As a consequence, we deduce the following:
Corollary 8.5.
Let $K$ satisfy (1.2), (1.3), (1.4), and (1.5). Let $g\in L^{\infty}(Q)$ be a $\mathbb{Z}^{n}$-periodic function such that $\int_{Q}g=0$.
Let $p\in\mathbb{S}^{n-1}$ and $t\in\mathcal{T}_{p}$, so that $E_{p,t}$ is a planelike minimizer for $\mathcal{J}$ in direction $p$ that satisfies (8.1) (according to Corollary 1.12).
Then, for any Lipschitz domain $\Omega\subseteq\mathbb{R}^{n}$,
(8.26) $\phi(p)\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p})=\lim_{\varepsilon\to 0}\mathscr{F}_{\varepsilon}(\varepsilon E_{p,t},\Omega),$
where $\mathcal{I}_{p}$ is defined as in (7.1).
Proof.
Let us consider sequences $\varepsilon_{j}$, $\rho_{j}$ such that
(8.27) $\lim_{j\to+\infty}\varepsilon_{j}=\lim_{j\to+\infty}\rho_{j}=\lim_{j\to+\infty}\frac{\varepsilon_{j}}{\rho_{j}}=0.$
Also, we consider the collection
$\mathcal{B}_{j}:=\{\rho_{j}\widetilde{Q}:=\varepsilon_{j}v_{j}+\rho_{j}Q^{p}\mbox{ s.t. }v_{j}\in\partial\mathcal{I}_{p}\mbox{, and }\rho_{j}\widetilde{Q}\subseteq\Omega\}.$
In particular, $N_{j}:=\sharp(\mathcal{B}_{j})=\lfloor\rho_{j}^{1-n}\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p})\rfloor$. Let us also denote
$R_{j}:=\bigcup_{\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}}\rho_{j}\widetilde{Q}$.
It is convenient to look at the family of cubes
$\mathcal{Q}_{E_{p}}(\Omega/\varepsilon_{j})_{1}:=\bigcup\big(\ell+Q\big)$
where the union is taken over all $\ell\in\mathbb{Z}^{n}$ such that $\ell+Q\subset\Omega/\varepsilon_{j}$ and $(\ell+Q)\cap\partial E_{p}\neq\varnothing$.
Thanks to (8.27), we can take $j$ sufficiently large such that $\rho_{j}/\varepsilon_{j}>M$, and we have that
(8.28) $\begin{split}&\varepsilon_{j}^{n-1}\int_{E_{p}\cap\mathcal{Q}(\Omega/\varepsilon)_{1}}g=\varepsilon_{j}^{n-1}\int_{E_{p}\cap\mathcal{Q}_{E_{p}}(\Omega/\varepsilon_{j})_{1}}g\\ &\qquad=\sum_{\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}}\left(\varepsilon_{j}^{n-1}\int_{E_{p}\cap\mathcal{Q}_{E_{p}}(\rho_{j}\widetilde{Q}/\varepsilon_{j})_{1}}g\right)+\varepsilon_{j}^{n-1}\int_{E_{p}\cap\left(\mathcal{Q}_{E_{p}}(\Omega/\varepsilon_{j})\setminus\bigcup_{\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}}\mathcal{Q}_{E_{p}}(\rho_{j}\widetilde{Q}/\varepsilon_{j})_{1}\right)}g\\ &\qquad=\sum_{\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}}\left(\varepsilon_{j}^{n-1}\int_{E_{p}\cap\mathcal{Q}(\rho_{j}\widetilde{Q}/\varepsilon_{j})_{1}}g\right)+\psi(j),\end{split}$
where
$\psi(j):=\varepsilon_{j}^{n-1}\int_{E_{p}\cap\left(\mathcal{Q}_{E_{p}}(\Omega/\varepsilon_{j})_{1}\setminus\bigcup_{\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}}\mathcal{Q}_{E_{p}}(\rho_{j}\widetilde{Q}/\varepsilon_{j})_{1}\right)}g.$
Now, let $\eta_{j}:=\varepsilon_{j}/\rho_{j}$, and define
$M_{\Omega,j}:=\sharp\left\{\ell+Q\subset\Omega/\varepsilon_{j}\mbox{ s.t. }\ell\in\mathbb{Z}^{n}\mbox{ and }(\ell+Q)\cap\partial E_{p}\neq\varnothing\right\}$
and $M_{\widetilde{Q},j}:=\sharp\left\{\ell+Q\subset\widetilde{Q}/\eta_{j}\mbox{ s.t. }\ell\in\mathbb{Z}^{n}\mbox{ and }(\ell+Q)\cap\partial E_{p}\neq\varnothing\right\},\quad\mbox{for every }\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}.$
Since $E_{p}$ is a planelike in direction $p$, we have that
(8.29) $\begin{split}&\lim_{j\to+\infty}\varepsilon^{n-1}M_{\Omega,j}=\lim_{j\to+\infty}\sharp\left\{\varepsilon_{j}(\ell+Q)\subset\Omega\mbox{ s.t. }\ell\in\mathbb{Z}^{n}\mbox{ and }\varepsilon_{j}(\ell+Q)\cap\partial\varepsilon_{j}E_{p}\neq\varnothing\right\}\\ &\qquad=\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p})\end{split}$
and, similarly, for every $\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}$,
(8.30) $\begin{split}&\lim_{j\to+\infty}\eta^{n-1}M_{\widetilde{Q},j}=\lim_{j\to+\infty}\sharp\left\{\eta_{j}(\ell+Q)\subset\widetilde{Q}\mbox{ s.t. }\ell\in\mathbb{Z}^{n}\mbox{ and }\eta_{j}(\ell+Q)\cap\partial(\eta_{j}E_{p})\neq\varnothing\right\}\\ &\qquad=\mathcal{H}^{n-1}(\widetilde{Q}\cap\partial\mathcal{I}_{p})=1.\end{split}$
Moreover, recalling that the fractional part of a real number is a bounded function, we deduce that
(8.31) $\lim_{j\to+\infty}\rho_{j}^{n-1}N_{j}=\lim_{j\to+\infty}\rho_{j}^{n-1}\left(\rho_{j}^{1-n}\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p})-\{\rho_{j}^{1-n}\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p}\}\right)=\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p}).$
Thus, from (8.29), (8.30), and (8.31) together, we infer that
(8.32) $\begin{split}&|\psi(j)|\leqslant\varepsilon_{j}^{n-1}\|g\|_{L^{1}(Q)}\left[M_{\Omega,j}-\sum_{\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}}M_{\tilde{Q},j}\right]\\ &\qquad=\|g\|_{L^{1}(Q)}\left[\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p})+o(1)-\sum_{\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}}\left(\rho^{n-1}+o(\rho^{n-1})\right)\right]\\ &\qquad=\|g\|_{L^{1}(Q)}\left[\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p})+o(1)-N_{j}\left(\rho^{n-1}+o(\rho^{n-1})\right)\right]\xrightarrow[j\to\infty]{}0.\end{split}$
Therefore, from (8.28) and recalling the definition of $\mathscr{E}$ from (1.20), we obtain that
$\begin{split}\varepsilon_{j}^{n-1}\mathscr{F}(E_{p},\Omega/\varepsilon_{j})&=\varepsilon_{j}^{n-1}\mathscr{E}(E_{p},\Omega/\varepsilon_{j})+\varepsilon_{j}^{n-1}\mathcal{L}_{K}(E_{p}\cap\Omega/\varepsilon_{j},E_{p}^{c}\cap\Omega^{c}/\varepsilon_{j})\\ &=\sum_{\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}}\left[\varepsilon_{j}^{n-1}\mathscr{E}\left(E_{p},\frac{\rho_{j}}{\varepsilon_{j}}\widetilde{Q}\right)\right]+\psi(j)+\varepsilon_{j}^{n-1}\operatorname{P}_{K}(E_{p},(\Omega\setminus R_{j})/\varepsilon_{j})\\ &\quad-\varepsilon_{j}^{n-1}\mathcal{L}_{K}(E_{p}\cap(\Omega\setminus R_{j})^{c}/\varepsilon_{j},E_{p}^{c}\cap(\Omega\setminus R_{j})/\varepsilon_{j})\\ &\quad+\varepsilon_{j}^{n-1}\mathcal{L}_{K}(E_{p}\cap\Omega/\varepsilon_{j},E_{p}^{c}\cap\Omega^{c}/\varepsilon_{j})\\ &=\sum_{\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}}\left[\varepsilon_{j}^{n-1}\mathscr{F}\left(E_{p},\frac{\rho_{j}}{\varepsilon_{j}}\widetilde{Q}\right)-\varepsilon_{j}^{n-1}\mathcal{L}_{K}\left(E_{p}\cap\frac{\rho_{j}}{\varepsilon_{j}}\widetilde{Q},E_{p}^{c}\cap\frac{\rho_{j}}{\varepsilon_{j}}\widetilde{Q}^{c}\right)\right]\\ &\quad+\psi(j)+\varepsilon_{j}^{n-1}\operatorname{P}_{K}(E_{p},(\Omega\setminus R_{j})/\varepsilon_{j})\\ &\quad-\varepsilon_{j}^{n-1}\mathcal{L}_{K}(E_{p}\cap(\Omega\setminus R_{j})^{c}/\varepsilon_{j},E_{p}^{c}\cap(\Omega\setminus R_{j})/\varepsilon_{j})\\ &\quad+\varepsilon_{j}^{n-1}\mathcal{L}_{K}(E_{p}\cap\Omega/\varepsilon_{j},E_{p}^{c}\cap\Omega^{c}/\varepsilon_{j}).\end{split}$
At this point, observe that it follows from Corollary 8.4, Lemma 7.1, and our choice of $\rho_{j}\widetilde{Q}$ that
$\varepsilon_{j}^{n-1}\mathscr{F}\left(E_{p},\frac{\rho_{j}}{\varepsilon_{j}}\widetilde{Q}\right)-\varepsilon_{j}^{n-1}\mathcal{L}_{K}\left(E_{p}\cap\frac{\rho_{j}}{\varepsilon_{j}}\widetilde{Q},E_{p}^{c}\cap\frac{\rho_{j}}{\varepsilon_{j}}\widetilde{Q}^{c}\right)=\rho_{j}^{n-1}\phi(p)+o(\rho_{j}^{n-1}).$
Thus, we infer that
(8.33) $\begin{split}&\varepsilon_{j}^{n-1}\mathscr{F}(E_{p},\Omega/\varepsilon_{j})\\ =&N_{j}\left(\rho_{j}^{n-1}\phi(p)+o(\rho_{j}^{n-1})\right)+\psi(j)+\varepsilon_{j}^{n-1}\operatorname{P}_{K}(E_{p},(\Omega\setminus R_{j})/\varepsilon_{j})\\ &\quad-\varepsilon_{j}^{n-1}\mathcal{L}_{K}(E_{p}\cap(\Omega\setminus R_{j})^{c}/\varepsilon_{j},E_{p}^{c}\cap(\Omega\setminus R_{j})/\varepsilon_{j})\\ &\quad+\varepsilon_{j}^{n-1}\mathcal{L}_{K}(E_{p}\cap\Omega/\varepsilon_{j},E_{p}^{c}\cap\Omega^{c}/\varepsilon_{j}).\end{split}$
Furthermore, by inspection of the proof of (8.13) in Lemma 8.3, we infer that
(8.34) $\begin{split}&\lim_{j\to+\infty}\varepsilon_{j}^{n-1}\operatorname{P}_{K}\left(E_{p},\Omega/\varepsilon_{j}\setminus\left(\bigcup_{\rho_{j}\widetilde{Q}\in\mathcal{B}_{j}}\frac{\rho_{j}}{\varepsilon_{j}}\widetilde{Q}\right)\right)\\ &\qquad\leqslant\lim_{j\to+\infty}c(\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p})-N_{j}\rho^{n-1}+\varepsilon_{j}^{n-1}(N_{j}\rho_{j}\varepsilon_{j}^{-1})^{2s_{2}-1-\alpha})=0.\end{split}$
Additionally, thanks to Lemma 7.1, we also have
(8.35) $\limsup_{j\to+\infty}\,\varepsilon_{j}^{n-1}\mathcal{L}_{K}(E_{p}\cap(\Omega\setminus R_{j})^{c}/\varepsilon_{j},E_{p}^{c}\cap(\Omega\setminus R_{j})/\varepsilon_{j})+\varepsilon_{j}^{n-1}\mathcal{L}_{K}(E_{p}\cap\Omega/\varepsilon_{j},E_{p}^{c}\cap\Omega^{c}/\varepsilon_{j})=0.$
Using (8.31), (8.32), (8.34), and (8.35) in (8.33), we conclude that
$\lim_{j\to+\infty}\varepsilon_{j}^{n-1}\mathscr{F}(E_{p},\Omega/\varepsilon_{j})=\mathcal{H}^{n-1}(\Omega\cap\partial\mathcal{I}_{p})\phi(p).\qed$
9. Proof of the $\Gamma$-$\liminf$ inequality
Here, we present a proof of the $\Gamma-\liminf$ inequality in Theorem 1.13-(i). To this purpose, let $E\subseteq\mathbb{R}^{n}$, and let $\{E_{\varepsilon}\}_{\varepsilon}$ be a sequence of sets such that $E_{\varepsilon}\to E$ in $L^{1}_{\rm{loc}}(\mathbb{R}^{n})$.
Moreover, up to extracting a subsequence, we suppose that
[…]
otherwise the result is trivial. In particular, without loss of generality, we assume that
$\sup_{\varepsilon}\mathscr{F}_{\varepsilon}(E_{\varepsilon},\Omega)<+\infty,$
Now, following the ideas of [MR1286918] (see also [MR1656477, MR1634336, chambolle_thouroude]), we consider the Radon measure
(9.1) $\lambda_{\varepsilon}(\cdot):=\iiint_{(\cdot)\cap\Omega\times\mathbb{R}^{n}}\chi_{E_{\varepsilon}}(x)\chi_{E_{\varepsilon}^{c}}(y)K_{\varepsilon}(x,y)\,dx\,dy+\sum_{\begin{subarray}{c}\ell\in\mathbb{Z}^{n}\\ \varepsilon(\ell+Q)\subseteq\Omega\end{subarray}}\delta_{\varepsilon\ell}(\cdot)\iint_{E_{\varepsilon}\cap\varepsilon(\ell+Q)}g_{\varepsilon}(x)dx,$
where, for any set $A\subseteq\mathbb{R}^{n}$,
$\delta_{\varepsilon\ell}(A):=\begin{cases}0,\quad&\mbox{if }\varepsilon\ell\not\in A,\\ 1,\quad&\mbox{if }\varepsilon\ell\in A.\end{cases}$
By Proposition 7.3, we have that $\lambda_{\varepsilon}\geqslant 0$. Moreover, the total mass of $\lambda_{\varepsilon}$ is bounded uniformly in $\varepsilon$. Indeed, by construction, we have that
$\lambda_{\varepsilon}(\mathbb{R}^{n})=\lambda_{\varepsilon}(\Omega)=\iint_{\Omega\times\mathbb{R}^{n}}\chi_{E_{\varepsilon}}(x)\chi_{E^{c}_{\varepsilon}}(y)K_{\varepsilon}(x,y)\,dx\,dy+\int_{E_{\varepsilon}\cap\mathcal{Q}(\Omega)_{\varepsilon}}g_{\varepsilon}(x)dx$
$\qquad\qquad=\mathscr{E}_{\varepsilon}(E_{\varepsilon},\Omega)\leqslant\mathscr{F}_{\varepsilon}(E_{\varepsilon},\Omega)\leqslant\sup_{\varepsilon}\mathscr{F}_{\varepsilon}(E_{\varepsilon},\Omega)<+\infty,$
where $\mathscr{E}_{\varepsilon}$ si defined as in (1.20).
Therefore, there exists a positive measure $\lambda$ such that
[…]
and $\lambda(A)\leqslant\liminf_{\varepsilon\to 0}\lambda_{\varepsilon}(A),\quad\mbox{for every set\penalty 10000\ $A$ such that\penalty 10000\ $\lambda(\partial A)=0$.}$
In particular, we have that
(9.2) $\lambda(\Omega)\leqslant\liminf_{\varepsilon\to 0}\lambda_{\varepsilon}(\Omega)\leqslant\liminf_{\varepsilon\to 0}\mathscr{F}_{\varepsilon}(E_{\varepsilon},\Omega).$
Then, let $\overline{x}$ be a regular point of $E$, namely a point $\overline{x}\in\partial E$ with (n1)(n-1)-density 11 and such that the blow-up of $E$ around $\overline{x}$ converges to $\{(x-\overline{x})\cdot\nu_{E}(\overline{x})>0\}$, e.g.
$\lim_{r\to 0}\frac{\mathcal{H}^{n-1}(\partial E\cap B_{r}(\overline{x}))}{\omega_{n-1}r^{n-1}}=1,$
and $\lim_{r\to 0^{+}}r^{-n}\int_{B_{2r}(\overline{x})}|\chi_{E}-\chi_{\{(x-\overline{x})\cdot\nu_{E}(\overline{x})>0\}}|dx=0,$
where $\omega_{n-1}:=\mathcal{H}^{n-1}(B^{n-1}_{1})$. We consider the Radon-Nikodyn derivative of $\lambda$ with respect to the (n1)(n-1)-dimensional Hausdorff measure at $\overline{x}$
$\frac{d\lambda}{d\mathcal{H}^{n-1}}(\overline{x}):=\lim_{r\to 0}\frac{\lambda(B_{r}(\overline{x}))}{\mathcal{H}^{n-1}(\partial E\cap B_{r}(\overline{x}))}=\lim_{r\to 0}\frac{\lambda(B_{r}(\overline{x}))}{\omega_{n-1}r^{n-1}},$
and observe that, in light of (9.2) and the Besicovitch derivation theorem (see [MR1857292, Theorem 5.52]), Theorem 1.13-(i) will follow if we show that
(9.3) $\frac{d\lambda}{d\mathcal{H}^{n-1}}(\overline{x})\geqslant\phi(\nu_{E}(\overline{x})).$
At this stage, let us consider subsequences $\{\varepsilon_{j}\}_{j}$, $\{r_{j}\}_{j}$, and $\{\eta_{j}:=\varepsilon_{j}/r_{j}\}_{j}$ such that
(9.4) $\begin{split}&\lim_{j\to+\infty}\frac{\lambda_{\varepsilon_{j}}(B_{r_{j}}(\overline{x}))}{\omega_{n-1}r_{j}^{n-1}}=\frac{d\lambda}{d\mathcal{H}^{n-1}}(\overline{x})\\ &\lim_{j\to+\infty}r_{j}^{-n}\int_{B_{2r_{j}}(\overline{x})}|\chi_{E}-\chi_{\{(x-\overline{x})\cdot\nu_{E}(\overline{x})>0\}}|dx=0,\\ \mbox{and}\quad&\lim_{j\to+\infty}\eta_{j}=0.\end{split}$
In what follows, in order to lighten the notation, we will write $E_{j}:=E_{\varepsilon_{j}}$, $\lambda_{\varepsilon_{j}}:=\lambda_{j}$, and $\nu:=\nu_{E}(\overline{x})$. We now focus on proving (9.3), that now is
$\lim_{j\to+\infty}\frac{\lambda_{j}(B_{r_{j}}(\overline{x}))}{\omega_{n-1}r_{j}^{n-1}}\geqslant\phi(\nu).$
To achieve this, let us introduce, for every set $A$, the notation $A^{\prime}:=(A-\overline{x})/r_{j}$. Thus, taking advantage of the scaling properties of $\lambda_{j}$, we obtain
(9.5) $\begin{split}&\frac{\lambda_{j}(B_{r_{j}}(\overline{x}))}{r_{j}^{n-1}}\\ =&r_{j}^{1-n}\left(\iint_{B_{r_{j}}(\overline{x})\times\mathbb{R}^{n}}\chi_{E_{j}}(x)\chi_{E^{c}_{j}}(y)K_{\varepsilon_{j}}(x,y)\,dx\,dy+\int_{E_{j}\cap\mathcal{Q}(B_{r_{j}}(\overline{x}))_{\varepsilon_{j}}}g_{\varepsilon_{j}}(x)dx\right)\\ =&\iint_{B_{1}\times\mathbb{R}^{n}}\chi_{E^{\prime}_{j}}(x)\chi_{E^{\prime c}_{j}}(y)K_{\eta_{j}}(x+\overline{x}/r_{j},y+\overline{x}/r_{j})\,dx\,dy\\ &+\int_{E^{\prime}_{j}\cap\mathcal{Q}^{\prime}(B_{r_{j}}(\overline{x}))_{\varepsilon_{j}}}g_{\eta_{j}}(x+\overline{x}/r_{j})dx,\end{split}$
where
$\mathcal{Q}^{\prime}(B_{r_{j}}(\overline{x}))_{\varepsilon_{j}}:=\left(\mathcal{Q}(B_{r_{j}}(\overline{x}))_{\varepsilon_{j}}-\overline{x}\right)/r_{j}=\bigcup\left\{\frac{\varepsilon_{j}(k+Q)-\overline{x}}{r_{j}}\mbox{s.t. }k\in\mathbb{Z}^{n},\ \varepsilon_{j}(k+Q)\subseteq B_{r_{j}}(\overline{x})\right\}.$
We also set $A^{\prime\prime}:=A^{\prime}+\eta_{j}\{\overline{x}/\varepsilon_{j}\}$, where $\{x\}_{j}:=\{x_{j}\}=x_{j}-\lfloor x_{j}\rfloor$ is the fractional part of the $j$-th component of a point $x$. Then, (9.5) boils down to
$\begin{split}&\frac{\lambda_{j}(B_{r_{j}}(\overline{x}))}{r_{j}^{n-1}}=\iint_{B_{1}(\eta_{j}\{\overline{x}/\varepsilon_{j}\})\times\mathbb{R}^{n}}\chi_{E^{\prime\prime}_{j}}(x)\chi_{E^{\prime\prime c}_{j}}(y)K_{\eta_{j}}(x,y)\,dx\,dy+\int_{E^{\prime\prime}_{j}\cap\mathcal{Q}^{\prime\prime}(B_{r_{j}}(\overline{x}))_{\varepsilon_{j}}}g_{\eta_{j}}(x)dx\\ &\qquad=\iint_{B^{\prime\prime}_{j}\times\mathbb{R}^{n}}\chi_{E^{\prime\prime}_{j}}(x)\chi_{E^{\prime\prime c}_{j}}(y)K_{\eta_{j}}(x,y)\,dx\,dy+\int_{E^{\prime\prime}_{j}\cap\mathcal{Q}^{\prime\prime}(B_{r_{j}}(\overline{x}))_{\varepsilon_{j}}}g_{\eta_{j}}(x)dx,\end{split}$
where, to be consistent with our notation, we wrote $B_{j}^{\prime\prime}:=B_{1}(\eta_{j}\{\overline{x}/\varepsilon_{j}\})$. Moreover, observe that
$\begin{split}&\mathcal{Q}^{\prime\prime}(B_{r_{j}}(\overline{x}))_{\varepsilon_{j}}=\bigcup\left\{\frac{\varepsilon_{j}(k+Q)-\overline{x}}{r_{j}}+\eta_{j}\left\{\frac{\overline{x}}{\varepsilon_{j}}\right\}\mbox{s.t. }k\in\mathbb{Z}^{n},\ \varepsilon_{j}(k+Q)\subseteq B_{r_{j}}(\overline{x})\right\}\\ &\qquad=\bigcup\left\{\eta_{j}\left(k-\left\lfloor\frac{\overline{x}}{\varepsilon_{j}}\right\rfloor+Q\right)\mbox{s.t. }k\in\mathbb{Z}^{n},\ \eta_{j}\left(k-\left\lfloor\frac{\overline{x}}{\varepsilon_{j}}\right\rfloor+Q\right)\subseteq B^{\prime\prime}_{j}(\overline{x})\right\}\\ &\qquad=\mathcal{Q}(B^{\prime\prime}_{j})_{\eta_{j}}.\end{split}$
Hence, it follows from the last two equations in display that
(9.6) $\frac{\lambda_{j}(B_{r_{j}}(\overline{x}))}{\omega_{n-1}r_{j}^{n-1}}=\frac{\mathscr{E}_{\eta_{j}}(E^{\prime\prime}_{j},B^{\prime\prime}_{j})}{\omega_{n-1}}.$
Furthermore, since $\eta_{j}\to 0$ and the fractional part is a bounded function, we deduce from (9.4) that
(9.7) $\begin{split}&\lim_{j\to+\infty}\int_{B_{3/2}}|\chi_{E^{\prime\prime}_{j}}-\chi_{\mathcal{I}_{\nu}}|dx\\ \leqslant\;&\lim_{j\to+\infty}\int_{B_{3/2}}|\chi_{E^{\prime\prime}_{j}}-\chi_{\{(x-\eta_{j}\{\overline{x}/\varepsilon_{j}\})\cdot\nu>0\}}|dx+\lim_{j\to+\infty}\int_{B_{3/2}}|\chi_{\{(x-\eta_{j}\{\overline{x}/\varepsilon_{j}\})\cdot\nu>0\}}-\chi_{\mathcal{I}_{\nu}}|dx\\ =\;&0.\end{split}$
In particular, $\chi_{E^{\prime\prime}_{j}}\to\chi_{\mathcal{I}_{\nu}}$ a.e. in $B_{3/2}$ (up to a subsequence).
Now, let $E_{\nu}$ be a planelike minimizer for $\mathcal{J}$ in direction $\nu$ constructed combining Theorem 1.4, Theorem 1.9, and Corollary 1.12. Then, let us define sets
(9.8) $F_{j}:=(\eta_{j}E_{\nu}\setminus\mathcal{Q}(B^{\prime\prime}_{j})_{1})\cup(E^{\prime\prime}_{j}\cap\mathcal{Q}(B^{\prime\prime}_{j})_{1}).$
Figure 2. The competitor $F_{j}$ as in (9.8) that produces the correct $\Gamma-\liminf$ inequality.
Since $F_{j}\setminus\mathcal{Q}(B^{\prime\prime}_{j})_{1}=\eta_{j}E_{\nu}\setminus\mathcal{Q}(B^{\prime\prime}_{j})_{1}$, by Theorem 1.9, Remark 1.8 and Corollary 8.5, we infer that
(9.9) $\mathscr{F}_{\eta_{j}}(F_{j},B^{\prime\prime}_{j})\geqslant\mathscr{F}_{\eta_{j}}(\eta_{j}E_{\nu},B^{\prime\prime}_{j})\to\mathcal{H}^{n-1}(\partial\mathcal{I}_{\nu}\cap B_{1})\phi(\nu)=\omega_{n-1}\phi(\nu),$
as $j\to+\infty$.
Also, observe that
(9.10) $\mathscr{E}_{\eta_{j}}(E^{\prime\prime}_{j},B^{\prime\prime}_{j})-\mathscr{F}_{\eta_{j}}(F_{j},B^{\prime\prime}_{j})\geqslant-\mathcal{L}_{K_{\eta_{j}}}(F_{j}\cap B^{\prime\prime}_{j},F_{j}^{c}\cap(B^{\prime\prime}_{j})^{c}).$
Since both $E_{j}^{\prime\prime}\to\mathcal{I}_{\nu}$ and $\eta_{j}E_{\nu}\to\mathcal{I}_{\nu}$ a.e. in $\mathbb{R}^{n}$, thanks to Lemma 7.1, we have that
(9.11) $\lim_{j\to\infty}\mathcal{L}_{K_{\eta_{j}}}(F_{j}\cap B^{\prime\prime}_{j},F_{j}^{c}\cap(B^{\prime\prime}_{j})^{c})=0.$
Therefore, putting together (9.9), (9.10), and (9.11), we find that
$\liminf_{j\to+\infty}\mathscr{E}_{\eta_{j}}(E^{\prime\prime}_{j},B^{\prime\prime}_{j})\geqslant\lim_{j\to+\infty}\mathscr{F}_{\eta_{j}}(\eta_{j}E_{\nu},B^{\prime\prime}_{j})-\mathcal{L}_{K_{\eta_{j}}}(F_{j}\cap B^{\prime\prime}_{j},F_{j}^{c}\cap(B^{\prime\prime}_{j})^{c})=\omega_{n-1}\phi(\nu).$
This, together with (9.6), entails the desired result.
10. Continuity of the stable norm $\phi$
In this section, we discuss the continuity of the stable norm $\phi$. This result is a version of [chambolle_thouroude, Corollary A.4] covering our setting, which will play a crucial role in proving the $\Gamma-\limsup$ inequality (see Theorem 1.13-(ii)) in Section 11.
Remark 10.1.
Before presenting our result, we point out that, in [chambolle_thouroude, Corollary A.4], the authors claim that the same argument of [MR1852978, Lemma 10.2] suffices to prove the convexity of the stable norm $\phi$. However, because of—plausibly—a typo in the proof of the latter, such argument cannot be directly applied in our context.
Proposition 10.2.
$\phi:\mathbb{S}^{n-1}\to\mathbb{R}$ is a continuous function.
Proof.
Let $p$, $q\in\mathbb{S}^{n-1}$ and define $\theta$ as the angle between $p$ and $q$ (i.e. $\theta$ is defined through $p\cdot q=\cos\theta$), so that if $\mathcal{R}_{\theta}\in SO(n)$ denotes a rotation of angle $\theta$, then $p=\mathcal{R}_{\theta}q$.
The continuity of $\phi$ will follow if we show the existence of a function $\psi$ such that
$|\phi(q)-\phi(p)|\leqslant\psi(\theta)\qquad\mbox{and}\qquad\lim_{\theta\to 0}\psi(\theta)=0.$
To this purpose, we use Theorem 1.13-(i) (notice indeed that $\varepsilon\mathcal{R}_{\theta}E_{q}$ converges to $\mathcal{I}_{p}$ as $\varepsilon\to 0$) and Proposition 8.1 to obtain
$\liminf_{\varepsilon\to 0}\mathscr{F}_{\varepsilon}(\varepsilon\mathcal{R}_{\theta}E_{q},Q^{p})\geqslant\mathscr{F}_{\phi}(\mathcal{I}_{p},Q^{p})=\phi(p)=\lim_{\varepsilon\to 0}\mathscr{F}_{\varepsilon}(\varepsilon E_{p},Q^{p}).$
From this and the fact that $K$ is rotation invariant (see (1.2)), it follows that
(10.1) $\begin{split}\phi(q)-\phi(p)&=\lim_{\varepsilon\to 0}\Big(\mathscr{F}_{\varepsilon}(\varepsilon E_{q},Q^{q})-\mathscr{F}_{\varepsilon}(\varepsilon E_{p},Q^{p})\Big)\\ &\geqslant\liminf_{\varepsilon\to 0}\Big(\mathscr{F}_{\varepsilon}(\varepsilon E_{q},Q^{q})-\mathscr{F}_{\varepsilon}(\varepsilon\mathcal{R}_{\theta}E_{q},Q^{p})\Big)\\ &=\liminf_{\varepsilon\to 0}\varepsilon^{n-1}\left(\int_{E_{q}\cap\mathcal{Q}(Q^{q}/\varepsilon)_{1}}g-\int_{\mathcal{R}_{\theta}E_{q}\cap\mathcal{Q}(Q^{p}/\varepsilon)_{1}}g\right).\end{split}$
We now recall that
$\partial E_{q}\subseteq\{|x\cdot q|\leqslant M\}=:S^{q}_{M}$
and hence
$\bigcup\big\{k+Q\mbox{ s.t. }k\in\mathbb{Z}^{n}\mbox{ and }(k+Q)\cap\partial E_{q}\neq\varnothing\big\}\subseteq S_{M+\sqrt{n}}^{q}.$
Thus, in light of (1.1), we find that
$\int_{E_{q}\cap\mathcal{Q}(Q^{q}/\varepsilon)_{1}}g-\int_{\mathcal{R}_{\theta}E_{q}\cap\mathcal{Q}(Q^{p}/\varepsilon)_{1}}g=\int_{E_{q}\cap\mathcal{Q}(Q^{q}/\varepsilon)_{1}\cap S^{q}_{M+\sqrt{n}}}g-\int_{\mathcal{R}_{\theta}E_{q}\cap\mathcal{Q}(Q^{p}/\varepsilon)_{1}\cap\mathcal{R}_{\theta}S^{q}_{M+\sqrt{n}}}g.$
Therefore, using also a Hölder inequality in (10.1) and observing that $\mathcal{R}_{\theta}S^{q}_{M+\sqrt{n}}=S^{p}_{M+\sqrt{n}}$, we obtain
$\begin{split}&\phi(q)-\phi(p)\\ &\quad\geqslant\liminf_{\varepsilon\to 0}\left(-\varepsilon^{n-1}\|g\|_{L^{\infty}(\mathbb{R}^{n})}\left|\Big(E_{q}\cap\mathcal{Q}(Q^{q}/\varepsilon)_{1}\cap S^{q}_{M+\sqrt{n}}\Big)\Delta\Big(\mathcal{R}_{\theta}E_{q}\cap\mathcal{Q}(Q^{p}/\varepsilon)_{1}\cap\mathcal{R}_{\theta}S^{q}_{M+\sqrt{n}}\Big)\right|\right)\\ &\quad\geqslant\liminf_{\varepsilon\to 0}\left(-\varepsilon^{n-1}\|g\|_{L^{\infty}(\mathbb{R}^{n})}\left|\Big(E_{q}\cap\mathcal{Q}(Q^{q}/\varepsilon)_{1}\cap S^{q}_{M+\sqrt{n}}\Big)\Delta\Big(\mathcal{R}_{\theta}E_{q}\cap\mathcal{Q}(Q^{p}/\varepsilon)_{1}\cap S^{p}_{M+\sqrt{n}}\Big)\right|\right).\end{split}$
Now, observe that, arguing as in (8.25), we have that
$\begin{split}&\left|(E_{q}\cap\mathcal{Q}(Q^{q}/\varepsilon)_{1}\cap S^{q}_{M+\sqrt{n}})\Delta(\mathcal{R}_{\theta}E_{q}\cap\mathcal{Q}(Q^{p}/\varepsilon)_{1}\cap S^{p}_{M+\sqrt{n}})\right|\\ &\qquad\leqslant\left|[E_{q}\cap S^{q}_{M+\sqrt{n}}\cap\frac{1}{\varepsilon}C^{q}]\Delta[\mathcal{R}_{\theta}E_{q}\cap S^{p}_{M+\sqrt{n}}\cap\frac{1}{\varepsilon}C^{p}]\right|+\varepsilon^{1-n}\zeta(\varepsilon),\end{split}$
where $C^{q}$ is the (unitary) square-based cylinder with axis $q$ and
$\begin{split}\varepsilon^{1-n}\zeta(\varepsilon):=&\left|S^{q}_{M+\sqrt{n}}\cap S^{p}_{M+\sqrt{n}}\cap\mathcal{R}_{\theta}E^{q}\cap\left[C^{p}/\varepsilon\setminus\mathcal{Q}(Q^{p}/\varepsilon)_{1}\right]\right|\\ &+\left|S^{q}_{M+\sqrt{n}}\cap S^{p}_{M+\sqrt{n}}\cap E^{q}\cap\left[C^{q}/\varepsilon\setminus\mathcal{Q}(Q^{q}/\varepsilon)_{1}\right]\right|,\end{split}$
with $\zeta(\varepsilon)\to 0$, as $\varepsilon\to 0$.
Thus,
$\begin{split}&\phi(q)-\phi(p)\\ \geqslant&\liminf_{\varepsilon\to 0}-\varepsilon^{n-1}\|g\|_{\infty}\left|[E_{q}\cap S^{q}_{M+\sqrt{n}}\cap\frac{1}{\varepsilon}C^{q}]\Delta[\mathcal{R}_{\theta}E_{q}\cap S^{p}_{M+\sqrt{n}}\cap\frac{1}{\varepsilon}C^{p}]\right|.\end{split}$
Let us define
$\psi(\theta):=\limsup_{\varepsilon\to 0}\varepsilon^{n-1}\left|[E_{q}\cap S^{q}_{M+\sqrt{n}}\cap\frac{1}{\varepsilon}C^{q}]\Delta[\mathcal{R}_{\theta}E_{q}\cap S^{p}_{M+\sqrt{n}}\cap\frac{1}{\varepsilon}C^{p}]\right|,$
so that
$\phi(q)-\phi(p)\geqslant-\|g\|_{\infty}\psi(\theta).$
Moreover, notice that $\psi(\theta)\to 0$, as $\theta\to 0$ (namely as $p\to q$).
Interchanging the roles of $p$ and $q$, we obtain again that
$\begin{split}&\phi(p)-\phi(q)\geqslant\liminf_{\varepsilon\to 0}-\|g\|_{\infty}\left|[E_{q}\cap S^{p}_{M+\sqrt{n}}\cap C^{p}]\Delta[\mathcal{R}_{-\theta}\cap S^{q}_{M+\sqrt{n}}\cap C^{q}]\right|\\ &\qquad\qquad=-\|g\|_{\infty}|J\mathcal{R}_{\theta}|\psi(\theta)=-\|g\|_{\infty}\psi(\theta),\end{split}$
and hence
$|\phi(q)-\phi(p)|\leqslant\|g\|_{\infty}\psi(\theta),$
as desired. ∎
Now, we define the $1$-homogeneous extension of $\phi$ as
(10.2) $\widetilde{\phi}(p):=\begin{cases}|p|\phi\left(\frac{p}{|p|}\right),\quad&{\mbox{ if }}p\neq 0,\\ 0,\quad&{\mbox{ if }}p=0.\end{cases}$
As a byproduct of Proposition 10.2, we obtain that also $\widetilde{\phi}$ is continuous (in the whole $\mathbb{R}^{n}$).
Corollary 10.3.
$\widetilde{\phi}:\mathbb{R}^{n}\to\mathbb{R}$ is a continuous function.
Proof.
Let $q\in\mathbb{R}^{n}$. Then, it follows from the homogeneity of $\widetilde{\phi}$ and Proposition 10.2 that
$\begin{split}&\lim_{p\to q}\widetilde{\phi}(p)=\lim_{p\to q}|p|{\phi}\left(\frac{p}{|p|}\right)=\lim_{p\to q}\left((|p|-|q|){\phi}\left(\frac{p}{|p|}\right)+|q|{\phi}\left(\frac{p}{|p|}\right)\right)\\ &\qquad=\lim_{p\to q}\left((|p|-|q|){\phi}\left(\frac{p}{|p|}\right)+|q|{\phi}\left(\frac{q}{|q|}\right)\right)=\lim_{p\to q}\left((|p|-|q|){\phi}\left(\frac{p}{|p|}\right)\right)+\widetilde{\phi}(q).\end{split}$
Since ${\phi}(p/|p|)$ is bounded, we conclude that
$\lim_{p\to q}\widetilde{\phi}(p)=\widetilde{\phi}(q),$
showing the continuity of $\widetilde{\phi}$ at $q$. ∎
11. Proof of the $\Gamma$-$\limsup$ inequality
This section is devoted to proving $\Gamma$-$\limsup$ inequality in Theorem 1.13-(ii). Our argument adapts to the setting under consideration here the ideas of [MR1852978, chambolle_thouroude], in which the result follows from a standard polyhedral approximation taking advantage of the existence of planelike minimizers. For this, we need the following technical result.
Lemma 11.1.
Let $E\subseteq\mathbb{R}^{n}$ be such that $E\cap\Omega$ is a polyhedron. Then, there exists a sequence of sets $\{E_{\varepsilon}\}_{\varepsilon}$ such that
$\limsup_{\varepsilon\to 0}\mathscr{F}_{\varepsilon}(E_{\varepsilon},\Omega)\leqslant\mathscr{F}_{\phi}(E,\Omega).$
Proof.
Since $E\cap\Omega$ is a polyhedron, there exist points $x_{j}\in\partial E\cap\Omega$ and directions $p_{j}\in\mathbb{S}^{n-1}$, with $j=1,\dots,N$, for some $N\in\mathbb{N}$, such that
$\partial E\cap\Omega=\bigcup_{j=1}^{N}(x_{j}+\mathcal{I}_{p_{j}})\cap\Omega.$
Here above and in the rest of the proof, the notation in (7.1) for $\mathcal{I}_{p_{j}}$ is used.
Also, there are $N$ disjoint Lipschitz domains $\Omega_{j}$ such that
$\Omega=\bigcup_{j=1}^{N}\Omega_{j}\qquad{\mbox{and}}\qquad(x_{j}+\mathcal{I}_{p_{j}})\cap\Omega\subseteq\Omega_{j}.$
Now, for any $\varepsilon>0$, we define
(11.1) $E_{\varepsilon}:=(E\setminus\Omega)\cap\bigcup_{j=1}^{N}\Big(\Omega_{j}\cap\big(x_{j}+\varepsilon E_{p_{j}}\big)\Big),$
where, for every $j=1,\dots,N$, $E_{p_{j}}$ is a planelike minimizer constructed as in Theorem 1.4 (see Figure 3).
Figure 3. The planelike approximation of the polygonal set $E$ as in (11.1).
Therefore, we obtain
(11.2) $\begin{split}\mathscr{F}_{\varepsilon}(E_{\varepsilon},\Omega)&\leqslant\sum_{j=1}^{N}\mathscr{F}_{\varepsilon}(x_{j}+\varepsilon E_{p_{j}},\Omega_{j})+\int_{E_{\varepsilon}\cap(\mathcal{Q}(\Omega)_{\varepsilon}\setminus\bigcup_{j=1}^{N}\mathcal{Q}(\Omega_{j})_{\varepsilon})}g_{\varepsilon}(x)\,dx\\ &\quad+\sum_{j=1}^{N}\mathcal{L}_{K_{\varepsilon}}(\varepsilon E_{p_{j}}\cap\Omega_{j},E_{\varepsilon}^{c}\cap\Omega_{j}^{c})+\sum_{j=1}^{N}\mathcal{L}_{K_{\varepsilon}}(\varepsilon E_{p_{j}}^{c}\cap\Omega_{j},E_{\varepsilon}\cap\Omega_{j}^{c})\\ &\quad-\sum_{j=1}^{N}\mathcal{L}_{K_{\varepsilon}}(\varepsilon E_{p_{j}}\cap\Omega_{j},\varepsilon E_{p_{j}}^{c}\cap\Omega_{j}^{c})-\sum_{j=1}^{N}\mathcal{L}_{K_{\varepsilon}}(\varepsilon E_{p_{j}}^{c}\cap\Omega_{j},\varepsilon E_{p_{j}}\cap\Omega_{j}^{c}).\end{split}$
By Corollary 8.5, for every $j\in\{1,\dots,N\}$, we have that
(11.3) $\lim_{\varepsilon\to 0}\mathscr{F}_{\varepsilon}(x_{j}+\varepsilon E_{p_{j}},\Omega_{j})=\phi(p_{j})\mathcal{H}^{n-1}\big(\Omega_{j}\cap(x_{j}+\partial\mathcal{I}_{p_{j}})\big).$
Furthermore,
$\begin{split}&\int_{E_{\varepsilon}\cap(\mathcal{Q}(\Omega)_{\varepsilon}\setminus\bigcup_{j=1}^{N}\mathcal{Q}(\Omega_{j})_{\varepsilon})}g_{\varepsilon}(x)\,dx\leqslant\varepsilon^{n-1}\int_{(E_{\varepsilon}/\varepsilon)\cap(\mathcal{Q}(\Omega/\varepsilon)_{1}\setminus\bigcup_{j=1}^{N}\mathcal{Q}(\Omega_{j}/\varepsilon)_{1})}|g(x)|\,dx\\ &\qquad=\varepsilon^{n-1}\sum_{j=1}^{N}\int_{(x_{j}/\varepsilon+E_{p_{j}})\cap(\Omega_{j}/\varepsilon)\cap(\mathcal{Q}(\Omega/\varepsilon)_{1}\setminus\mathcal{Q}(\Omega_{j}/\varepsilon)_{1}}|g(x)|\,dx\\ &\qquad\leqslant 2(M+\sqrt{n})\|g\|_{L^{1}(Q)}\varepsilon^{n-1}\sum_{j=1}^{N}\Big(\varepsilon^{1-n}\mathcal{H}^{n-1}(\Omega_{j}\cap\partial\mathcal{I}_{p_{j}})-\lfloor\varepsilon^{1-n}\mathcal{H}^{n-1}(\Omega_{j}\cap\partial\mathcal{I}_{p_{j}})\rfloor\Big).\end{split}$
Thus,
(11.4) $\limsup_{\varepsilon\to 0}\int_{E_{\varepsilon}\cap(\mathcal{Q}(\Omega)_{\varepsilon}\setminus\bigcup_{j=1}^{N}\mathcal{Q}(\Omega_{j})_{\varepsilon})}g_{\varepsilon}(x)\,dx\leqslant 0.$
Moreover, thanks to Lemma 7.1, we infer that
(11.5) $\begin{split}\lim_{\varepsilon\to 0}&\sum_{j=1}^{N}\mathcal{L}_{K_{\varepsilon}}(\varepsilon E_{p_{j}}\cap\Omega_{j},E_{\varepsilon}^{c}\cap\Omega_{j}^{c})+\sum_{j=1}^{N}\mathcal{L}_{K_{\varepsilon}}(\varepsilon E_{p_{j}}^{c}\cap\Omega_{j},E_{\varepsilon}\cap\Omega_{j}^{c})\\ &-\sum_{j=1}^{N}\mathcal{L}_{K_{\varepsilon}}(\varepsilon E_{p_{j}}\cap\Omega_{j},\varepsilon E_{p_{j}}^{c}\cap\Omega_{j}^{c})-\sum_{j=1}^{N}\mathcal{L}_{K_{\varepsilon}}(\varepsilon E_{p_{j}}^{c}\cap\Omega_{j},\varepsilon E_{p_{j}}\cap\Omega_{j}^{c})=0,\end{split}$
for all $j\in\{1,\dots,N\}$.
Hence, gathering (11.2), (11.3), (11.4), and (11.5), we conclude that
$\limsup_{\varepsilon\to 0}\mathscr{F}_{\varepsilon}(E_{\varepsilon},\Omega)\leqslant\sum_{j=1}^{N}\phi(p_{j})\mathcal{H}^{n-1}\big(\Omega_{j}\cap(x_{j}+\partial\mathcal{I}_{p_{j}})\big)=\int_{\partial E\cap\Omega}\phi(\nu_{E})\,d\mathcal{H}^{n-1}.$
as desired. ∎
Proof of Theorem 1.13-(ii).
By a standard approximation argument, there exists a sequence of sets $\{E_{j}\}_{j}$ such that $E_{j}\cap\Omega$ is a polyhedron, for every $j$, and
$\lim_{j\to+\infty}\mathcal{H}^{n-1}(E_{j}\cap\Omega)=\mbox{Per}(E,\Omega).$
Therefore, recalling that (the $1$-homogeneous extension of) $\phi$ is continuous by Corollary 10.3, we deduce from [MR1857292, Theorem 2.39] that
$\lim_{j\to+\infty}\mathscr{F}_{\phi}(E_{j},\Omega)=\mathscr{F}_{\phi}(E,\Omega).$
Using Lemma 11.1 and taking a diagonal sequence, we construct a sequence of sets $\{E_{\varepsilon_{j}}\}_{j}$ such that $E_{\varepsilon_{j}}\to E$ in $L^{1}_{\rm{loc}}(\mathbb{R}^{n})$, and
$\limsup_{j\to+\infty}\mathscr{F}_{\varepsilon_{j}}(E_{\varepsilon_{j}},\Omega)\leqslant\mathscr{F}_{\phi}(E,\Omega),$
showing Theorem 1.13-(ii). ∎
Remark 11.2.
We point out that, in light of Lemma 7.1, one can equivalently consider energies $\mathscr{F}_{\varepsilon}((\cdot),\Omega)$, $\mathscr{E}_{\varepsilon}((\cdot),\Omega)$, or even $\mathscr{L}_{\varepsilon}((\cdot),\Omega):=\mathcal{L}_{K_{\varepsilon}}((\cdot)\cap\Omega,(\cdot)^{c}\cap\Omega)$ in Theorem 1.13. Indeed, since the rescaled planelike $\varepsilon E_{p}$ converges to the straight line $\mathcal{I}_{p}$, as $\varepsilon\to 0$, such limit crosses transversely any Lipschitz domain. Therefore, the energy contribution due to the interactions with the exterior of $\Omega$ in the limit.
As a corollary of the $\Gamma$-convergence, we improve Proposition 10.2. Indeed, since a $\Gamma$-limit is lower semi-continuous by construction, we infer that the stable norm must be convex.
Recall the definitions of the stable norm $\phi:\mathbb{S}^{n-1}\to\mathbb{R}$ in (1.21), and of its homogenization $\widetilde{\phi}$ in (10.2). Then, the following holds true:
Corollary 11.3.
The function $\widetilde{\phi}:\mathbb{R}^{n}\to\mathbb{R}$ is convex.
Appendix A Norms subject to short-range interactions
We point out a useful equivalence of norms when a kernel is restricted to close-by interactions (see e.g. [MR5007887, Lemma 2.1] for related results). For the rest of this section, let $s\in(0,1/2)$.
Lemma A.1.
There exists $C>0$, depending only on $\delta$, $n$, and $s$, such that, for all $u:\mathbb{R}^{n}\to\mathbb{R}$ such that
$\int_{Q}u(x)\,dx=0,$
we have that
$\iint_{Q\times Q}\frac{|u(x)-u(y)|}{|x-y|^{n+2s}}\,dx\,dy\leqslant C\iint_{(Q\times Q)\cap\{|x-y|<\delta\}}\frac{|u(x)-u(y)|}{|x-y|^{n+2s}}\,dx\,dy.$
Proof.
Up to renaming constants, the desired result follows if we prove that
$\iint_{(Q\times Q)\cap\{|x-y|\geqslant\delta\}}\frac{|u(x)-u(y)|}{|x-y|^{n+2s}}\,dx\,dy\leqslant C\iint_{(Q\times Q)\cap\{|x-y|<\delta\}}\frac{|u(x)-u(y)|}{|x-y|^{n+2s}}\,dx\,dy.$
Moreover, since
$\begin{split}\iint_{(Q\times Q)\cap\{|x-y|\geqslant\delta\}}\frac{|u(x)-u(y)|}{|x-y|^{n+2s}}\,dx\,dy&\leqslant\frac{1}{\delta^{n+2s}}\iint_{Q\times Q}\big(|u(x)|+|u(y)|\big)\,dx\,dy\\ &=\frac{2}{\delta^{n+2s}}\int_{Q}|u(x)|\,dx,\end{split}$
it suffices to show that
$\int_{Q}|u(x)|\,dx\leqslant C\iint_{(Q\times Q)\cap\{|x-y|<\delta\}}\frac{|u(x)-u(y)|}{|x-y|^{n+2s}}\,dx\,dy.$
We argue for the sake of contradiction and suppose that this is not true. Namely, we suppose that there exists a sequence of functions $u_{j}:\mathbb{R}^{n}\to\mathbb{R}$ such that
(A.1) $\int_{Q}u_{j}(x)\,dx=0$
and
(A.2) $1=\int_{Q}|u_{j}(x)|\,dx>j\iint_{(Q\times Q)\cap\{|x-y|<\delta\}}\frac{|u_{j}(x)-u_{j}(y)|}{|x-y|^{n+2s}}\,dx\,dy.$
We decompose $Q$ as the disjoint union of cubes $Q_{1},\dots,Q_{N}$ of side less than $\frac{\delta}{2}$ and we deduce from (A.2) that, for all $\ell\in\{1,\dots,N\}$,
(A.3) $\frac{1}{j}>\iint_{\begin{subarray}{c}Q_{\ell}\times(Q_{\ell})_{\delta}\\ (Q\times Q)\cap\{|x-y|<\delta\}\end{subarray}}\frac{|u_{j}(x)-u_{j}(y)|}{|x-y|^{n+2s}}\,dx\,dy\geqslant\iint_{Q_{\ell}\times Q_{\ell}}\frac{|u_{j}(x)-u_{j}(y)|}{|x-y|^{n+2s}}\,dx\,dy,$
where
$(Q_{\ell})_{\delta}:=\{x\in\mathbb{R}^{n}\mbox{ s.t. }\mbox{dist}(x,Q_{\ell})<\delta\}.$
Then (see e.g. [MR2944369, Theorem 7.1]), the sequence $u_{j}$ is precompact in $L^{1}(Q_{\ell})$ and consequently, up to a subsequence, $u_{j}$ converges to some function $u_{0}$ in $L^{1}(Q_{\ell})$ for all $\ell\in\{1,\dots,N\}$, and thus in $L^{1}(Q)$ and a.e. in $Q$.
Thus, from (A.3) and Fatou Lemma,
$\iint_{Q_{\ell}\times Q_{\ell}}\frac{|u_{0}(x)-u_{0}(y)|}{|x-y|^{n+2s}}\,dx\,dy=0,$
which implies that $u_{0}$ is constant. As a result, since we have from (A.1) that
$\int_{Q}u_{0}(x)\,dx=0,$
we conclude that $u_{0}$ is the null function.
However, according to (A.2),
$1=\lim_{j\to+\infty}\int_{Q}|u_{j}(x)|\,dx=\int_{Q}|u_{0}(x)|\,dx,$
providing the desired contradiction. ∎
Appendix B Uniform density estimates for $\left(\Lambda,\frac{2s_{1}}{n}\right)$-minimizers of $\operatorname{P}_{K}$
In this section, we revisit the proof of [sequoia, Theorem 2.2] adapting it to the setting under consideration here. The precise statement goes as follows:
Proposition B.1.
Let $K$ satisfy (1.2), (1.3), (1.4), and (1.5) and let $\Omega\subseteq\mathbb{R}^{n}$ be a Lipschitz domain.
Then, there exists $\Lambda_{0}>0$ such that for all $\Lambda\in(0,\Lambda_{0}]$ the following statement holds true.
Let $E$ be a $\left(\Lambda,\frac{2s_{1}}{n}\right)$-minimal set for $\operatorname{P}_{K}$ in $\Omega$ (in the sense of Definition 5.2).
Then, there exists a constant $c_{0}\in(0,1)$, depending only on $n$, $s_{1}$, and $\kappa_{2}$, such that, for any $x_{0}\in(\partial E)\cap\Omega$ and $r\in(0,\min\{\delta/4,\mbox{dist}(x_{0},\partial\Omega)\})$,
(B.1) $c_{0}r^{n}\leqslant|E\cap B_{r}(x_{0})|\leqslant(1-c_{0})r^{n}.$
For the proof of Proposition B.1, we employ the following auxiliary results in Lemmata B.2 and B.3.
Lemma B.2 (Lemma 7.1, [MR1707291]).
Let $\beta\in(0,1)$, $N>1$, and $M>0$. Let $\{x_{k}\}_{k}$ be a decreasing sequence in $\mathbb{R}$ such that
$x_{k+1}^{1-\beta}\leqslant N^{k}Mx_{k}.$
If $x_{0}\leqslant N^{\frac{1}{\beta}-\frac{1}{\beta^{2}}}M^{-\frac{1}{\beta}}$, then $x_{k}\to 0$ as $k\to+\infty$.
Lemma B.3 (Isoperimetric lower bound for the kernel $K$).
Suppose that $K$ satisfy (1.2), (1.3), (1.4), and (1.5).
Then, there exists a positive constant $C$, depending on $n$, $s_{1}$, $\kappa_{1}$, and $\delta$, such that
(B.2) $\operatorname{P}_{K}(B_{r})\geqslant C\begin{cases}r^{n-2s_{1}},\quad&{\mbox{if }}\;0<r\leqslant\delta/4,\\ r^{n-1},\quad&{\mbox{if }}\;r>\delta/4.\end{cases}\quad$
Proof.
To show (B.2), let us start by assuming $0<r\leqslant\delta/4$. In this case, by (1.4) and the scaling properties of the kernel $|x|^{-n-2s_{1}}$, we have that
(B.3) $\begin{split}&P_{K}(B_{r})=\iint_{B_{r}\times B_{r}^{c}}K(x,y)\,dx\,dy\geqslant\iint_{B_{r}\times(B_{2r}\setminus B_{r})}K(x,y)\,dx\,dy\\ &\qquad\geqslant\kappa_{1}\iint_{B_{r}\times(B_{2r}\setminus B_{r})}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}\\ &\qquad=\kappa_{1}r^{n-2s_{1}}\iint_{B_{1}\times(B_{2}\setminus B_{1})}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}=:C_{1}(n,s_{1},\kappa_{1})r^{n-2s_{1}}.\end{split}$
Also, if $r>\delta/4$, let us consider the covering $\{B_{\delta/8}(x_{0})\}_{x_{0}\in\partial B_{r}}$ of $\partial B_{r}$. By compactness, we extract a finite sub-covering $\mathcal{B}$. In particular,
(B.4) $\sharp(\mathcal{B})=c\left(\frac{r}{\delta}\right)^{n-1},$
for some constant $c>0$ depending only on $n$. Thus, we have that
(B.5) $\begin{split}&P_{K}(B_{r})=\iint_{B_{r}\times B_{r}^{c}}K(x,y)\,dx\,dy\geqslant\sum_{B_{\delta/8}(x_{0})\in\mathcal{B}}\iint_{(B_{r}\cap B_{\delta/8}(x_{0}))\times(B_{r}^{c}\cap B_{\delta/8}(x_{0}))}K(x,y)\,dx\,dy\\ &\qquad\geqslant\kappa_{1}\sum_{B_{\delta/8}(x_{0})\in\mathcal{B}}\iint_{(B_{r}\cap B_{\delta/8}(x_{0}))\times(B_{r}^{c}\cap B_{\delta/8}(x_{0}))}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}\\ &\qquad\geqslant\kappa_{1}\delta^{n-2s_{1}}\sum_{B_{\delta/8}(x_{0})\in\mathcal{B}}\iint_{(B_{8r/\delta}\cap B_{1}(8x_{0}/\delta))\times(B_{8r/\delta}^{c}\cap B_{1}(8x_{0}/\delta))}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}.\end{split}$
We now use the notation
$x_{\delta}:=\frac{8x_{0}}{\delta}$
and we claim that
(B.6) $(\partial B_{8r/\delta})\cap B_{1}(x_{\delta})\subseteq\left\{-\frac{1}{4}<(x-x_{\delta})\cdot\frac{x_{0}}{|x_{0}|}\leqslant 0\right\}.$
To check this, we pick $x\in\partial B_{8r/\delta}\cap B_{1}(x_{\delta})$ and we observe that
$1>|x-x_{\delta}|^{2}=|x|^{2}+|x_{\delta}|^{2}-2x\cdot x_{\delta}=\left(\frac{8r}{\delta}\right)^{2}+\left(\frac{8}{\delta}\right)^{2}|x_{0}|^{2}-\frac{16x\cdot x_{0}}{\delta}=2\left(\frac{8r}{\delta}\right)^{2}-\frac{16x\cdot x_{0}}{\delta}$
and accordingly
(B.7) $x\cdot\frac{x_{0}}{|x_{0}|}=\frac{\delta}{16r}\;\frac{16x\cdot x_{0}}{\delta}>\frac{\delta}{16r}\left[2\left(\frac{8r}{\delta}\right)^{2}-1\right]=\frac{8r}{\delta}-\frac{\delta}{16r}.$
Moreover,
(B.8) $(x-x_{\delta})\cdot\frac{x_{0}}{|x_{0}|}=x\cdot\frac{x_{0}}{|x_{0}|}-|x_{\delta}|=x\cdot\frac{x_{0}}{|x_{0}|}-\frac{8|x_{0}|}{\delta}=x\cdot\frac{x_{0}}{|x_{0}|}-\frac{8r}{\delta},$
from which we arrive at
(B.9) $(x-x_{\delta})\cdot\frac{x_{0}}{|x_{0}|}\leqslant|x|-\frac{8r}{\delta}=0.$
Also, from (B.7) and (B.8),
$(x-x_{\delta})\cdot\frac{x_{0}}{|x_{0}|}>-\frac{\delta}{16r}>-\frac{1}{4}.$
Combining this and (B.9), we obtain (B.6), as desired.
It thereby follows from (B.6) that
$\iint_{(B_{8r/\delta}\cap B_{1}(x_{\delta}))\times(B_{8r/\delta}^{c}\cap B_{1}(x_{\delta}))}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}$
$\qquad\qquad\geqslant\iint_{\left(\left\{(x-x_{\delta})\cdot\frac{x_{0}}{|x_{0}|}>0\right\}\cap B_{1}(x_{\delta})\right)\times\left(\left\{(y-x_{\delta})\cdot\frac{x_{0}}{|x_{0}|}<-\frac{1}{4}\right\}\cap B_{1}(x_{\delta})\right)}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}.$
Plugging this information into (B.5), we obtain that
$\begin{split}&P_{K}(B_{r})\geqslant\kappa_{1}\delta^{n-2s_{1}}\sum_{B_{\delta/8}(x_{0})\in\mathcal{B}}\iint_{\left(\left\{(x-x_{\delta})\cdot\frac{x_{0}}{|x_{0}|}>0\right\}\cap B_{1}(x_{\delta})\right)\times\left(\left\{(y-x_{\delta})\cdot\frac{x_{0}}{|x_{0}|}<-\frac{1}{4}\right\}\cap B_{1}(x_{\delta})\right)}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}.\end{split}$
By the translation and rotation invariance of the fractional kernel $|x|^{-n-2s_{1}}$ and (B.4), we infer that
(B.10) $\begin{split}P_{K}(B_{r})&\geqslant\kappa_{1}\delta^{n-2s_{1}}\sum_{B_{\delta/8}(x_{0})\in\mathcal{B}}\iint_{\left(B_{1}\cap\{x_{n}>0\}\right)\times\left(B_{1}\cap\left\{y_{n}<-\frac{1}{4}\right\}\right)}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}\\ &\geqslant c\kappa_{1}\delta^{1-2s_{1}}r^{n-1}\iint_{(B_{1}\cap\{x_{n}>0\})\times(B_{1}\cap\{y_{n}<-1/4\})}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}\\ &=:C_{2}(n,\kappa_{1},s_{1},\delta)r^{n-1}.\end{split}$
We conclude the proof from (B.3) and (B.10) by setting $C:=\min\{C_{1},C_{2}\}$. ∎
Proof of Proposition B.1.
Let us recall that, by definition, if $E$ is a $\left(\Lambda,\frac{2s_{1}}{n}\right)$-minimal set, for some $\Lambda\geqslant 0$, then
(B.11) $\operatorname{P}_{K}(E,\Omega)\leqslant\operatorname{P}_{K}(F,\Omega)+\Lambda|E\Delta F|^{1-\frac{2s_{1}}{n}},$
for every $F$ such that $F\setminus\Omega=E\setminus\Omega$.
Let $x_{0}\in(\partial E)\cap\Omega$. Up to a translation, we suppose that $x_{0}$ coincides with the origin.
Define $A_{r}:=E\cap B_{r}$, with $r\in(0,\delta/4)$, where $\delta$ is as in (1.4), and observe that $A_{r}\subseteq\Omega$. Also, let $\mu(r):=|A_{r}|$ and notice that, by the co-area formula, $\mu^{\prime}(r)=\mathcal{H}^{n-1}(E\cap\partial B_{r})$.
Our strategy now is to provide an estimate for $\mu^{1-\frac{2s_{1}}{n}}(r)$ in terms of $\mu(r)$. For this, let us consider $\rho_{r}\in(0,r)$ such that $|A_{r}|=\omega_{n}\rho_{r}^{n}$. Then, observe that, thanks to Lemma B.3 and [MR3732175, Proposition 3.1],
(B.12) $\|\chi_{A_{r}}\|_{L^{\frac{n}{n-2s_{1}}}(\mathbb{R}^{n})}=|A_{r}|^{1-\frac{2s_{1}}{n}}\leqslant C\operatorname{P}_{K}(B_{\rho_{r}})\leqslant C\operatorname{P}_{K}(A_{r})=C\mathcal{L}_{K}(A_{r},A_{r}^{c}),$
up to renaming $C$, and we stress that $C$ depends only [^4] Notice indeed that, once we assume that $\rho_{r}<\delta/4$, we can select $C:=C_{1}$ given by (B.3) in Lemma B.3. Thus, the constant $C$ in (B.12) is independent of $\delta$. on $n$, $s_{1}$, and $\kappa_{1}$.
Now, since
$\mathcal{L}_{K}(A_{r},A_{r}^{c})=\mathcal{L}_{K}(A_{r},E\cap A_{r}^{c})+\mathcal{L}_{K}(A_{r},E^{c}),$
it follows from the subsolution property in [sequoia, Definition 1.2] that
(B.13) $\mathcal{L}_{K}(A_{r},A_{r}^{c})\leqslant 2\mathcal{L}_{K}(A_{r},E\cap A_{r}^{c})+\Lambda\mu^{1-\frac{2s_{1}}{n}}(r)\leqslant 2\mathcal{L}_{K}(A_{r},B_{r}^{c})+\Lambda\mu^{1-\frac{2s_{1}}{n}}(r).$
Moreover, by the co-area formula and (1.4), we have that
$\begin{split}&\mathcal{L}_{K}(A_{r},B_{r}^{c})=\iint_{A_{r}\times B_{r}^{c}}K(x,y)dx\,dy\leqslant\kappa_{2}\iint_{A_{r}\times B_{r}^{c}}\frac{dy\,dx}{|x-y|^{n+2s_{1}}}\leqslant C\int_{A_{r}}\left(\int_{r-|x|}^{+\infty}\frac{dz}{z^{2s_{1}+1}}\right)\,dx\\ &\qquad\quad\leqslant C\int_{A_{r}}\frac{dx}{(r-|x|)^{2s_{1}}}\leqslant C\int_{0}^{r}\frac{\mu^{\prime}(\rho)}{(r-\rho)^{2s_{1}}}\,d\rho,\end{split}$
for some $C>0$, depending on $n$, $s_{1}$, and $\kappa_{2}$ and possibly changing from line to line.
Plugging this into (B.13), we obtain that
$\mathcal{L}_{K}(A_{r},A_{r}^{c})\leqslant C\int_{0}^{r}\frac{\mu^{\prime}(\rho)}{(r-\rho)^{2s_{1}}}\,d\rho+\Lambda\mu^{1-\frac{2s_{1}}{n}}(r).$
From the last inequality and (B.12) we deduce that
(B.14) $\mu^{1-\frac{2s_{1}}{n}}(r)=\|\chi_{A_{r}}\|_{L^{\frac{n}{n-2s_{1}}}(\mathbb{R}^{n})}\leqslant C\left(\int_{0}^{r}\frac{\mu^{\prime}(\rho)}{(r-\rho)^{2s_{1}}}\,d\rho+\Lambda\mu^{1-\frac{2s_{1}}{n}}(r)\right),$
up to renaming $C$.
Furthermore, we assume that $\Lambda$ is so small that
$C\Lambda\leqslant\frac{1}{2}.$
Using this into (B.14), we thus obtain that
$\mu^{1-\frac{2s_{1}}{n}}(r)\leqslant C\int_{0}^{r}\frac{\mu^{\prime}(\rho)}{(r-\rho)^{2s_{1}}}\,d\rho.$
Hence, integrating the latter inequality in $r\in(0,t)$, we deduce that, for all $t\in(0,\delta/4]$,
(B.15) $\begin{split}\int_{0}^{t}\mu^{1-\frac{2s_{1}}{n}}(r)\,dr&\leqslant\int_{0}^{t}\left(C\int_{0}^{r}\frac{\mu^{\prime}(\rho)}{(r-\rho)^{2s_{1}}}\,d\rho\right)\,dr\\ &=C\int_{0}^{t}\left(\mu^{\prime}(\rho)\int_{\rho}^{t}\frac{dr}{(r-\rho)^{2s_{1}}}\right)\,d\rho\\ &=C\int_{0}^{t}\mu^{\prime}(\rho)(t-\rho)^{1-2s_{1}}\,d\rho\\ &\leqslant C\mu(t)\,t^{1-2s_{1}}.\end{split}$
Now, we set
$c_{0}:=2^{\frac{n}{2s_{1}}-\frac{n^{2}}{(2s_{1})^{2}}}(4C)^{-\frac{n}{2s_{1}}}$
and we claim that
(B.16) $\mu(t)\geqslant c_{0}t^{n}\quad{\mbox{for all }}t\in(0,\delta/4].$
To prove this, we argue by contradiction and assume that there exists $t_{0}\in(0,\delta/4]$ such that
(B.17) $\mu(t_{0})<c_{0}t_{0}^{n}.$
We define $t_{k}:=\frac{t_{0}}{2}+\frac{t_{0}}{2^{k+1}}$. Then, using (B.15), we have that
$\frac{t_{0}}{2^{k+2}}\mu^{1-\frac{2s_{1}}{n}}(t_{k+1})=(t_{k}-t_{k+1})\mu^{1-\frac{2s_{1}}{n}}(t_{k+1})\leqslant\int_{t_{k+1}}^{t_{k}}\mu^{1-\frac{2s_{1}}{n}}(r)\,dr\leqslant C\mu(t_{k})t_{k}^{1-2s_{1}}\leqslant C\mu(t_{k})t_{0}^{1-2s_{1}}.$
Notice that, by continuity,
$\lim_{k\to+\infty}\mu(t_{k})=\mu\left(\frac{t_{0}}{2}\right)=|E\cap B_{t_{0}/2}|>0.$
Therefore, using Lemma B.2 with $x_{k}:=\mu(t_{k})$, $\beta:=s/n$, $M:=4Ct_{0}^{-2s_{1}}$, and $N:=2$, we find that
$\mu(t_{0})>N^{\frac{1}{\beta}-\frac{1}{\beta^{2}}}M^{-\frac{1}{\beta}}=2^{\frac{n}{2s_{1}}-\frac{n^{2}}{(2s_{1})^{2}}}\big(4Ct_{0}^{-2s_{1}}\big)^{-\frac{n}{2s_{1}}}.$
Thus, thanks to (B.17) we deduce that
$c_{0}t_{0}^{n}>\mu(t_{0})>2^{\frac{n}{2s_{1}}-\frac{n^{2}}{(2s_{1})^{2}}}\big(4Ct_{0}^{-2s_{1}}\big)^{-\frac{n}{2s_{1}}}=2^{\frac{n}{2s_{1}}-\frac{n^{2}}{(2s_{1})^{2}}}(4C)^{-\frac{n}{2s_{1}}}t_{0}^{n}=c_{0}t_{0}^{n},$
which gives the desired contradiction and proves (B.16).
So, (B.16) yields
$|E\cap B_{r}|\geqslant c_{0}r^{n}\quad\text{ for all }r\in(0,\delta/4],$
proving the first inequality in (B.1).
Moreover, if $E$ is an almost minimal set, then also $E^{c}$ is almost minimal . Thus, we exploit (B.16) replacing $E$ with $E^{c}$, obtaining that, for $r$ sufficiently small,
$|E^{c}\cap B_{r}|\geqslant c_{0}r^{n},$
from which we infer the second inequality in (B.1). ∎
Appendix C Behavior of $\operatorname{P}_{K}$ in large domains
Here, we show that if $\Omega$ is a Lipschitz domain and
[…]
then $\operatorname{P}_{K}(\Omega_{R})$ is bounded by $R^{n-1}$.
First, let us consider the space of Lipschitz function on $B_{\rho}^{n-1}$, for some $\rho>0$,
$\mbox{Lip}(B_{\rho}^{n-1}):=\{f:B_{\rho}^{n-1}\to\mathbb{R}\mbox{ is Lipschitz-continuous}\},$
endowed with the norm
$\|f\|_{Lip(B_{\rho}^{n-1})}:=\|f\|_{\mathcal{C}^{0}(B_{\rho}^{n-1})}+[f]_{\mathcal{C}^{0}(B_{\rho}^{n-1})},$
where
$f]_{\mathcal{C}^{0}(B_{\rho}^{n-1})}:=\sup_{\begin{subarray}{c}x^{\prime},y^{\prime}\in B_{\rho}^{n-1}\\ x^{\prime}\neq y^{\prime}\end{subarray}}\frac{|f(x^{\prime})-f(y^{\prime})|}{|x-y^{\prime}|}$
denotes the usual Lipschitz seminorm.
Then, we recall the following notion of Lipschitz set (see [MR3791463, Definition 2.4.5]).
Definition C.1.
We say that an open set $\Omega\subset R^{n}$ has Lipschitz boundary (or that $\Omega$ is a Lipschitz set) if there exist $r_{\Omega}>0$, $L_{\Omega}>0$, and $\alpha_{\Omega}>0$ such that for every $x_{0}\in\partial\Omega$ there exist a rotation $\mathcal{R}_{\theta_{x_{0}}}$ (of an angle $\theta_{x_{0}}\in[0,2\pi)$) and a Lipschitz function $\phi_{x_{0}}\in\mbox{Lip}(B_{r_{\Omega}}^{n-1})$ such that:
  1. i)
    $\phi_{x_{0}}]_{Lip(B_{r_{\Omega}}^{n-1})}\leqslant L_{\Omega}$;
  2. ii)
    $\mathcal{R}_{\theta_{x_{0}}}\Omega(-x_{0})\cap\left(B_{r_{\Omega}}^{n-1}\times(-\alpha_{\Omega},\alpha_{\Omega})\right)=\left\{(x^{\prime},x_{n})\mbox{ s.t. }x^{\prime}\in B_{r_{\Omega}}^{n-1}\mbox{ and }x_{n}\in(-\alpha_{\Omega},\phi_{x_{0}}(x^{\prime}))\right\}$;
  3. iii)
    $\partial\left(\mathcal{R}_{\theta_{x_{0}}}\Omega(-x_{0})\right)\cap\left(B_{r_{\Omega}}^{n-1}\times(-\alpha_{\Omega},\alpha_{\Omega})\right)=\left\{(x^{\prime},\phi_{x_{0}}(x^{\prime}))\mbox{ s.t. }x^{\prime}\in B_{r_{\Omega}}^{n-1}\right\}$.
The main result of this section is the following:
Proposition C.2.
There exist positive constants $\eta\in(0,1)$ and $C=C(n,s_{1},s_{2},\kappa_{2},\delta,\eta,\Omega)$ such that
$\operatorname{P}_{K}(\Omega_{R})\leqslant CR^{n-1},$
for every $R\geqslant 1$ large enough.
The proof relies on the following one-dimensional observation:
Lemma C.3.
Let $s\in(1/2,1)$ and $\delta\in(0,1)$. Then,
$\int_{R}^{+\infty}\int_{0}^{R}\frac{\chi_{(\delta,+\infty)}(|x-y|)}{|x-y|^{1+2s}}\,dx\,dy\leqslant\frac{\delta^{1-2s}}{2s-1}.$
Proof.
The claim follows from a direct computation that we provide here for the facility of the reader.
We observe that
$\int_{R}^{+\infty}\int_{0}^{R}\frac{\chi_{(\delta,+\infty)}(|x-y|)}{|x-y|^{1+2s}}\,dx\,dy$ $=\int_{R}^{+\infty}\int_{0}^{\min\{R,x-\delta\}}\frac{dx\,dy}{(x-y)^{1+2s}}$
$=\frac{1}{2s}\int_{R}^{+\infty}\left[\frac{1}{(x-y)^{2s}}\right]_{0}^{\min\{R,x-\delta\}}\,dx.$
Since $x-\delta=\min\{R,x-\delta\}$ if and only if $x\in(R,R+\delta]$, we obtain that
$\int_{R}^{+\infty}\int_{0}^{R}\frac{\chi_{(\delta,+\infty)}(|x-y|)}{|x-y|^{1+2s}}\,dx\,dy=\frac{1}{2s}\left(\int_{R}^{R+\delta}\frac{dx}{\delta^{2s}}+\int_{R+\delta}^{+\infty}\frac{dx}{(x-R)^{2s}}-\int_{R}^{+\infty}\frac{dx}{x^{2s}}\right)$
$\qquad\leqslant\frac{1}{2s}\left(\int_{R}^{R+\delta}\frac{dx}{\delta^{2s}}+\int_{R+\delta}^{+\infty}\frac{dx}{(x-R)^{2s}}\right)=\frac{1}{2s}\left(\delta^{1-2s}+\frac{\delta^{1-2s}}{2s-1}\right)=\frac{\delta^{1-2s}}{2s-1}.\qed$
Another important tool for proving Proposition C.2 is an adaptation of [MR2782803, Lemma 10] for Lipschitz domains. For this, let us define for any $\rho>0$ and $\lambda\geqslant 1$
(C.1) $\begin{split}&C_{\rho,\lambda}:=\{(x^{\prime},x_{n})\in\mathbb{R}^{n}\mbox{ s.t. }|x^{\prime}|<\rho,\ |x_{n}|<\lambda\rho\}=B_{\rho}^{n-1}\times(-\lambda\rho,\lambda\rho),\\ &C_{\rho,\lambda}^{+}:=\{(x^{\prime},x_{n})\in\mathbb{R}^{n}\mbox{ s.t. }|x^{\prime}|<\rho,\ 0<x_{n}<\lambda\rho\}=B_{\rho}^{n-1}\times(0,\lambda\rho),\\ &C_{\rho,\lambda}^{-}:=\{(x^{\prime},x_{n})\in\mathbb{R}^{n}\mbox{ s.t. }|x^{\prime}|<\rho,\ -\lambda\rho<x_{n}<0\}=B_{\rho}^{n-1}\times(-\lambda\rho,0),\\ \mbox{and}\quad&C_{\rho}^{0}:=C_{\rho,\lambda}\cap\{x_{n}=0\}=B_{\rho}^{n-1}\times\{0\}.\end{split}$
Lemma C.4.
Let us consider $s\in(0,1)$, $\rho>0$, and $\lambda\geqslant 1$, and let $\phi\in\mbox{Lip}(B_{\rho}^{n-1})$ be a Lipschitz function such that $\phi=0$ and $\nabla\phi(0)=0$.
Let also
$C_{\rho\lambda}^{+}(\phi):=\{(x^{\prime},x_{n})\in\mathbb{R}^{n}\mbox{ s.t. }|x^{\prime}|<\rho,\ \phi(x^{\prime})<x_{n}<\lambda\rho\},$
and $C_{\rho,\lambda}^{-}(\phi):=\{(x^{\prime},x_{n})\in\mathbb{R}^{n}\mbox{ s.t. }|x^{\prime}|<\rho,\ -\rho\lambda<x_{n}<\phi(x^{\prime})\}).$
If
(C.2) $\phi]_{Lip(B_{\rho}^{n-1})}\leqslant\frac{\lambda}{10},$
then there exist a transformation $\Psi:\mathbb{R}^{n}\to\mathbb{R}^{n}$ such that $\Psi(C_{\rho}^{\pm}(\phi))=C_{\rho}^{\pm}$ and
(C.3) $\iint_{C_{\rho,\lambda}^{+}(\phi)\times C_{\rho,\lambda}^{-}(\phi)}\frac{dxdy}{|x-y|^{n+2s}}\leqslant c\iint_{C_{\rho,\lambda}^{+}\times C_{\rho,\lambda}^{-}}\frac{dxdy}{|x-y|^{n+2s}},$
for some positive constant $c$ independent of $\rho$.
Proof.
Since $\phi$ is Lipschitz and $\phi(0)=0$, we have that
$\sup_{|x^{\prime}|<\rho}\phi(x^{\prime})\leqslant[\phi]_{Lip(B_{\rho}^{n-1})}\rho.$
Then, we define function $\psi\in\mathcal{C}^{\infty}(\mathbb{R}^{n})$ such that
$\psi(x):=\begin{cases}1,\quad&\mbox{if }|x_{n}-\phi(x^{\prime})|\leqslant\frac{\lambda-[\phi]_{Lip(B_{\rho}^{n-1})}}{8}\rho,\\ 0,\quad&\mbox{if }|x_{n}-\phi(x^{\prime})|\geqslant(\lambda-[\phi]_{Lip(B_{\rho}^{n-1})})\rho,\end{cases}$
and
$|\nabla\psi|\leqslant\frac{2}{(\lambda-[\phi]_{Lip(B_{\rho}^{n-1})})\rho}.$
Notice that, for any $|x^{\prime}|\leqslant\rho$, we have that
$|\phi(x^{\prime})\pm\lambda\rho|\geqslant\lambda\rho-|\phi(\rho)|\geqslant(\lambda-[\phi]_{Lip(B_{\rho}^{n-1})})\rho,$
and hence
(C.4) $\psi(x^{\prime},\pm\lambda\rho)=0.$
On the other hand, for any $|x^{\prime}|\leqslant\rho$, we have that
(C.5) $\psi(x^{\prime},\phi(x^{\prime}))=1.$
Now, we consider the transformation $\Psi:\mathbb{R}^{n}\to\mathbb{R}^{n}$ defined as
$\Psi(x):=(x^{\prime},x_{n}-\psi(x)\phi(x^{\prime}))=x-\psi(x)\phi(x^{\prime})e_{n},$
where $e_{n}:=(0,\dots,0,1)$ is the usual unit vector in the $n^{th}$-direction.
By Rademacher’s Theorem, $\phi$ is differentiable almost everywhere, and, if $\widetilde{x}:=\Psi(x)$ and $\widetilde{y}:=\Psi(y)$, we have
$\begin{split}&d\widetilde{x}\,d\widetilde{y}=|1-\partial_{n}\Psi(x)\phi(x^{\prime})|\,|1-\partial_{n}\Psi(y)\phi(y^{\prime})|dx\,dy\\ &\qquad\geqslant\left(1-\left|\partial_{n}\Psi(x)\phi(x^{\prime})+\partial_{n}\Psi(y)\phi(y^{\prime})-\partial_{n}\Psi(x)\phi(x^{\prime})\partial_{n}\Psi(y)\phi(y^{\prime})\right|\right)dx\,dy\\ &\qquad\geqslant(1-c_{1})dx\,dy,\end{split}$
(see also [MR3409135, Theorem 2, Sec. 3.3.3]) with
(C.6) $c_{1}:=\frac{4[\phi]_{Lip(B_{\rho}^{n-1})}^{2}}{(\lambda-[\phi]_{Lip(B_{\rho}^{n-1})})^{2}}+\frac{4[\phi]_{Lip(B_{\rho}^{n-1})}}{\lambda-[\phi]_{Lip(B_{\rho}^{n-1})}}\leqslant\frac{1}{16}+\frac{1}{2}<1,$
where we have employed (C.2) to obtain the last inequality. It thereby follows that
(C.7) $dx\,dy\leqslant\frac{1}{1-c_{1}}d\widetilde{x}\,d\widetilde{y}$.
Besides, recalling (C.4) and (C.5), we deduce that
(C.8) $\Psi(C_{\rho,\lambda}^{\pm}(\phi))=C_{\rho,\lambda}^{\pm}.$
Moreover, for almost every $x,y\in C_{\rho}$, we have that
$\begin{split}&|\widetilde{x}-\widetilde{y}|=|x-\psi(x)\phi(x^{\prime})e_{n}-y+\psi(y)\phi(y^{\prime})e_{n}|\\ &\qquad\leqslant|x-y|+|\psi(x)|\,|\phi(x^{\prime})-\phi(y^{\prime})|+|\phi(y^{\prime})|\,|\psi(x)-\psi(y)|\\ &\qquad\leqslant|x-y|+[\phi]_{Lip(B_{\rho}^{n-1})}|x^{\prime}-y^{\prime}|+[\phi]_{Lip(B_{\rho}^{n-1})}\rho\frac{2}{(\lambda-[\phi]_{Lip(B_{\rho}^{n-1})})\rho}|x-y|\\ &\qquad\leqslant c_{2}|x-y|\end{split}$
for some constant $c_{2}>0$ independent of $\rho$. This yields
(C.9) $|x-y|^{-n-2s}\leqslant c_{2}|\widetilde{x}-\widetilde{y}|^{-n-2s},$
up to renaming $c_{2}$.
Using the change of variable $\widetilde{x}:=\Psi(x)$ and $\widetilde{y}:=\Psi(y)$ together with (C.8), (C.7), and (C.9), we conclude that
$\iint_{C_{\rho,\lambda}^{+}(\phi)\times C_{\rho,\lambda}^{-}(\phi)}\frac{dxdy}{|x-y|^{n+2s}}\leqslant\frac{c_{2}}{1-c_{1}}\iint_{C_{\rho,\lambda}^{+}\times C_{\rho,\lambda}^{-}}\frac{d\widetilde{x}d\widetilde{y}}{|\widetilde{x}-\widetilde{y}|^{n+2s}},$
showing (C.3). ∎
Proof of Proposition C.2.
Let $R>0$ and take $\eta\in(0,1)$ such that
$\delta<\eta R<6\sqrt{n}\eta R<R.$
Since $\Omega$ is bounded, there exists a constant $\zeta>0$ such that $\Omega\subseteq B_{\zeta}$. Up to scaling, we can suppose that $\zeta=1$, otherwise we consider $\widetilde{\eta}:=\eta\zeta\in(0,1)$.
Moreover, let $r_{\Omega}$, $L_{\Omega}$, and $\alpha_{\Omega}$ be as in Definition C.1. We define
$\lambda:=\max\{10L_{\Omega},1\}.$
Wa assume that $R$ is so large that
$3\sqrt{n}\delta<Rr_{\Omega}\quad\mbox{and}\quad 3\sqrt{n}\delta\lambda<R\alpha_{\Omega},$
and that $\eta$ is so small that
$3\sqrt{n}\eta<r_{\Omega}\quad\mbox{and}\quad 3\sqrt{n}\eta\lambda<\alpha_{\Omega}.$
Then, we have
(C.10) $\begin{split}&\operatorname{P}_{K}(\Omega_{R})\\ =&\iint_{\begin{subarray}{c}\Omega_{R}\times\Omega_{R}^{c}\\ \{|x-y|<\delta\}\end{subarray}}K(x,y)\,dx\,dy+\iint_{\begin{subarray}{c}\Omega_{R}\times\Omega_{R}^{c}\\ \{\delta\leqslant|x-y|<\eta R\}\end{subarray}}K(x,y)\,dx\,dy+\iint_{\begin{subarray}{c}\Omega_{R}\times\Omega_{R}^{c}\\ \{|x-y|\geqslant\eta R\}\end{subarray}}K(x,y)\,dx\,dy\\ =&:\mathcal{I}_{1}+\mathcal{I}_{2}+\mathcal{I}_{3}.\end{split}$
We aim to provide estimates for each $\mathcal{I}_{j}$, with $j=1,2,3$.
First, notice that, by compactness, there exists a finite covering of $\partial\Omega_{R}$ made of balls of radius $\delta$,
$\mathcal{B}_{\delta}:=\{B_{\delta}(x_{j})\mbox{ s.t. }x_{j}\in\partial\Omega_{R}\}_{j=1}^{N_{\delta}},$
where $N_{\delta}:=\sharp(\mathcal{B}_{\delta})\leqslant c_{n}\delta^{1-n}R^{n-1}$, for some $c_{n}>0$ depending only on the dimension $n$.
Moreover,
$(\Omega_{R}\times\Omega_{R}^{c})\cap\{|x-y|<\delta\}\subseteq\bigcup_{i,j=0}^{N_{\delta}}B_{\sqrt{n}\delta}(x_{i})\times B_{\sqrt{n}\delta}(x_{j})\cap\{|x-y|<\delta\}$
$\qquad\qquad\subseteq\bigcup_{i=0}^{N_{\delta}}B_{3\sqrt{n}\delta}(x_{i})\times B_{3\sqrt{n}\delta}(x_{i}).$
Besides, since $\Omega_{R}$ has Lipschitz boundary, for every $i$, let $\mathcal{R}_{\theta_{i}}:=\mathcal{R}_{\theta_{x_{i}}}$ be a rotation and let $\phi_{i}:=\phi_{x_{i}}$ be a Lipschitz function as in Definition C.1, so that
$\partial(\mathcal{R}_{\theta_{i}}\Omega_{R}(-x_{i}))\cap C_{3\sqrt{n}\delta,\lambda}=\left\{(x^{\prime},\phi_{i}(x^{\prime}))\mbox{ s.t. }x^{\prime}\in B_{3\sqrt{n}\delta}^{n-1}\right\}.$
Thus, recalling also (1.2) and (1.4), and taking advantage of the rotational symmetries, we infer that
(C.11) $\begin{split}\mathcal{I}_{1}&=\iint_{\begin{subarray}{c}\Omega_{R}\times\Omega_{R}^{c}\\ \{|x-y|<\delta\}\end{subarray}}K(x,y)\,dx\,dy\leqslant\sum_{i=0}^{N_{\delta}}\kappa_{2}\iint_{(\Omega_{R}\cap B_{3\sqrt{n}\delta}(x_{i}))\times(\Omega_{R}^{c}\cap B_{3\sqrt{n}\delta}(x_{i}))}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}\\ &\leqslant\sum_{i=0}^{N_{\delta}}\kappa_{2}\iint_{(\mathcal{R}_{\theta_{i}}\Omega_{R}(-x_{i})\cap B_{3\sqrt{n}\delta})\times(\mathcal{R}_{\theta_{i}}\Omega_{R}^{c}(-x_{i})\cap B_{3\sqrt{n}\delta})}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}\\ &\leqslant\sum_{i=0}^{N_{\delta}}\kappa_{2}\iint_{(\mathcal{R}_{\theta_{i}}\Omega_{R}(-x_{i})\cap C_{3\sqrt{n}\delta,\lambda})\times(\mathcal{R}_{\theta_{i}}\Omega_{R}^{c}(-x_{i})\cap C_{3\sqrt{n}\delta,\lambda})}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}.\end{split}$
Now, thanks to Lemma C.4, for every $i$, there exists a transformation $\Phi_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n}$} such that
$\Phi_{i}(\mathcal{R}_{\theta_{i}}\Omega_{R}(-x_{i})\cap C_{3\sqrt{n}\delta,\lambda})=\{x_{n}<0\}\cap C_{3\sqrt{n}\delta,\lambda}$
and $\Phi_{i}(\mathcal{R}_{\theta_{i}}\Omega_{R}^{c}(-x_{i})\cap C_{3\sqrt{n}\delta,\lambda})=\{x_{n}>0\}\cap C_{3\sqrt{n}\delta,\lambda}.$
and
(C.12) $\begin{split}&\iint_{(\mathcal{R}_{\theta_{i}}\Omega_{R}(-x_{i})\cap C_{3\sqrt{n}\delta,\lambda})\times(\mathcal{R}_{\theta_{i}}\Omega_{R}^{c}(-x_{i})\cap C_{3\sqrt{n}\delta,\lambda})}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}\\ &\qquad\qquad\leqslant c_{1}^{i}\iint_{(\{x_{n}<0\}\cap C_{3\sqrt{n}\delta})\times(\{x_{n}>0\}\cap C_{3\sqrt{n}\delta})}\frac{dx\,dy}{|x-y|^{n+2s_{1}}},\end{split}$
for some positive constant $c_{1}^{i}$ depending only on $n$, $s_{1}$, and $\Phi_{i}$.
Moreover, recalling (C.6) and (C.9) from the proof of Lemma C.4, we infer that
$c_{1}^{i}\leqslant\left(1+[\phi_{i}]_{Lip(B_{r_{\Omega}}^{n-1})}+\frac{2[\phi_{i}]_{Lip(B_{r_{\Omega}}^{n-1})}}{\lambda-[\phi_{i}]_{Lip(B_{r_{\Omega}}^{n-1})}}\right)^{-(n+2s_{1})}\frac{1}{1-\frac{9}{16}}\leqslant\frac{16}{7},$
for every $i$. Let us set
(C.13) $c_{1}:=\frac{16}{7}\iint_{(\{x_{n}<0\}\cap C_{3\sqrt{n}\delta})\times(\{x_{n}>0\}\cap C_{3\sqrt{n}\delta})}\frac{dx\,dy}{|x-y|^{n+2s_{1}}}<+\infty,$
which is independent of $i$.
Thus, plugging(C.12) and (C.13) into (C.11), we arrive at
(C.14) $\mathcal{I}_{1}\leqslant\sum_{i=0}^{N_{\delta}}\kappa_{2}c_{1}^{i}\leqslant\kappa_{2}c_{1}N_{\delta}\leqslant c_{1}R^{n-1},$
up to renaming $c_{1}>0$, now also depending on $\kappa_{2}$ and $\delta$.
Let us now focus on estimating $\mathcal{I}_{2}$. To do this, we cover $\partial\Omega_{R}$ with balls of radius $\eta R$. By compactness, we find a finite covering
$\mathcal{B}_{\eta R}:=\{B_{\eta R}(x_{i})\mbox{ s.t. }x_{i}\in\partial\Omega_{R}\}_{i=0}^{N_{\eta R}},$
with $N_{\eta R}\leqslant c_{n,\Omega}\eta^{1-n}$ (independent of $R$), and
$(\Omega_{R}\times\Omega_{R}^{c})\cap\{|x-y|<\eta R\}\subseteq\bigcup_{i=0}^{N_{\eta R}}\left((\Omega_{R}\cap B_{3\sqrt{n}\eta R})\times(\Omega_{R}^{c}\cap B_{3\sqrt{n}\eta R})\right)\cap\{|x-y|<\eta R\}.$
Therefore,
[…]
Thus, in the same spirit as above, recalling the notation in (C.1), we find a transformation $\Psi_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n}$} such that
$\Psi_{i}(\mathcal{R}_{\theta_{i}}\Omega_{R}(x_{i})\cap C_{3\sqrt{n}\eta R,\lambda})=\{x_{n}<0\}\cap C_{3\sqrt{n}\eta R,\lambda}=B_{3\sqrt{n}\eta R}^{n-1}\times(-3\sqrt{n}\eta\lambda R,0)$
and $\Psi_{i}(\mathcal{R}_{\theta_{i}}\Omega_{R}^{c}(x_{i})\cap C_{3\sqrt{n}\eta R,\lambda})=\{x_{n}>0\}\cap C_{3\sqrt{n}\eta R,\lambda}=B_{3\sqrt{n}\eta R}^{n-1}\times(0,3\sqrt{n}\eta\lambda R).$
For the sake of simplicity, let us call
(C.16) $r:=3\sqrt{n}\eta R.$
Then, using again Lemma C.4 and setting $c_{2}:=\max_{i}c_{2}^{i}$ (which does not depend on $R$), we have
(C.17) $\begin{split}&\iint_{\begin{subarray}{c}(\mathcal{R}_{\theta_{i}}\Omega_{R}(-x_{i})\cap B_{r})\times(\mathcal{R}_{\theta_{i}}\Omega_{R}^{c}(-x_{i})\cap B_{r})\\ \{\delta\leqslant|x-y|<\eta R\}\end{subarray}}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}\\ \leqslant&c_{2}^{i}\iint_{\begin{subarray}{c}(B_{r}^{n-1}\times(-\lambda r,0))\times(B_{r}^{n-1}\times(0,\lambda r))\\ \{\delta\leqslant|x-y|<\eta R\}\end{subarray}}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}\\ \leqslant&c_{2}\iint_{\begin{subarray}{c}(B_{r}^{n-1}\times(-\lambda r,0))\times(B_{r}^{n-1}\times(0,\lambda r))\\ \{\delta\leqslant|x-y|<\eta R\}\end{subarray}}\frac{dx\,dy}{(|x^{\prime}-y^{\prime}|^{2}+|x_{n}-y_{n}|^{2})^{\frac{n+2s_{2}}{2}}},\end{split}$
where we have used the notation $x=(x^{\prime},x_{n})\in\mathbb{R}^{n-1}\times\mathbb{R}$.
Adopting the change of variable $z^{\prime}:=\frac{y^{\prime}-x^{\prime}}{|x_{n}-z_{n}|}$ and $z_{n}:=y_{n}$, we obtain
$\iint_{\begin{subarray}{c}(B_{r}^{n-1}\times(-\lambda r,0))\times(B_{r}^{n-1}\times(0,\lambda r))\\ \{\delta\leqslant|x-y|<\eta R\}\end{subarray}}\frac{dx\,dy}{(|x^{\prime}-y^{\prime}|^{2}+|x_{n}-y_{n}|^{2})^{\frac{n+2s_{2}}{2}}}$
$\qquad\leqslant\int_{B_{r}^{n-1}}\int_{\mathbb{R}^{n-1}}\int_{-\lambda r}^{0}\int_{0}^{\lambda r}\chi_{(\delta,\eta R)}\left(|x_{n}-z_{n}|\sqrt{1+|z^{\prime}|^{2}}\right)\frac{|x_{n}-z_{n}|^{-1-2s_{2}}}{(1+|z^{\prime}|^{2})^{\frac{n+2s_{2}}{2}}}\,dx^{\prime}\,dz^{\prime}\,dx_{n}\,dz_{n}$
$\qquad\leqslant\int_{B_{r}^{n-1}}\int_{\mathbb{R}^{n-1}}\int_{-\lambda r}^{0}\int_{0}^{\lambda r}\chi_{(\delta,+\infty)}\left(|x_{n}-z_{n}|\sqrt{1+|z^{\prime}|^{2}}\right)\frac{|x_{n}-z_{n}|^{-1-2s_{2}}}{(1+|z^{\prime}|^{2})^{\frac{n+2s_{2}}{2}}}\,dx^{\prime}\,dz^{\prime}\,dx_{n}\,dz_{n}$
$\qquad=\int_{B_{r}^{n-1}}\int_{\mathbb{R}^{n-1}}\int_{-\lambda r}^{0}\int_{0}^{\lambda r}\chi_{\left(\delta/\sqrt{1+|z^{\prime}|^{2}},+\infty\right)}(|x_{n}-z_{n}|)\frac{|x_{n}-z_{n}|^{-1-2s_{2}}}{(1+|z^{\prime}|^{2})^{\frac{n+2s_{2}}{2}}}\,dx^{\prime}\,dz^{\prime}\,dx_{n}\,dz_{n},$
where the notation $B_{r}^{n-1}$ stands for the ball of radius $r$ in $\mathbb{R}^{n-1}$.
Now, thanks to Lemma C.3, we have that
$\int_{-\lambda r}^{0}\int_{0}^{\lambda r}\frac{\chi_{\left(\delta/\sqrt{1+|z^{\prime}|^{2}},+\infty\right)}(|x_{n}-z_{n}|)}{|x_{n}-z_{n}|^{1+2s_{2}}}\,dx_{n}\,dz_{n}\leqslant\frac{1}{2s_{2}-1}\left(\frac{\delta}{\sqrt{1+|z^{\prime}|^{2}}}\right)^{1-2s_{2}}.$
Therefore,
$\iint_{\begin{subarray}{c}(B_{r}^{n-1}\times(-\lambda r,0))\times(B_{r}^{n-1}\times(0,\lambda r))\\ \{\delta\leqslant|x-y|<\eta R\}\end{subarray}}\frac{dx\,dy}{(|x^{\prime}-y^{\prime}|^{2}+|x_{n}-y_{n}|^{2})^{\frac{n+2s_{2}}{2}}}$
$\qquad\leqslant\frac{\delta^{1-2s_{2}}}{2s_{2}-1}\int_{B_{r}^{n-1}}\int_{\mathbb{R}^{n-1}}\frac{dx^{\prime}\,dz^{\prime}}{(1+|z^{\prime}|^{2})^{\frac{n+1}{2}}}\leqslant c_{2}r^{n-1}\leqslant c_{2}R^{n-1},$
up to changing $c_{2}>0$ at every step, where we used (C.16) in the last inequality.
Plugging this information into (C.17), we conclude that
$\iint_{\begin{subarray}{c}(\mathcal{R}_{\theta_{i}}\Omega_{R}(-x_{i})\cap B_{r})\times(\mathcal{R}_{\theta_{i}}\Omega_{R}^{c}(-x_{i})\cap B_{r})\\ \{\delta\leqslant|x-y|<\eta R\}\end{subarray}}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}\leqslant c_{2}R^{n-1}.$
It follows from this and (C.15) that
(C.18) $\mathcal{I}_{2}\leqslant\sum_{i=0}^{N_{\eta R}}\kappa_{2}c_{2}R^{n-1}=N_{\eta R}\kappa_{2}c_{2}R^{n-1}\leqslant c_{n,\Omega}\eta^{1-n}\kappa_{2}c_{2}R^{n-1}\leqslant c_{2}R^{n-1},$
up to renaming $c_{2}$.
At last, we exhibit an estimate for $\mathcal{I}_{3}$. By virtue of (1.4), we have that
(C.19) $\begin{split}\mathcal{I}_{3}&\leqslant\kappa_{2}\iint_{\begin{subarray}{c}\Omega_{R}\times\Omega_{R}^{c}\\ \{|x-y|\geqslant\eta R\}\end{subarray}}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}\\ &=\kappa_{2}\iint_{\begin{subarray}{c}\Omega_{R}\times(B_{(1+\eta)R}\setminus\Omega_{R})\\ \{|x-y|\geqslant\eta R\}\end{subarray}}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}+\kappa_{2}\iint_{\begin{subarray}{c}\Omega_{R}\times B_{(1+\eta)R}^{c}\\ \{|x-y|\geqslant\eta R\}\end{subarray}}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}.\end{split}$
Notice that
(C.20) $\iint_{\begin{subarray}{c}\Omega_{R}\times(B_{(1+\eta)R}\setminus\Omega_{R})\\ \{|x-y|\geqslant\eta R\}\end{subarray}}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}\leqslant\frac{|\Omega_{R}||B_{(1+\eta)R}\setminus\Omega_{R}|}{(\eta R)^{n+2s_{2}}}\leqslant c_{3}R^{n-2s_{2}},$
for some positive $c_{3}=c_{3}(n,s_{2},\eta)$.
Additionally, by the change of variable z:=yxz:=y-x,
$\iint_{\begin{subarray}{c}\Omega_{R}\times B_{(1+\eta)R}^{c}\\ \{|x-y|\geqslant\eta R\}\end{subarray}}\frac{dx\,dy}{|x-y|^{n+2s_{2}}}\leqslant\iint_{\Omega_{R}\times B_{\eta R}^{c}}\frac{dx\,dz}{|z|^{n+2s_{2}}}\leqslant c_{3}R^{n-2s_{2}}.$
This inequality and (C.20), together with (C.19), yield that
(C.21) $\mathcal{I}_{3}\leqslant\kappa_{2}c_{3}R^{n-2s_{2}}.$
As a result of (C.10), (C.11), (C.18), and (C.21), we obtain that
$\operatorname{P}_{K}(\Omega_{R})\leqslant c_{1}R^{n-1}+c_{2}R^{n-1}+c_{3}R^{n-2s_{2}},$
concluding the proof. ∎
Appendix D Well-posedness of the Plateau problem for the energy $\mathscr{F}$
In this section, we use a Sobolev embedding and a simple covering argument to show that the minimization problem for the energy functional $\mathscr{F}$ with fixed external datum is well-posed. To do this, let us consider a Lipschitz domain $\Omega\subseteq\mathbb{R}^{n}$ and a set $\widetilde{E}\subseteq\mathbb{R}^{n}$ of finite $K$-perimeter. Suppose $K$ satisfies (1.2), (1.3), (1.4), and (1.5). Let also $g\in L^{\infty}(\mathbb{R}^{n})$ be a $\mathbb{Z}^{n}$-periodic function satisfying (1.1).
We recall that the functional $\mathscr{F}$ is defined in (1.7) as
$\mathscr{F}(E,\Omega):=\operatorname{P}_{K}(E,\Omega)+\int_{E\cap\mathcal{Q}(\Omega)_{1}}g(x)\,dx$
and we investigate the problem:
[…]
We claim that:
Proposition D.1 (Well-posedness of the minimization problem (D.1)).
There exists a minimizer for $\mathscr{F}$ in $\Omega$ with external datum $\widetilde{E}\setminus\Omega$.
This result is a consequence of a standard compactness argument, whose precise statement goes as follows:
Lemma D.2.
Let $\{E_{j}\}_{j}$ be a sequence of sets such that $E_{j}\setminus\Omega=\widetilde{E}\setminus\Omega$ and
$\sup_{j}\mathscr{F}(E_{j},\Omega)<+\infty.$
Then, there exist a set $E\subseteq\mathbb{R}^{n}$ and a subsequence $\{E_{j_{m}}\}_{m}$ such that $E\setminus\Omega=\widetilde{E}\setminus\Omega$ and
$E_{j_{m}}\to E\mbox{ in }L^{1}_{\rm{loc}}(\mathbb{R}^{n}),\quad\mbox{as }m\to+\infty.$
Proof.
Let us consider the covering $\{B_{\delta/2}(x)\}_{x\in\overline{\Omega}}$ of $\Omega$ made of balls of radius $\delta/2$. By compactness, there exist finitely many points $x_{1},\dots,x_{N}\in\overline{\Omega}$ such that, setting $B^{(\ell)}:=B_{\delta/2}(x_{\ell})\cap\Omega$, we have
$\Omega\subseteq\bigcup_{\ell=1}^{N}B^{(\ell)},$
and $B^{(\ell)}\cap B^{(\ell+1)}\neq\varnothing,\mbox{ for every }\ell\in\{1,\dots,N-1\}.$
Now, observe that, in virtue of (1.4), for any $\ell\in\{1,\dots,N\}$ and for any set $F$,
$\begin{split}&\frac{\kappa_{1}}{2}[\chi_{F}]^{2}_{H^{s_{1}}(B^{(\ell)}}=\frac{\kappa_{1}}{2}\iint_{B^{(\ell)}\times B^{(\ell)}}\frac{|\chi_{F}(x)-\chi_{F}(y)|^{2}}{|x-y|^{n+s_{1}}}\,dx\,dy\\ &\qquad\leqslant\frac{1}{2}\iint_{B^{(\ell)}\times B^{(\ell)}}|\chi_{F}(x)-\chi_{F}(y)|^{2}K(x,y)\,dx\,dy\\ &\qquad=\iint_{B^{(\ell)}\times B^{(\ell)}}\chi_{F}(x)\chi_{F^{c}}(y)K(x,y)\,dx\,dy\leqslant\iint_{\Omega\times\Omega}\chi_{F}(x)\chi_{F^{c}}(y)K(x,y)\,dx\,dy\\ &\qquad\leqslant\operatorname{P}_{K}(F,\Omega)=\mathscr{F}(F,\Omega)-\int_{F\cap\mathcal{Q}(\Omega)_{1}}g(x)\,dx\leqslant\mathscr{F}(F,\Omega)+\|g\|_{L^{\infty}(\Omega)}|\Omega|,\end{split}$
where $\cdot]_{H^{s}}$ denotes the $(2s)$-Gagliardo seminorm.
In particular, we deduce that, for all $j\in\mathbb{N}$,
$\frac{\kappa_{1}}{2}[\chi_{E_{j}}]^{2}_{H^{s_{1}}(B^{(1)})}\leqslant\sup_{j}\mathscr{F}(E_{j},\Omega)+\|g\|_{L^{\infty}(\Omega)}|\Omega|.$
Hence, by the compactness of the Sobolev embedding $H^{s_{1}}(B^{(1)})\hookrightarrow\hookrightarrow L^{2}(B^{(1)})\subseteq L^{1}(B^{(1)})$ (see [MR2944369, Corollary 7.2]), there exist a set $E_{1}\subseteq B^{(1)}$ and a subsequence $\{E_{j_{m}}\}_{m}$ such that $E_{j_{m}}\to E_{1}$ in $L^{1}(B^{(1)})$ and pointwise in $B^{(1)}$.
Similarly,
$\frac{\kappa_{1}}{2}[\chi_{E_{j_{m}}}]^{2}_{H^{s_{1}}(B^{(2)})}\leqslant\sup_{m}\mathscr{F}(E_{j_{m}},\Omega)+\|g\|_{L^{\infty}(\Omega)}|\Omega|.$
Therefore, up to considering a further subsequence, we also have that $E_{j_{m}}\to E_{2}$ in $L^{1}(B^{(2)})$ and pointwise in $B^{(2)}$, for some set $E_{2}\subseteq B^{(2)}$. Moreover, by the uniqueness of the limit, it follows that $E_{1}=E_{2}$ in $B^{(1)}\cap B^{(2)}$.
Repeating the same argument in every ball $B^{(\ell)}$, up to considering further subsequences, we infer that $E_{j_{m}}\to E$ in $L^{1}(\Omega)$, where
$E:=\left(\bigcup_{\ell=1}^{N}E_{\ell}\right)\cup\widetilde{E}\setminus\Omega,$
concluding the proof. ∎
Proof of Proposition D.1.
First, notice that the energy functional $\mathscr{F}$ is uniformly bounded from below. Indeed, for any set $F$ we have
$\mathscr{F}(F,\Omega)\geqslant\int_{F\cap\mathcal{Q}(\Omega)_{1}}g(x)\,dx\geqslant-\|g\|_{L^{\infty}(\Omega)}|\Omega|.$
Thus, we consider a minimizing sequence $\{E_{j}\}_{j}$ such that $E_{j}\setminus\Omega=\widetilde{E}\setminus\Omega$, and
$\mathscr{F}(E_{j},\Omega)\to\inf\{\mathscr{F}(F,\Omega)\quad\mbox{s.t. }F\setminus\Omega=\widetilde{E}\setminus\Omega\},\quad\mbox{as }j\to+\infty.$
Now we take
$F:=\begin{cases}\Omega&\quad{\mbox{ in }}\Omega,\\ \widetilde{E}\setminus\Omega&\quad{\mbox{ in }}\mathbb{R}^{n}\setminus\Omega\end{cases}$
and we observe that, for any $j$ large enough,
$\mathscr{F}(E_{j},\Omega)\leqslant\mathscr{F}(F,\Omega)=\operatorname{P}_{K}(F,\Omega)+\int_{F\cap\mathcal{Q}(\Omega)_{1}}g(x)\,dx\leqslant\operatorname{P}_{K}(\Omega)+\|g\|_{L^{\infty}(\Omega)}|\Omega|<+\infty.$
Therefore, the desired result follows from Lemma D.2 and the lower semi-continuity of the functional $\mathscr{F}$. ∎
Acknowledgments
Matteo Novaga is member of the INDAM-GNAMPA, and was supported by Next Generation EU, PRIN 2022E9CF89. Serena Dipierro and Enrico Valdinoci are members of the Australian Mathematical Society (AustMS). Serena Dipierro and Riccardo Villa are supported by the Australian Research Council Future Fellowship FT230100333 “New perspectives on nonlocal equations”. Enrico Valdinoci is supported by the Australian Laureate Fellowship FL190100081 “Minimal surfaces, free boundaries and partial differential equations”. Riccardo Villa is supported by a Scholarship for International Research Fees at the University of Western Australia.